### v1

# A Survey on Autonomous Driving Datasets: Data Statistic, Annotation, and Outlook

Abstract
Autonomous driving has rapidly developed and shown promising performance with recent advances in hardware and deep learning methods. High-quality datasets are fundamental for developing reliable autonomous driving algorithms. Previous dataset surveys tried to review the datasets but either focused on a limited number or lacked detailed investigation of the characters of datasets. To this end, we present an exhaustive study of over 200 autonomous driving datasets from multiple perspectives, including sensor modalities, data size, tasks, and contextual conditions. We introduce a novel metric to evaluate the impact of each dataset, which can also be a guide for establishing new datasets. We further analyze the annotation process and quality of datasets. Additionally, we conduct an in-depth analysis of the data distribution of several vital datasets. Finally, we discuss the development trend of the future autonomous driving datasets.
Index Terms:
Dataset, influence, annotation, autonomous driving.
I Introduction
Autonomous driving (AD) aims to revolutionize the transportation system by creating vehicles that can accurately perceive their environment, make intelligent decisions, and drive safely without human intervention. Due to thrilling technical development, various autonomous driving products have been implemented in several fields, such as robotaxi [1]. These rapid advancements in autonomous driving rely heavily on extensive datasets, which help autonomous driving systems be robust and reliable in complex driving environments.
In recent years, there has been a significant increase in the quality and variety of autonomous driving datasets. The first apparent phenomenon in the development of datasets is the various data collection strategies, including synthetic datasets [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12] generated by simulators and recorded from the real world [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], to name just a few. Secondly, the datasets vary in composition, including but not limited to multiple sensory data (like camera images and LiDAR point clouds) and different annotation types for various tasks in autonomous driving. Fig. 1 depicts the statistic of the 3D object bounding box distribution of six real-world datasets (Argoverse 2[28], KITTI [13], nuScenes [22], ONCE [29], Waymo [23], and ZOD [30]) under a bird’s-eye-view perspective, showing each dataset’s distinguished annotation characters. According to the equipment position of sensors, the variety of datasets also reflects in the sensing domains, containing onboard, V2X, drone, and others. Furthermore, the geometrical diversity and weather conditions improve the generalization of the autonomous driving datasets.
Figure 1: BEV object distribution of datasets. Each heatmap corresponds to a dataset plotted with X and Y coordinates. Y is the driving direction of the ego-vehicle. The unique annotation characters of each dataset are reflected in the distribution range, density, and number of bounding boxes.
I-A Research Gap & Motivation
We demonstrate the yearly published number of perception datasets in Fig. 2 to reflect the trend of autonomous driving datasets from one perspective. Since numerous and continually increasing numbers of publicly published datasets exist, an exhaustive survey on autonomous driving datasets is valuable for advancing academic and industrial research. In prior work, Yin et al. [31] summarized 27 publicly available datasets containing data collected on public roads. As a sequential work of  [31], [32] extended the dataset number. [33] and [34] proposed systematical introductions to the existing datasets from an application perspective. Beyond describing existing datasets, [35] discussed the domain adaptation between synthetic and real data and automatic labeling methods. [36] summarized existing datasets and undertook an exhaustive analysis of the characters of the next-generation datasets. However, these surveys only summarized a small number of datasets, causing a non-wide scope. AD-Dataset [37] collected a large number of datasets while lacking detailed analysis for the attributes of these datasets. Compared to studies on all types of datasets, some researchers presented surveys on a particular type of autonomous driving dataset, such as anomaly detection [38], synthetic datasets [39], 3D semantic segmentation [40], and decision-making [41].
To this end, we aim to propose a comprehensive and systematic study on a large number of datasets in autonomous driving, covering all tasks from perception to control, considering real-world as well as synthetic data, and providing insight into the data modality and quality of several crucial datasets. We illustrate the comparison between other dataset surveys and ours in Tab I.
Survey General Data Annotation
Year #Datasets Tasks S. domain S. moda. Geo. Env. Data analysis Quality Process
When to use what dataset [31] 2017 27 Perc $\checkmark$ $\checkmark$ $\checkmark$
Self-driving Algorithm [32] 2019 37 Perc $\checkmark$ $\checkmark$ $\checkmark$
Is it safe to drive [33] 2019 54 Perc, Pred, E2E $\checkmark$ $\checkmark$ $\checkmark$
CV for AVs [34] 2020 33 Perc $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
A Survey on AD Datasets [35] 2021 30 Perc $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
3D Semantic Segmentation [40] 2021 29 Perc $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AD-Dataset [37] 2022 204 Perc, Pred, Pl, C $\checkmark$ $\checkmark$ $\checkmark$
Anomaly Detection [38] 2023 16 Perc $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Synthetic Datasets for AD [39] 2023 17 Perc, Pred $\checkmark$ $\checkmark$ $\checkmark$
Decision-making [41] 2023 25 Pl, C $\checkmark$ $\checkmark$ $\checkmark$
Open-sourced Data Ecosystem [36] 2023 70 Perc, Pred, Pl $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Ours 235 Perc, Pred, Pl, C, E2E $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
TABLE I: We compare our survey paper with other AD dataset surveys in the following perspectives: collected dataset number (#Dataset), relevant tasks, sensing domain (S. domain), sensor modality (S. moda.), geometric conditions (Geo.), environmental conditions (Env.), analyzing data distribution, introducing annotation quality and process. We describe the task types in a coarse granularity, including perception (Perc.), prediction (Pred.), planning (Pl.), control (C.) and end-to-end (E2E).
Figure 2: Overview of the trend of publishing datasets. The diagram exhibits a rapid growth of onboard datasets between 2015 and 2020 and then slowly down. In contrast, there is an increment trend for the V2X datasets, showing the research trend on the cooperative perception systems.
I-B Main Contributions
The main contributions of this paper can be summarized as follows:
  • We present an exhaustive survey on autonomous driving datasets. We consider publicly available datasets as comprehensively as possible, recording their fundamental characteristics, such as published year, data size, sensor modalities, sensing domains, geometrical and environmental conditions, and support tasks. To the best of our knowledge, we provide an overview of the most extensive collection of autonomous driving datasets recorded to date.
  • We systematically illustrate the sensors and sensing domains for collecting autonomous driving data. Furthermore, we describe the main tasks in autonomous driving, including task goals, required data modalities, and evaluation metrics.
  • We summarize and divide the datasets according to their sensing domains and supporting tasks, which assists researchers in efficiently selecting and gathering information for the goal datasets. Thereby facilitating more targeted and effective research and development efforts.
  • Additionally, we introduce an impact score metric to evaluate the influence of published perception datasets in the community. This metric can also be leveraged as a guidance for developing future datasets. We deeply analyze datasets with the highest grades based on the impact score, highlighting their advantages and utility.
  • We investigate the annotation quality of datasets and the existing labeling procedures for various autonomous driving tasks.
  • Detailed data statistics is conducted to demonstrate the data distribution of various datasets from different perspectives, exhibiting their inherent limitations and suitable use situations.
  • We analyze the recent technology trend and demonstrate the development direction of the next generation of datasets. We also provide a prospect on the potential effect of the advancement of Large Language Models on future autonomous driving.
I-C Scope & Limitations
We aimed to conduct an exhaustive survey on the existing autonomous driving datasets to provide assistance and guidance for developing future algorithms and datasets in the domain. We collected datasets focusing on the four basic autonomous driving tasks: perception, prediction, planning, and control. Because several versatile datasets support multiple tasks, we only explained them in the main scope they support to avoid repeat introduction. Additionally, we collected a large number of datasets and exhibited them with their primary characters in tables. Nevertheless, a detailed explanation of all collected datasets could not highlight the most popular ones, and it can hurt assisting researchers in finding valuable datasets through this survey. Hence, we only describe the most impactful datasets in detail.
I-D Structure of the Article
The rest of the article is structured as follows: Section II introduces the approach we leveraged to obtain public datasets and the evaluation metrics for datasets. Section III demonstrates the primary sensors used in autonomous driving and their modalities. Section IV discusses autonomous driving tasks, related challenges, and required data. We further discuss several important datasets in Section V. The process of annotations and factors affecting annotation quality are exhibited in Section VI. Moreover, we statistic the data distribution of several datasets in section VII. In section VIII, we investigate the developing trend and future works of autonomous driving datasets. In the end, we conclude our work in Section IX. The taxonomy of this survey is shown in Fig. 3.
Figure 3: This survey’s primary taxonomy includes impact score, sensors and modalities, autonomous driving tasks, important datasets, and annotation process.
II Methodology
This section consists of 1) How we collect and filter datasets (II-A), and 2) How to evaluate the impact of a dataset on the autonomous driving domain (II-B).
II-A Datasets Collection
Following [42], we conducted a systematic review to exhaustively collect published autonomous driving datasets.
To ensure source diversity, we utilized well-known search engines such as Google [^1] https://www.google.com/, Google Scholar [^2] https://scholar.google.com/ and Baidu [^3] https://www.baidu.com/ to search datasets. To ensure a thorough dataset collection from various countries and regions, we conducted searches in English, Chinese, and German using keywords such as ”autonomous driving datasets,” ”intelligent vehicle datasets,” and terms related to object detection, classification, tracking, segmentation, prediction, planning, and control.
Furthermore, we explored IEEE Xplore [^4] https://ieeexplore.ieee.org/Xplore/home.jsp and pertinent conferences in autonomous driving and intelligent transportation systems to collate datasets from journals and conference proceedings. We verified datasets from these sources through keyword searches and manual title reviews.
Finally, to ensure the inclusion of specialized or lesser-known datasets, we searched through Github [^5] https://github.com/ repositories and Paperwithcodes [^6] https://paperswithcode.com/. Similar to databases, we performed both manual and keyword-based searches for datasets.
II-B Dataset Evaluation Metrics
We introduce a novel metric, impact score, to assess the significance of a published dataset, which can also be a guide to preparing a new dataset. In this section, we explain in detail the approach to calculate the impact score of autonomous driving datasets.
For a fair and compatible comparison, we only consider datasets related to the perception domain, which takes up a large portion of autonomous driving datasets. Additionally, to ensure the objectivity and comprehensibility of our scoring system, we take into account various factors, including citation score, data dimension, and environmental diversity. All the values are gathered from official papers or open-source dataset websites.
Citation Score. First, we calculate the citation scores from the total citation number and average annual citation. To gain fair citation counts, we choose the time of the earliest version of a dataset as its publish time. Moreover, all citation counts were collected as of September 20, 2023, to ensure the comparison is based on a consistent timeframe. The total citation number $\mathbf{c}^{t}$ reflects the overall influence of a dataset. A higher count of this metric means that the dataset has been widely recognized and utilized by researchers. However, datasets published in earlier years can accumulate more citations. To address this unfairness, we leverage average annual citation $\mathbf{c}^{a}$, which describes the yearly citation increase speed of a dataset. The calculate function is shown in Eq. 1.
$\mathbf{c}^{a}=\begin{cases}\mathbf{c}^{t}/\left(\mathbf{y}_{curr}-\mathbf{y}_{pub}\right)&\text{if }\mathbf{y}_{curr}\neq\mathbf{y}_{pub}\\ \mathbf{c}^{t}&\text{if }\mathbf{y}_{curr}=\mathbf{y}_{pub}\end{cases}$ (1)
where $\mathbf{y}_{curr}$ and $\mathbf{y}_{pub}$ represent the current year and the dataset published year, respectively. On the other hand, the citation number of distastes has a wide distribution range from single digits to tens of thousands. To alleviate the extreme unbalance and highlight the differences of each dataset, we apply logarithmic transformation followed by Min-Max normalization to both $\mathbf{c}^{t}$ and $\mathbf{c}^{a}$, described in Eq. 2.
$\mathbf{c}_{norm}=\min\text{-}\max\left(\log(\mathbf{c})\right)$ (2)
Finally, the citation score $\mathbf{c}_{score}$ is the summation of $\mathbf{c}^{t}_{norm}$ and $\mathbf{c}^{a}_{norm}$:
$\mathbf{c}_{score}=0.5\*\mathbf{c}^{t}_{norm}+0.5\*\mathbf{c}^{a}_{norm}$ (3)
Data Dimension Score. We measure the data dimension across four perspectives: dataset size, temporal information, task number, and labeled categories. Dataset size $\mathbf{f}$ is represented by the frame number of a dataset, reflecting its volume and comprehensiveness. To get the dataset size score $\mathbf{f_{normal}}$, we leverage the same method as the citation score to process the frame number to overcome the extreme imbalance between different datasets.
Temporal information is essential for autonomous driving as it enables the vehicle to understand how the surrounding environment changes over time. We use $\mathbf{t}\in\left\{0,1\right\}$ to indicate whether a dataset includes temporal information. Regarding the task number, we only consider datasets related to the six fundamental tasks in the autonomous driving perception domain, such as 2D object detection, 3D object detection, 2D semantic segmentation, 3D semantic segmentation, tracking, and lane detection. Therefore, the task number score is recorded as $\mathbf{t_{n}}\in\left\{1,2,3,4,5,6\right\}$. A large number of categories is critical for the robustness and versatility of a dataset. During the statistic, if a dataset supports multiple tasks and includes various types of annotation, we choose the largest number of categories. Afterward, the categories are divided into five levels, $\mathbf{l}=\left\{1,2,3,4,5\right\}$, based on quintiles. To simplify the calculation, we normalize $\mathbf{t_{n}}$ and $\mathbf{l}$ before the following process.
In order to reflect the data dimension score $\mathbf{d}_{score}$ as objectively as possible, we give different weights to the four components, as shown in Eq. 4.
$\mathbf{d}_{score}=0.5\*\mathbf{f}_{norm}+0.1\*\mathbf{t}+0.2\*\mathbf{t_{n}}_{norm}+0.2\*\mathbf{l}_{norm}$ (4)
Environmental Diversity Score. We evaluate the environmental diversity of a dataset according to the following factors: 1) weather conditions, such as rain or snow. 2) time of day refers to when the data is collected, like morning or dusk. 3) the types of driving scenarios, e.g., urban or rural. 4) the geometric scope means the number of countries or cities where the data is recorded. It is worth noting that we treat the geometric scope for synthetic datasets as missing. We follow the granularity with which a paper categorizes its data to quantify the diversity. Moreover, for the missing value, if a dataset announces that the data is recorded under diversity conditions, we use the median value as the missing value. Otherwise, we set the missing value of this attribute to one. We apply quintiles to quantify each factor into five distinct levels. After that, the environmental diversity score $\mathbf{e}_{score}$ is the sum of these four factors.
In the end, we leverage Eq. 5 to calculate the impact score $\mathbf{i}_{score}$.
$\mathbf{i}_{score}=60\*\mathbf{c}_{score}+20\*\mathbf{d}_{score}+20\*\mathbf{e}_{score}$ (5)
The total impact score is 100, and 60 percent belongs to the citation score $\mathbf{c}_{score}$, data dimension score $\mathbf{d}_{score}$ and environmental diversity score $\mathbf{e}_{score}$ takes 40 percent.
III Data sources and Cooperative Perception in Autonomous Driving
In this section, we introduce the sensors and their modalities mainly used in autonomous driving (III-A). Furthermore, in III-B, we analyze the data acquisition and communication domain, such as onboard, drone, and vehicle-to-everything.
III-A Sensors and Modalities of Data
(a) Camera
(b) LiDAR
(c) Radar
(d) Event camera
(e) IMU
(f) Thermal camera
Figure 4: Sensors on autonomous driving vehicle. The type of each sensor is (a) Camera: Basler ace acA1600-20uc, (b) LiDAR: Velodyne Puck LITE, (c) Radar: Ainstein Launches K-79, (d) Event-based camera: Evaluation Kit 4 HD, (e) IMU: IMU383_Aceinna-W and (f) Thermal camera: FLIR_2nd_Gen_ADK. All figures are extracted from the websites hosting the sensors.
Figure 5: Modalities of sensors. We demonstrate each type of sensor’s modality to intuitively understand each sensor’s characters.
Efficiently and precisely collecting data from the surrounding environment is the key to a reliable perception system in autonomous driving. To achieve this goal, various types of sensors are utilized on self-driving vehicles and infrastructures. The sensor examples are shown in Fig. 4. The most used sensors are cameras, LiDAR, and radars. Event-based and thermal cameras are also installed in vehicles or roadsides to improve the perception capability further.
RGB Images. RGB images are usually recorded by monocular, stereo, or fisheye cameras. Monocular cameras offer a 2D view without depth; stereo cameras, with their dual lenses, provide depth perception; fisheye cameras use wide-angle lenses to capture a broad view. All these cameras channel light through a lens onto an image sensor, e.g., CMOS, converting this light into electronic signals representing an image. As shown in Fig. 5 (a), the 2D images capture color information, rich textures, patterns, and visual details of the environment. Due to these characters, RGB images are mainly used to detect vehicles and pedestrians and recognize road signs. However, the RGB images are vulnerable to conditions like low illumination, rain, fog, or glare [43].
LiDAR Point Clouds. LiDAR uses laser beams to measure the distance between the sensor and an object, creating a 3D environment representation [44]. The LiDAR point clouds (Fig. 5 (b)) provide precise spatial information with high resolution and can detect objects over long distances. However, the density of these points can decrease with increasing distance, leading to sparser representations for distant objects. The weather conditions, e.g., fog, also limit the performance of LiDAR. In general, LiDAR is suitable in cases that require 3D concise information.
Radar Point Clouds. Radars detect objects, distance, and relative speed by emitting radio waves and analyzing their reflection. Moreover, radars are robust under various weather conditions [45]. Nevertheless, radar point clouds are generally coarser than LiDAR data, lacking the detailed shape or texture information of objects. Therefore, radars are generally used to assist other sensors. Fig. 5 (c) exhibits the radar point clouds.
Event of Event-based Camera. Event-based cameras asynchronously capture data, activating only when a pixel detects a change in brightness. The captured data is called events (Fig. 5 (d)). Thanks to the specific data generation method, the recorded data has extremely high temporal resolution and captures fast motion without blur [46].
Infrared Images of Thermal Camera. Thermal cameras (see Fig. 5 (e)) detect heat signatures by capturing infrared radiation [47]. Due to producing images based on temperature differences, thermal cameras can work in total darkness and are unaffected by fog or smoke. However, thermal cameras cannot discern colors or detailed visual patterns evident. Furthermore, the resolution of infrared images is lower compared to optical cameras.
Inertial Measurement Unit (IMU). An IMU is an electronic device that measures and reports an object’s specific force, angular rate, and sometimes magnetic field surrounding the object [48]. In autonomous driving, it is used to track the movement and orientation of the vehicle. Although IMU does not include visual information of the surrounding environment, the perception systems can achieve more accurate and robust tracking of a vehicle’s motion and orientation by fusing the data from an IMU with data from other sensors.
Figure 6: Sensor number distribution. We statistic the distribution of different sensors. Overall, RGB cameras and LiDAR are the most used sensors in autonomous driving datasets.
We analyze the sensor distribution from the collected datasets, shown in Fig. 6. More than half of the sensors are monocular cameras (53.85%) due to their low price and reliable performance. Additionally, 93 datasets include LiDAR data, valued for its high resolution and precise spatial information. However, its high-cost limits LiDAR’s widespread use. Beyond LiDAR point clouds, 29 datasets leverage stereo cameras to capture depth information. Furthermore, 5.41%, 3.42%, and 1.71% datasets include radar, thermal camera, and fisheye camera, respectively. Given the temporal efficiency of capturing dynamic scenes, three datasets generate data based on event-based cameras.
III-B Sensing Domains and Cooperative Perception Systems
The sensory data and communication between the ego vehicle and other entities in the surrounding environment play a pivotal role in ensuring self-driving systems’ safety, efficiency, and overall functionality. Therefore, the positioning of sensors is crucial as it determines the quality, angle, and scope of data that can be collected. Generally, sensors in the autonomous driving environment can be categorized into the following domains: onboard, Vehicle-to-Everything (V2X), drone, and others.
Figure 7: Overview of cooperative perception systems in autonomous driving. A complete autonomous driving perception system consists of the ego vehicle and the cooperation between vehicles, infrastructure, and networks.
Onboard. Onboard sensors are installed directly on the autonomous driving vehicle and usually consist of cameras, LiDAR, radars, and IMUs. These sensors provide a direct perspective from the vehicle’s standpoint, offering immediate feedback on the surroundings. Nevertheless, due to the limitation of the vehicle detection scope, onboard sensors may have limitations in providing advanced warning about obstacles in blind spots or detecting hazards around sharp bends.
V2X. Vehicle-to-Everything (V2X) encompasses communications between a vehicle and any other components in the transport system [49], including V2V, V2I, and V2N (as shown in Fig. 7). Beyond the immediate sensory input, the cooperative systems ensure multiple entities work harmoniously.
  • Vehicle-to-Vehicle (V2V)
    V2V enables nearby vehicles to share data, including their positions and velocity and sensory data, like camera images or LiDAR scans, which assists in creating a more comprehensive understanding of the driving scenarios.
  • Vehicle-to-Infrastructure (V2I)
    V2I facilitates communications between the autonomous vehicle and infrastructure components such as traffic lights, signs, or roadside sensors. The sensors embedded in the road infrastructure, including cameras, LiDAR, radars, or event-based cameras, work synergistically to extend the perception range and improve the situational awareness of autonomous vehicles. In this survey, we consider both perceptions through infrastructure or V2I belonging to V2I.
  • Vehicle-to-Network (V2N)
    V2N refers to exchanging information between a vehicle and a broader network infrastructure, often leveraging cellular networks to provide vehicles with access to cloud data. V2N aids the cooperation perception of V2V and V2I by sharing cross-area data or offering real-time updates about traffic congestion or road closures.
Drone. Drones, or UAVs, offer an aerial perspective, providing data essential for trajectory prediction and route planning [18] . For example, the real-time data from drones can be integrated into traffic management systems to optimize the traffic flow and alert autonomous vehicles of accidents ahead.
Others. Data not collected by the previous three types is defined as others, such as other devices installed on non-vehicle objects or multiple domains.
IV Tasks in Autonomous Driving
This section offers insight into the pivotal tasks in autonomous driving, such as perception and localization (IV-A), prediction (IV-B), and planning and control (IV-C). The overview of the autonomous driving pipeline is demonstrated in Fig. 8. We detail their objectives, the nature of data they rely upon, and inherent challenges. Fig. 9 illustrates examples of several main tasks in autonomous driving.
Figure 8: The overview of autonomous driving pipeline. Autonomous driving systems can be categorized into two types: modular-based and end-to-end. Both rely on the data collected by various sensors installed on the vehicles or infrastructures and interact with the surrounding driving experiments.
IV-A Perception and Localization
(a) 3D object detection
(b) Semantic segmentation
(c) Trajectory prediction
(d) Lane detection
(e) Visual referring
(f) Point cloud segmentation
(g) Driver attention prediction
(h) HD map
Figure 9: Examples of various autonomous driving tasks. (a) is from KITTI [13], (b) is from Cityscapes [14], (c) is from V2X-Seq [50], (d) is from BDD100K [24], (e) is from Refer-KITTI [51], (f) is from KITTI-360 [52], (g) is from Dr(eye)ve [53], (h) is from Argoverse 2 [28]. All figures are collected from the open-source data of datasets or the websites hosting the datasets.
Perception focuses on understanding the environment based on the sensory data, while localization determines the autonomous vehicle’s position within that environment.
2D/3D Object Detection. 2D or 3D object detection aims to identify and classify other entities within the driving environment. While 2D object detection identifies objects in image space, 3D object detection further incorporates precise depth information, often provided by LiDAR. Although the detection technologies have significantly advanced, several challenges remain, such as object occlusions, varying light conditions, and diverse object appearances.
Usually, the Average Precision (AP) metric [54] is applied to evaluate the object detection performance. According to [1], the AP metric can be formulated as
[…]
where $p(r)$ is the precision-recall curve.
2D/3D Semantic Segmentation. Semantic segmentation involves classifying each pixel of an image or point of a point cloud to its semantic category. From a dataset perspective, maintaining fine-grained object boundaries while managing extensive labeling requirements presents significant challenges for this task.
As mentioned in [55], the main metrics used for segmentation are mean Pixel Accuracy (mPA):
$mPA=\frac{1}{k+1}\sum_{i=0}^{k}\frac{p_{ii}}{\sum_{j=0}^{k}p_{ij}}$ (7)
and the mean Intersection over Union (mIoU):
$mIoU=\frac{1}{k=1}\sum_{i=0}^{k}\frac{p_{ii}}{\sum_{j=0}^{k}p_{ij}+\sum_{j=0}^{k}p_{ji}-p_{ii}}$ (8)
where $k\in\mathbb{N}$ is the number of classes, and $p_{ii}$, $p_{ij}$, and $p_{ji}$ represent true positives, false positives, and false negatives, respectively.
Object Tracking. Object Tracking monitors the trajectories of a single or multiple objects over time. This task necessitates time-series RGB data, LiDAR, or radar sequences. Usually, object tracking includes single-object tracking or multi-object tracking (MOT).
Multi-Object-Tracking Accuracy (MOTA) is a widely utilized metric for multiple object tracking, which combines false negatives, false positives, and mismatch rate [56] (see Eq. 9).
$MOTA=1-\frac{\sum_{t}\left(fp_{t}+fn_{t}+e_{t}\right)}{\sum_{t}gt_{t}}$ (9)
where $fp$, $fn$, and $e$ are the number of false positives, false negatives, and mismatch errors over time $t$. $gt$ is the ground truth.
Furthermore, instead of considering a single threshold, Average MOTA (AMOTA) is calculated based on all object confidence thresholds [57].
HD Map. HD mapping aims to construct detailed, highly accurate representations that include information about road structures, traffic signs, and landmarks. A dataset should provide LiDAR data for precise spatial information and camera data for visual details to ensure established map accuracy. According to [28], HD map automation [58] and HD map change detection [59] have received more and more attention. Normally, the HD map quality is estimated using the accuracy metric.
SLAM. Simultaneous Localization And Mapping (SLAM) entails building a concurrent map of the surrounding environment and localizing the vehicle within this map. Hence, data from cameras, IMUs for position tracking, and real-time LiDAR point clouds are vital. [60] introduces two evaluation metrics, relative pose error (RPE) and absolute trajectory error (ATE), for evaluating the quality of the estimated trajectory from the input RGB-D images.
IV-B Prediction
Prediction refers to forecasting the future states or actions of surrounding agents. This capacity ensures safer navigation in dynamic environments. Several evaluation metrics are used for prediction [61, 62], such as Root Mean Squared Error (RMSE):
$RMSE=\sqrt{\frac{1}{N}\sum_{n=1}^{N}\left(T_{pred}^{n}-T_{gt}^{n}\right)^{2}}$ (10)
where $N$ is the total number of samples, $T_{pred}$ and $T_{gt}$ represent the predicted trajectory and ground truth.
Negative Log Likelihood (NLL) (see Eq. 11) is another metric focusing on determining the correctness of the trajectory, which can be used to compare the uncertainty of different models [63].
$NLL=-\sum_{c=1}^{C}n_{c}log\left(\hat{n}_{c}\right)$ (11)
Where $C$ is the total classes, $n_{c}$ is the binary indicator of the correctness of prediction, and $\hat{n}_{c}$ is the corresponding prediction probability.
Trajectory Prediction. Using time-series data from sensors like cameras and LiDAR, trajectory prediction pertains to anticipating the future paths or movement patterns of other entities [61], such as pedestrians, cyclists, or other vehicles.
Behavior Prediction. Behavior prediction anticipates the potential actions of other road users [62], e.g., whether a vehicle will change lanes. Training behavior prediction models rely on extensive annotated data due to entities’ vast range of potential actions within various scenarios.
Intention Prediction. Intention prediction focuses on inferring the higher-level goals of the intention behind the actions of objects, involving a deeper semantic comprehension of the physical or mental activities of humans [64]. Because of the task’s complexity, it requires data not only from perception sensors like cameras but also additional information, such as traffic signals and hand gestures, to infer the intentions of other agents.
IV-C Planning and Control
IV-C1 Planning
Planning represents the decision-making process in reaction to the perceived environment and predictions. A classic three-level hierarchical planning framework consists of path, behavioral, and motion planning [53].
Path Planning Path planning, also known as route planning, involves setting long-term objectives. It is a high-level process of determining the best path to the destination.
Behavior Planning. Behavior planning sits at the mid-level of the framework and is related to decision-making, including lane changes, overtaking, merging, and intersection crossing. This process relies on the correct understanding and interaction with the behavior of other agents.
Motion Planning. Motion planning deals with the actual trajectory the vehicle should follow in real time, considering obstacles, road conditions, and the predicted behavior of other road agents. In contrast to path planning, motion planning generates appropriate paths to achieve local objectives [53].
IV-C2 Control
Control mechanisms in autonomous driving govern how the self-driving car executes the decided path or behavior from the motion planning system and corrects tracking errors [65]. It translates high-level commands into actionable throttle, brake, and steering commands.
IV-D End-to-end Autonomous Driving
End-to-end approaches in autonomous driving are those where a single deep learning model handles everything from perception to control, bypassing the traditional modular pipeline. Such models can often be more adaptive as they rely on adjusting the whole model through learning. Their inherent promise is in their simplicity and efficiency by reducing the need for hand-crafted components [66]. However, implementing end-to-end models faces critical limitations such as extensive training data requirements, low interpretability, and inflexible modular tuning.
Large-scale benchmarking for end-to-end AD can be categorized into closed-loop and open-loop evaluation [67]. The closed-loop evaluation is based on a simulated environment, while open-loop evaluation involves assessing a system’s performance against expert driving behavior from real-world datasets.
V High-Influence Datasets
This section describes the milestone autonomous driving datasets in the field of perception (V-A), prediction, planning, and control (V-B). We also exhibit datasets for end-to-end autonomous driving (V-C).
Figure 10: Chronological overview. We show the top 50 datasets with the highest impact scores. This diagram records datasets from 2007 to 2023 (until the writing of this paper).
Dataset Year Size Temp Tasks Categories Weather Time of day Scenario Geometrical Impact
2D Det 3D Det 2D Seg 3D Seg Tracking Lane Det number conditions type scope score
Onboard
nuScenes [22] 2019 40K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23 3 2 4 2 81.20
Cityscapes [14] 2016 25K $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 30 1 1 3 1 79.00
BDD100K [24] 2020 12M $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 40 5 2 4 1 77.82
Waymo [23] 2019 230K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23 2 3 5 1 77.81
KITTI [13] 2012 41K $\checkmark$ $\checkmark$ $\checkmark$ 8 1 1 2 1 77.30
SYNTHIA [2] 2016 13.4K $\times$ $\checkmark$ 13 2 3 5 - 70.44
Apolloscape [17] 2018 143,906 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 28 4 3 3 1 68.97
Virtual KITTI [3] 2016 21,260 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 8 3 2 5 - 67.61
VIPER [5] 2017 254,064 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 11 4 3 4 - 67.07
SemanticKITTI [6] 2019 43,552 $\checkmark$ $\checkmark$ 28 1 1 4 1 66.66
GTA5 [4] 2016 24,966 $\times$ $\checkmark$ 19 2 2 2 - 66.10
Argoverse 2 [28] 2023 6M $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 30 2 1 1 6 61.11
Lane Det [68] 2017 133,235 $\times$ $\checkmark$ 1 1 3 3 1 59.62
CityPersons [69] 2017 25K $\times$ $\checkmark$ $\checkmark$ 30 2 1 - 27 58.84
CamVid [70] 2009 701 $\times$ $\checkmark$ 32 2 1 2 1 58.69
IDD [71] 2019 10,004 $\times$ $\checkmark$ 34 3 3 5 1 58.12
Foggy Cityscapes [72] 2018 20,550 $\times$ $\checkmark$ $\checkmark$ 19 1 1 2 1 58.09
A2D2 [25] 2020 41,277 $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 38 2 1 3 3 58.06
Virtual KITTI 2 [9] 2020 20,992 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 8 4 2 3 1 56.74
RADIATE [26] 2021 44,140 $\checkmark$ $\checkmark$ $\checkmark$ 8 5 2 4 1 56.70
GTSDB [73] 2013 900 $\times$ $\checkmark$ 4 2 2 3 1 56.68
Caltech Ped [74] 2009 250K $\times$ $\checkmark$ 1 1 1 1 2 56.64
SHIFT [10] 2022 2,5M $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23 5 5 3 - 55.79
CAOS [7] 2022 13K $\times$ $\checkmark$ 13 3 2 3 - 55.74
STF [75] 2020 13,500 $\times$ $\checkmark$ $\checkmark$ 1 4 2 3 4 54.91
KITTI-360 [52] 2021 150K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 37 1 1 1 1 54.77
ACDC [76] 2021 4,006 $\times$ $\checkmark$ 19 4 2 3 1 54.62
VPGNet [77] 2017 20K $\times$ $\checkmark$ 17 3 2 1 1 54.37
ONCE [29] 2021 1M $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ 5 3 2 5 1 54.09
V2X
DAIR-V2X [27] 2021 71,254 $\checkmark$ $\checkmark$ $\checkmark$ 10 2 2 2 1 49.21
V2XSet [11] 2022 11,447 $\checkmark$ $\checkmark$ $\checkmark$ 1 1 1 1 - 45.39
V2VNet [78] 2020 51,2K $\checkmark$ $\checkmark$ 1 1 1 1 - 44.94
T&J [79] 2019 100 $\checkmark$ $\checkmark$ $\checkmark$ 1 1 1 2 1 44.58
Rope3D [80] 2022 50K $\times$ $\checkmark$ $\checkmark$ 12 3 3 2 1 44.39
V2X-Sim [81] 2022 10K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23 1 1 1 - 41.00
V2V4Real [82] 2023 40K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 5 2 1 2 1 40.28
Co-Percep [83] 2020 10K $\times$ $\checkmark$ $\checkmark$ 1 1 1 1 1 40.06
A9-Dataset [84] 2022 1,098 $\checkmark$ $\checkmark$ $\checkmark$ 9 4 2 1 1 39.12
LUMPI [85] 2022 200K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 3 6 3 1 1 35.87
DOLPHINS [86] 2022 42,376 $\checkmark$ $\checkmark$ $\checkmark$ 2 3 2 4 - 35.21
Drone
UAVDT [87] 2018 80K $\checkmark$ $\checkmark$ $\checkmark$ 3 2 2 6 1 61.63
DroneVehicle [88] 2021 28,439 $\times$ $\checkmark$ 5 1 3 4 1 44.42
Others
Mapillary Vistas [89] 2017 25K $\times$ $\checkmark$ 66 5 3 3 2 68.63
TT 100K [15] 2016 100K $\times$ $\checkmark$ 45 2 2 2 10 61.99
Pascal3D+ [90] 2014 30,899 $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 12 2 2 1 - 58.07
WildDash [91] 2018 1,800 $\times$ $\checkmark$ 28 2 2 7 1 50.02
TorontoCity [92] 2016 56K $\times$ $\checkmark$ 4 2 2 1 1 45.82
DAWN [93] 2020 4,543 $\times$ $\checkmark$ 5 6 3 3 - 43.99
RAD [94] 2019 60 $\times$ $\checkmark$ 19 1 1 1 1 37.86
STCrowd [95] 2022 10,891 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 1 3 1 1 1 35.19
TABLE II: High-impact perception datasets. For a more comprehensive demonstration, we exhibit 50 perception datasets from different sensing domains instead of those with the highest scores.
V-A Perception Datasets
Perception datasets are critical for developing and optimizing autonomous driving systems. They enhance vehicle reliability and robustness by providing rich, multimodal sensory data, ensuring effective perception and understanding of surroundings.
We leverage the proposed dataset evaluation metrics (see II-B) to calculate the impact scores of the collected perception datasets, subsequently selecting the top 50 datasets based on these scores to create a chronological overview, shown in Fig. 10. Concurrently, as described in III-B, we categorize datasets into onboard, V2X, drone, and others, choosing a subset from each to compile a comprehensive table of 50 datasets (Tab. II). It is worth noting that the datasets in the table are sorted by impact score within their respective categories and do not represent the overall top 50. In the following part, we choose several datasets with the highest impact scores within each sensing source and exhibit them considering their published year.
V-A1 Onboard
KITTI. KITTI [13] has profoundly influenced the autonomous driving domain since its release in 2012. KITTI encompasses a variety of real-world driving scenarios recorded by various sensors, including cameras, LiDAR, and GPS/IMU. Its rich annotations and high-resolution sensor data have facilitated the development and benchmarking of numerous algorithms across a range of autonomous driving tasks, such as object detection, tracking, optical flow, depth estimation, and visual odometry.
Cityscapes. Cityscapes [14] comprises a vast collection of images captured explicitly in intricate urban environments. Through meticulous labeling, Cityscapes provides pixel-level segmentation for 30 different object classes, which include various vehicle types, pedestrians, roads, and traffic sign information. Given its complexity and richness, Cityscapes has become a standard benchmark for tasks such as semantic segmentation in urban scenarios.
SYNTHIA. SYNTHIA [2] is a synthetic dataset in the realm of autonomous driving. The dataset includes 13,400 images with pixel-wise annotations for semantic segmentation. One of the standout attributes of SYNTHIA is its capability to bridge the gap between real-world and synthetic data, fostering the development of robust and transferable approaches across different domains.
Virtual KITTI. Virtual KITTI [3] crafts to closely mimic the original KITTI dataset through a virtual environment, distinguishing itself by delivering high-resolution video sequences encapsulating various traffic situations and environmental conditions. Similar to [13], Virtual KITTI supports crucial autonomous driving tasks, including object detection, semantic segmentation, and object tracking.
VIPER. VIPER [5] is a synthetic dataset collected from driving, riding, and walking perspectives in a realistic virtual world, addressing the challenge of data scarcity and the high cost of labeling real-world data. VIPER contains over 250,000 video frames annotated with ground-truth data for low- and high-level vision tasks while encapsulating various weather conditions, lighting scenarios, and complex urban landscapes. Overall, VIPER presents a valuable and cost-effective tool for researchers to expedite the development of reliable and safe autonomous driving.
Apolloscapes. Apolloscapes [17] offers more than 140,000 high-resolution frames with accurate bounding boxes and pixel-wise semantic labels, critical for training and validating autonomous vehicles’ perception and navigation systems. Apolloscapes supports images and point clouds semantic segmentation, 2D/3D object detection, multi-object tracking, and lane segmentation, enabling the creation and evaluation of advanced and safe autonomous driving systems.
SemanticKITTI. SemanticKITTI [6] is a notable extension of the KITTI family, focusing on semantic segmentation in the autonomous driving field. SemanticKITTI consists of over 43,000 LiDAR point cloud frames, making it one of the largest datasets for 3D semantic segmentation in outdoor environments. SemanticKITTI provides precise labels for 28 categories, such as car, road, building, etc., achieving a robust benchmark for evaluating the performance of point cloud semantic segmentation methods, underpinning numerous studies and innovations in related domains.
nuScenes. nuScenes [22] stands as an essential contribution to the field of autonomous driving, offering a rich repository of data that addresses the diverse needs of perception systems. nuScenes leverages LiDAR, radars, and cameras to record data from different urban scenes from Boston and Singapore. It is worth mentioning that its six cameras provide a comprehensive perspective of the surrounding environment, making them widely utilized in multi-view object detection tasks. In conclusion, the nuScenes dataset is a cornerstone in developing autonomous driving technologies, supporting multi-tasks and applications, and setting new benchmarks in the field.
Waymo. The Waymo Open Dataset [23], introduced in 2019, significantly influences research and advancement in autonomous driving by providing an extensive size of multimodal sensory data with high-quality annotations. Key contributions of the Waymo dataset include its comprehensive coverage of driving conditions and geographics, which are pivotal for the robustness and generability of different tasks, such as detection, tracking, and segmentation.
BDD100K. BDD100K [24] dataset, released by the Berkeley DeepDrive center in 2018, is a substantial and diverse driving dataset renowned for its size and diversity. It comprises 100,000 videos, each about 40 seconds in duration. Meanwhile, it offers various annotated labels for object detection, tracking, semantic segmentation, and lane detection. This extensive compilation of data prompts advancements in the autonomous driving community, establishing itself as a challenging and versatile platform for researchers and engineers to propose and refine algorithms.
RADIATE. RADIATE [26] is the first public radar dataset, which contains 44,140 annotated frames gathered under several adversarial weather conditions, such as rain, fog, overcast, and snow. It also incorporates LiDAR and camera data, allowing full perception and understanding of the driving surroundings.
Argoverse 2. As a sequel to Argoverse 1 [19], Argoverse 2 [28] introduces more diversified and complex driving scenarios, presenting the largest autonomous driving taxonomy to date. It captures various real-world driving scenarios across six cities and varying conditions. Argoverse 2 supports a wide range of essential tasks, including but not limited to 3D object detection, semantic segmentation, and tracking. In summary, the Argoverse 2 dataset offers numerous multimodal data of real-world driving scenarios, fostering innovations and advancements of algorithms and showing its substantial potential as a vital resource in autonomous driving.
V-A2 V2X
V2VNet. The dataset introduced by V2VNet [78] focuses on leveraging V2V communication to allow autonomous vehicles to share information and perceive the environment from multiple viewpoints, which is crucial for detecting occluded objects and predicting the behavior of other traffic participants. This dataset was created using a high-fidelity LiDAR simulator named Lidarsim [96] that utilizes real-world data to generate realistic LiDAR point clouds for various traffic scenes. In conclusion, this work brings attention to the V2V as a promising avenue for improving the capabilities of autonomous vehicles.
DAIR-V2X. DAIR-V2X [27] is a pioneering resource in the Vehicle-Infrastructure Cooperative Autonomous Driving, providing large-scale, multi-modality, multi-view real-world data. The dataset is designed to tackle challenges such as the temporal asynchrony between vehicle and infrastructure sensors and the data transmission costs involved in such cooperative systems. The impact of the DAIR-V2X dataset on autonomous driving is significant because it sets a benchmark for complexities of vehicle-infrastructure cooperation thanks to its multiple scenarios from the real world.
Rope3D. Rope3D presented in [80] is an essential contribution to the perception systems, which addresses the critical gap in autonomous driving by leveraging data gathered from roadside cameras. Rope3D consists of 50,000 images under various environmental conditions, including different illumination (daytime, night, dusk) and weather scenarios (rainy, sunny, cloudy), ensuring a high diversity of data that reflects real-world complexities. Overall, the Rope3D dataset is a pioneering work that accelerates advancements in roadside perception for autonomous driving while being a critical tool for researchers and engineers to develop more robust and intelligent autonomous driving systems.
V2V4Real. V2V4Real [82] is the first large-scale, real-world dataset to address V2V cooperative perception. The data was collected from two vehicles with multimodal sensors like LiDAR and cameras. V2V4Real focuses on a range of perception tasks, such as cooperation 3D object detection, cooperative 3D object tracking, and Sim2Real domain adaptation. This versatility makes it an invaluable resource for developing and benchmarking autonomous driving algorithms.
V-A3 Drone
UAVDT. The UAVDT [87] dataset consists of 80,000 accurately annotated frames with up to 14 kinds of attributes, such as weather conditions, flying attitude, camera view, vehicle category, and occlusion levels. The dataset focuses on UAV-based object detection and tracking in urban environments. Moreover, the UAVDT benchmark includes high-density scenes with small objects and significant camera motion, all challenging for the current state-of-the-art methods.
DroneVehicle. DroneVehicle [88] proposes a large-scale drone-based dataset, which provides 28,439 RGB-Infrared image pairs to address object detection, especially under low-illumination conditions. Furthermore, it covers a variety of scenarios, such as urban roads, residential areas, and parking lots. This dataset is a significant step forward in developing autonomous driving technologies due to its unique drone perspective across a broad range of conditions.
V-A4 Others
Pascal3D+. Pasacal3D+ [90] is an extension of the PASCAL VOC 2022 [97], overcoming the limitations of previous datasets by providing a richer and more varied set of annotations for images. Pasacal3D+ augments 12 rigid object categories, such as cars, buses, and bicycles, with 3D pose annotations and adds more images from ImageNet [98], resulting in a high degree of variability.
TT 100K. Tsinghua-Tencent 100K [15] solves the challenges of detecting and classifying traffic signs in realistic driving conditions. It provides 100K images, including 30,000 traffic-sign instances. Beyond its large data size, the high-resolution images encompass a wide range of illumination and weather conditions, making it robust for training and validating traffic sign recognition.
Mapillary Vistas. Mapillary Vistas dataset, proposed by [89] in 2017, particularly aims at semantic segmentation of street scenes. The 25,000 images in the dataset are labeled with 66 object categories and include instance-specific annotations for 37 classes. It contains images from diverse weather, time of day, and geometric locations, which helps mitigate the bias towards specific regions or conditions.
V-B Prediction, Planning, and Control Datasets
Prediction, planning, and control datasets serve as the foundation for facilitating the training and evaluation of driving systems to forecast traffic dynamics, pedestrian movements, and other essential factors that influence driving decisions. By simulating myriad driving scenarios, they empower autonomous vehicles to make informed decisions, navigate complex environments, and maintain safety and efficiency on the road. Hence, we demonstrate in detail several high-impact datasets related to these tasks according to the data size, modalities, and citation number. We summarize the prediction, planning, and control datasets into task-specific and multi-task groups.
Dataset Year Sensing domain Size Tasks Weather conditions Time of day Scenario conditions
Task-Specific
JAAD [16] 2017 onboard 75K frames pedestrian IT sunny, rainy, cloudy, snowy day, afternoon, night urban
Dr(eye)ve [53] 2018 onboard 500K frames driver’s attention prediction sunny, rainy, cloudy day, night urban, countryside, highway
highD [18] 2018 drone 45K km distance TP sunny 8 am to 5 pm highway
PIE [21] 2019 onboard 293K frames pedestrian IT sunny, overcast day urban
USyd [99] 2019 onboard 24K trajectories driver IT 5 intersections
Argoverse [19] 2019 onboard 300K trajectories TP variety variety urban
inD [100] 2020 drone 11.5K trajectories road user prediction sunny day 4 urban intersections
PePscenes [101] 2020 onboard 719 frames pedestrian BP
openDD [102] 2020 drone 84,774 trajectories pedestrian BP 7 roundabouts
nuPlan [103] 2021 drone 1.5K hours data MPlan 4 cities
exiD [104] 2022 drone 16 hours data TP sunny daytime 7 locations on highway
MONA [105] 2022 drone 702K trajectories TP sunny, overcast, rain 8 am to 5 pm urban
Multi-Task
INTERACTION [20] 2019 drone, V2X 110K trajectories MPlan and MP, DM (un)signalized intersection
BLVD [106] 2019 onboard 120K frames 4D OT, 5D event recognition day, night urban, highway
rounD [107] 2019 drone 13,746 road users scenario classification, BP sunny daytime (sub-)urban
Lyft Level 5 [108] 2021 drone 1.1K hours data MPlan, MP suburban
LOKI [109] 2021 onboard 644 scenarios TP, BP (sub-)urban
SceNDD [110] 2022 onboard 68 driving scenes MPlan, MP urban
DeepAccident [12] 2023 V2X 57K frames MP, accident prediction sunny, rainy, cloudy, wet noon, sunset, night synthetic
Talk2BEV [111] 2023 onboard 20K QA pairs DM, IT, spatial reasoning
V2X-Seq (forecasting) [50] 2023 V2X 50K scenarios online/offline VIC TP 28 urban intersections
TABLE III: Prediction, planning, and control datasets. We demonstrate several crucial datasets related to prediction, planning, and control. BP: behavior prediction, IP: intention prediction, MP: motion prediction, TP: trajectory prediction, MPlan: motion planning, DM: decision-making, OT: object tracking, QA: question-answering
V-B1 Task-Specific Datasets
highD. The drone-based highD [18] dataset provides a large-scale collection of naturalistic vehicle trajectories on German highways, containing post-processed trajectories of 110,000 cars and trucks. The dataset aims to overcome the limitations of existing measurement methods for scenario-based safety validation, which often fail to capture the naturalistic behavior of road users or to include all relevant data with sufficient quality.
PIE. The Pedestrian Intention Estimation (PIE) dataset proposed by [21] represents a significant advancement in understanding pedestrian behaviors in urban environments. It encompasses over 6 hours of driving footage recorded in downtown Toronto under various lighting conditions. The PIE dataset offers rich annotations for perception and visual reasoning, including bounding boxes with occlusion flags, crossing intention confidence, and text labels for pedestrian actions. The long continuous sequences and annotations facilitate multiple tasks like trajectory prediction and pedestrian intention prediction.
USyd. USyd [99] pushes the progress of the driver intention prediction in the context of urban intersections without traffic signals, which are common in urban settings and represent a challenge due to the lack of clear road rules and signals. The dataset incorporates over 23,000 vehicles traversing five different intersections, collected using a vehicle-mounted LiDAR-based tracking system. The data modalities included are exhaustive, providing vehicle tracks with lateral and longitudinal coordinates, heading, and velocity. This information is vital to predict driver behavior, accounting for the uncertainty inherent in human driving patterns.
Argoverse. Argoverse [19] is a crucial dataset in 3D object tracking and motion forecasting. Argoverse provides ${}^{\circ}$ images from 7 cameras, forward-facing stereo imagery, and LiDAR point clouds. The recorded data covers over 300K extracted vehicle trajectories from 290km of mapped lanes. With the assistance of rich sensor data and semantic maps, Argoverse is pivotal in advancing research and development in prediction systems.
inD. The importance of the inD [100] lies in its large-scale, high-quality, and diverse set of trajectory data crucial for several applications, including road user prediction models and scenarios-based safety validation for autonomous vehicles in urban intersection environments. It covers around 11,500 different road user trajectories, e.g., vehicles, bicyclists, and pedestrians. These trajectories have a positioning error of less than 0.1 meters, which is pivotal for the reliability of the data.
PePscenes. PePscenes [101] tackles the requirement for understanding and anticipating pedestrian actions in dynamic driving surroundings. This dataset enhances the nuScenes [22] dataset by adding per-frame 2D/3D bounding box and behavior annotations, focusing on pedestrian crossing actions. One of the key attributes of [101] is incorporating various data types involving semantic maps, scene images, trajectories, and ego-vehicle states, which are necessary for creating robust models capable of understanding complex traffic scenarios.
openDD. The openDD [102] dataset focuses on analyzing and predicting traffic scenes around roundabouts, which are complex and unregulated by traffic lights. It is unique in leveraging drone-captured imagery with high resolution (4K), which spans over 62 hours of trajectory data from 501 separate flights. The dataset contains not only trajectories but also shapefiles and an extensible markup language (XML) file describing the road topology, along with a reference image for each underlying intersection.
nuPlan. nuPlan [103] is the world’s first closed-loop machine learning-based planning benchmark in autonomous driving. This multimodal dataset comprises around 1,500 hours of human driving data from four cities across America and Asia, featuring different traffic patterns, such as merges, lane changes, interactions with cyclists and pedestrians, and driving in construction zones. These characters of the nuPlan dataset take into account the dynamic and interactive nature of actual driving, allowing for a more realistic evaluation.
exiD. The exiD [104] trajectory dataset, presented in 2022, is a pivotal contribution to the highly interactive highway scenarios. It takes advantage of drones to record traffic without occlusion, minimizing the influence on traffic and ensuring high data quality and efficiency. This drone-based dataset surpasses previous datasets in terms of the diversity of interactions captured, especially those involving lane changes at highway entries and exits.
MONA. The Munich Motion Dataset of Natural Driving (MONA) [105] is an extensive dataset, with 702K trajectories from 130 hours of videos, covering urban roads with multiple lanes, an inner-city highway stretch, and their transitions. This dataset boasts an average overall position accuracy of 0.51 meters, which exhibits the quality of the data collected using highly accurate localization and LiDAR sensors.
V-B2 Multi-Task Datasets
INTERACTION. The INTERACTION [20] dataset covers diverse, complex, and critical driving scenes, coupled with its comprehensive semantic map, allowing it to be a versatile platform for a multitude of tasks, such as motion prediction, imitation learning, and validation of decision and planning. Its inclusion of different countries further improves the robustness of analyzing the driving behavior across different cultures, which is critical for the global development of autonomous driving.
BLVD. The BLVD [106] benchmark facilitates tasks such as dynamic 4D (3D+temporal) tracking, 5D (4D+interactive) interactive event recognition, and intention prediction, which are essential for a deeper understanding of traffic scenes. BLVD offers around 120K frames from different traffic scenes, comprising object density (low and high) and illumination conditions (daytime and nighttime). These frames are fully annotated with a large number of 3D labels incorporating vehicles, pedestrians, and riders.
rounD. The rounD dataset presented by [107] is pivotal for scenario classification, road user behavior prediction, and driver modeling because of the large number of collections of road user trajectories at roundabouts. The dataset utilizes a drone equipped with a 4K resolution camera to collect over six hours of video, recording more than 13K road users. The broad recorded traffic situations and the high-quality recordings make rounD an essential dataset in autonomous driving, facilitating the study of naturalistic driving behaviors in public traffic.
Lyft Level 5. Lyft Level 5 [108] represents one of the most extensive autonomous driving datasets for motion prediction to date, with over 1,000 hours of data. It encompasses 17,000 25-second long scenes, a high-definition semantic map with over 15,000 human annotations, 8,500 lane segments, and a high-resolution aerial image of the area. It supports multiple tasks like motion forecasting, motion planning, and simulation. The numerous multimodal data with detailed annotations make the Lyft Level 5 dataset a vital benchmark for prediction and planning.
LOKI. LOKI [109], standing for Long Term and Key Intentions, is an essential dataset in multi-agent trajectory prediction and intention prediction. LOKI tackles a crucial gap in intelligent and safety-critical systems by proposing large-scale, diverse data for heterogeneous traffic agents, including pedestrians and vehicles. This dataset makes a multidimensional view of traffic scenarios available by utilizing camera images with corresponding LiDAR point clouds, making it a highly flexible resource for the community.
SceNDD. SceNDD [110] introduces real driving scenarios that feature diverse trajectories and driving behaviors for developing efficient motion planning and path-following algorithms. It is also adaptable to different configurations of the ego-vehicle and contains a time horizon of prediction that can be broken down into timestamps for detailed analysis. In conclusion, the SceNDD dataset is a significant addition to autonomous driving prediction and planning research.
DeepAccident. The synthetic dataset, DeepAccident [12], is the first work that provides direct and explainable safety evaluation metrics for autonomous vehicles. The extensive dataset with 57K annotated frames and 285K annotated samples supports end-to-end motion and accident prediction, which is vital for improving the predictive capabilities of autonomous driving systems in avoiding collisions and ensuring safety. Moreover, this multimodal dataset is versatile for various V2X-based perception tasks, such as 3D object detection, tracking, and bird’s-eye-view (BEV) semantic segmentation.
Talk2Bev. The innovative dataset, Talk2BEV [111], promotes the shift from traditional autonomous driving tasks to incorporate large vision-language models with BEV maps in the context of autonomous driving. Talk2BEV utilizes recent advances in vision-language models, allowing for a more flexible and comprehensive understanding of road scenarios. This dataset comprises over 20,000 diverse question categories, all human-annotated and derived from [22]. The proposed Talk2BEV-Bench benchmark can be leveraged across multiple tasks, covering decision-making, visual and spatial reasoning, and intent prediction.
V2X-Seq (Forecasting). The trajectory forecasting dataset is a substantial part of the real-world dataset V2X-Seq [50], containing about 80K infrastructure-view and 80K vehicle-view scenarios, and a further 50K cooperative-view scenarios. This diversity of sensing domains creates a more holistic view of the traffic environment, exhibiting huge potentiality for research and analysis on vehicle-infrastructure-cooperative (VIC) trajectory prediction.
V-C End-to-End Datasets
End-to-end has become a growing trend as an alternative to modular-based architecture in autonomous driving [66]. Several versatile datasets (like nuScenes [22] and Waymo [23]) or simulators like CARLA [112] provide the opportunity to develop end-to-end autonomous driving. Meanwhile, some works present datasets that are especially for end-to-end driving.
DDD17. The DDD17 [113] dataset is notable for its use of event-based cameras, which provide a concurrent stream of standard active pixel sensor (APS) images and dynamic vision sensors (DVS) temporal contrast events, offering a unique blend of visual data. Additionally, DDD17 captures diverse driving scenarios, including highway and city driving and various weather conditions, thus providing exhaustive and realistic data for training and testing end-to-end autonomous driving algorithms.
Other datasets summarized in this survey are shown in Tab. IV Tab. V, Tab. VI.
Dataset Year Size Temporal Sensing domain Tasks Real/Synthetic
KAIST MPD [114] 2015 95K color-thermal pair frames $\times$ onboard pedestrian detection real
BelgiumTS [115] 2011 13K traffic sign annotations $\times$ onboard traffic sign detection real
nighttime drive [116] 2018 35K $\times$ onboard SS real
[…]
Caltech Lanes [118] 2019 1,224 frames $\times$ onboard lane detection real
A*3D [119] 2020 39,179 point cloud frames $\times$ onboard (3D) OD real
PreSIL [8] 2019 50K frames $\checkmark$ onboard OD, 3D SS synthetic
H3D [120] 2019 27,721 frames $\checkmark$ onboard (3D) OD, MOT real
ROAD [121] 2021 122K frames $\checkmark$ onboard OD, SS real
ETH Ped [122] 2007 2,293 frames $\checkmark$ onboard pedestrian detection real
CADC [123] 2020 56K $\times$ onboard (3D) OD, OT real
RadarScenes [124] 2021 40K radar frames $\checkmark$ onboard OD, classification real
All-in-One Drive [125] 2021 100K $\checkmark$ onboard (3D) OD, (3D) SS, trajectory prediction real
CURE-TSR [126] 2017 2.2M annotated images $\times$ onboard traffic sign detection real
Paris-Lille-3D [127] 2018 2,479 frames $\times$ onboard 3D SS, classification real
PandaSet [128] 2021 8,240 frames $\checkmark$ onboard (3D) OD, SS, OT real
NightOwis [129] 2018 56K frames $\checkmark$ onboard pedestrian detection, tracking real
SynWoodScape [130] 2022 80K frames $\times$ onboard (3D) OD, segmentation synthetic
Boreas [131] 2023 7,111 frames $\checkmark$ onboard (3D) OD, localization real
TUD-Brussels [132] 2009 1,600 frames $\times$ onboard (3D) OD real
VEIS [133] 2018 61,305 frames $\times$ onboard OD, SS synthetic
CCTSDB 2021 [134] 2021 16,356 frames $\times$ onboard traffic sign detection real
SemanticPOSS [135] 2020 2,988 point cloud frames $\checkmark$ onboard 3D SS synthetic
IDDA [136] 2020 1M frames $\times$ onboard segmentation synthetic
TJ4RadSet [137] 2022 7,757 frames $\checkmark$ onboard OD, OT real
TME Motorway [138] 2012 30K frames $\checkmark$ onboard OD, OT real
Stanford Tack [139] 2011 14K tracks $\checkmark$ onboard classification real
CARRADA [140] 2020 7,193 radar frames $\checkmark$ onboard SS real
SODA10M [141] 2021 20K labeled images $\checkmark$ onboard OD real
Robo3D [142] 2023 476K frames $\checkmark$ onboard (3D) OD, 3D SS real
LostAndFound [143] 2016 2,104 frames $\checkmark$ onboard obstacle detection real
Titan [144] 2020 75,262 frames $\checkmark$ onboard OD, action recognition real
CODA [145] 2022 1,500 frames $\times$ onboard corner case detection real
PixSet [146] 2021 29K point cloud frames $\times$ onboard (3D) OD real
ZOD [30] 2023 100K frames $\checkmark$ onboard (3D) OD, segmentation real
K-Radar [147] 2023 35K radar frames $\times$ onboard 3D OD, OT real
RoadObstacle21 [148] 2021 327 scenes $\times$ onboard anomaly segmentation synthetic
VIL-100 [149] 2021 10K frames $\times$ onboard lane detection real
EuroCity Persons [150] 2018 47,300 frames $\times$ onboard OD real
TuSimple [151] 2017 6,408 frames $\times$ onboard lane detection, velocity estimation real
OpenMPD [152] 2021 15K frames $\times$ onboard (3D) OD, 3D OT, semantic segmentation real
WADS [153] 2021 1K point cloud frames $\checkmark$ onboard SS real
NightCity [154] 2020 4,297 frames $\times$ onboard nighttime SS real
LiDAR Snowfall [155] 2022 7,385 point cloud frames $\checkmark$ onboard 3D OD synthetic
LLMAS [156] 2019 100K frames $\checkmark$ onboard lane detection real
Fishyscapes [157] 2021 1,030 frames $\times$ onboard SS, anomaly detection real
Toronto-3D [158] 2020 4 scenarios $\times$ onboard 3D SS real
MIT-AVT [159] 2020 1.15M 10s video clips $\checkmark$ onboard SS, anomaly detection real
LISA [160] 2014 6,610 frames $\checkmark$ onboard traffic sign detection real
SynthCity [161] 2019 367.9M points in 30 scans $\times$ onboard (3D) OD, (3D) SS synthetic
SemanticUSL [162] 2021 1.2K frames $\times$ onboard domain adaptation 3D SS real
RANUS [163] 2018 4K frames $\times$ onboard SS, scene understanding real
PePScenes [101] 2020 40K frames $\checkmark$ onboard (3D) OD, pedestrian action prediction real
MUAD [164] 2022 10.4K frames $\checkmark$ onboard OD. SS, depth estimation synthetic
AUTOCASTSIM [165] 2022 52K frames $\checkmark$ V2X (3D) OD, OT, SS real
Mcity [166] 2019 1,7500 frames $\checkmark$ onboard SS real
comma2k19 [167] 2018 2M images $\times$ onboard pose estimation, end-to-end driving real
DET [168] 2019 5,424 event-based camera images $\times$ onboard lane detection real
TRoM [169] 2017 712 frames $\times$ onboard road marking detection real
Cirrus [170] 2020 6,285 frames $\checkmark$ onboard (3D) OD real
KITTI InstanceMotSeg [171] 2020 12,919 frames $\checkmark$ onboard moving instance segmentation real
CARTI [172] 2022 11K frames $\checkmark$ V2X cooperative perception synthetic
Multifog KITTI [173] 2021 15K frames $\times$ onboard 3D OD synthetic
UNDD [174] 2019 7.2K frames $\checkmark$ onboard SS real
CeyMo [175] 2021 2,887 frames $\times$ onboard road marking detection real
RaidaR [176] 2021 58,542 rainy street scenes $\times$ onboard SS real
K-Lane [177] 2022 15,382 frames $\times$ onboard lane detection real
aiMotive [178] 2023 26,583 frames $\checkmark$ onboard 3D OD, MOT real
SAP [179] 2016 19K frames $\times$ drone OD, OT real
Astyx [180] 2019 500 radar frames $\times$ onboard 3D OD real
Comap [181] 2021 4,391 frames $\checkmark$ V2X 3D OD synthetic
Ithaca365 [182] 2022 7K frames $\times$ onboard 3D OD, SS, depth estimation real
Small Obstacles [183] 2020 3K frames $\times$ onboard small obstacle segmentation real
GLARE [184] 2022 2,157 frames $\times$ onboard traffic sign detection real
WIBAM [185] 2021 33,092 frames $\checkmark$ V2X 3D OD real
MIT DriveSeg [186] 2019 5K frames $\times$ onboard SS real
SUPS [187] 2023 5K frames $\checkmark$ onboard SS, depth estimation, SLAM synthetic
R3 [188] 2021 369 scenes $\times$ onboard out-of-distribution detection real
TABLE IV: Autonomous driving dataset 1
Dataset Year Size Temporal Sensing domain Tasks Real/Synthetic
V2X-Seq (perception) [50] 2023 15K frames $\checkmark$ V2X cooperative perception real
DRIV100 [189] 2021 100K frames $\times$ onboard domain adaptation SS real
NEOLIX [190] 2021 30K frames $\checkmark$ onboard 3D OD, OT real
CARLANE [191] 2022 118K frames $\checkmark$ onboard lane detection synthetic
IPS3000+ [192] 2021 14,198 frames $\checkmark$ V2X 3D OD real
TUM-Traffic [193] 2023 4.8K frames $\checkmark$ V2X 3D OD, OT real
Amodal Cityscapes [194] 2022 5K frames $\times$ onboard amodal SS real
WEDGE [195] 2023 3,360 frames $\times$ others OD, classification synthetic
HEV [196] 2019 230 video clips $\checkmark$ onboard object localization real
Aachen Day-Night [197] 2018 4,328 images, 1.65M points $\checkmark$ onboard visual localization real
HAD [198] 2019 5,675 video clips $\checkmark$ onboard end-to-end driving real
CARLA-100 [199] 2019 100 hours driving $\checkmark$ onboard path planning, behavior cloning synthetic
Collective activity [200] 2009 44 short videos $\checkmark$ others human activity recognition real
DA4AD [201] 2020 9 sequences $\checkmark$ onboard visual localization real
Bosch STL [202] 2017 13,334 images $\checkmark$ onboard traffic light detection and classification real
UrbanLaneGraph [203] 2023 around 5,220 km lane spans $\checkmark$ drone lane graph estimation real
CrashD [204] 2022 15,340 scenes $\times$ onboard 3D OD synthetic
CODD [205] 2022 108 sequences $\checkmark$ V2X multi-agent SLAM synthetic
MIT DSS [186] 2020 10K video frames $\checkmark$ onboard SS real
CPIS [83] 2020 10K frames $\times$ V2X cooperative 3D OD synthetic
CarlaScenes [206] 2022 7 sequences $\checkmark$ onboard (3D) SS, SLAM, depth estimation synthetic
FlyingThings3D [207] 2015 26,066 frames $\checkmark$ others scene flow estimation synthetic
San Francisco Landmark [208] 2011 150K panoramic images $\times$ others landmark identification real
DDAD [209] 2020 21,200 frames $\times$ onboard depth estimation real
Ua-detrac [210] 2015 140K frames $\checkmark$ V2X OD, MOT real
NCLT [211] 2015 34.9 hours $\checkmark$ onboard odometry real
NVSEC [212] 2018 around 28 km distances $\checkmark$ others SLAM real
MSLU [213] 2013 36.8 km distances $\checkmark$ onboard SLAM real
Oxford Radar RobotCar [214] 2020 240K scans $\checkmark$ onboard odometry real
A2D2 [25] 2020 433,833 frames $\checkmark$ onboard 3D OD, SS, SLAM real
VERI-Wild [215] 2019 416,314 images $\checkmark$ V2X onboard re-identification real
OTOH [108] 2020 170K scenes $\checkmark$ drone trajectory prediction, planning real
TRANCOS [216] 2015 1.2K images $\times$ V2X onboard number estimation real
Complex Urban [217] 2017 around 190km paths onboard SLAM real
Syncapes [218] 2018 25K images $\times$ onboard OD, SS synthetic
SydneyUrbanObject [219] 2013 588 object scans $\times$ onboard classification real
ApolloCar3D [220] 2019 5,277 driving images $\times$ onboard 3D instance understanding real
ACDC [76] 2021 4,006 images $\times$ onboard SS on adverse conditions real
MulRan [221] 2020 41.2km paths $\checkmark$ onboard place recognition real
Paris-rue-Madame [222] 2014 643 objects $\times$ onboard OD, SS real
Ground Truth SitXel [223] 2013 78,500 frames $\checkmark$ onboard stereo confidence real
LiVi-Set [224] 2018 10K frames $\checkmark$ onboard driving behavior prediction real
Newer College [225] 2020 290M points, 2300 seconds $\checkmark$ others SLAM real
CADP [226] 2018 1,416 scenes $\times$ V2X traffic accident analysis real
LIBRE [227] 2020 40 frames $\times$ onboard LiDAR performance benchmark real
OPV2V [228] 2022 11,464 frames $\times$ V2X onboard-to-onboard perception synthetic
NYC3DCars [229] 2013 2K images $\times$ onboard OD real
RUGD [230] 2019 37K frames $\checkmark$ others SS real
EU LTD [231] 2020 around 37 hours $\checkmark$ onboard odometry real
MOTSynth [232] 2021 768 driving sequences $\checkmark$ onboard Pedestrian detection and tracking synthetic
LLamas [156] 2019 100,042 images $\times$ onboard lane marker detection, lane segmentation real
PedX [233] 2018 5K images $\checkmark$ onboard pedestrian detection and tracking real
CCD [234] 2020 4.5K videos $\checkmark$ onboard accident prediction real
MAVD [235] 2021 113,283 images $\checkmark$ onboard OD and OT with sound real
Gated2Depth [236] 2020 17,686 frames $\checkmark$ onboard depth estimation real
DDD20 [237] 2017 51 hours event frames $\checkmark$ onboard end-to-end driving real
TCGR [238] 2020 839,350 frames $\checkmark$ others traffic control gesture recognition real
 4Seasons [238] 2020 350km recordings $\checkmark$ onboard SLAM real
CCSAD [239] 2015 96K frames $\checkmark$ onboard scene understanding real
TAF-BW [240] 2018 2 scenarios $\checkmark$ V2X MOT, V2X communication real
Boxy [241] 2019 200K frames $\checkmark$ onboard OD real
AMUSE [242] 2013 117,440 frames $\checkmark$ onboard SLAM real
Brno Urban [243] 2019 375.7 km $\checkmark$ onboard recognition real
AUTOMATUM [244] 2021 30 hours $\checkmark$ drone trajectory prediction real
DurLAR [245] 2021 100K frames $\checkmark$ onboard depth estimation real
Reasonable-Crowd [246] 2021 92 scenarios $\checkmark$ onboard driving behavior prediction synthetic
Daimler Ped [247] 2013 12,485 frames $\checkmark$ onboard pedestrian path prediction real
PVDN [248] 2021 59,746 frames $\checkmark$ onboard nighttime OD, OT real
CARLA-WildLife [249] 2022 26 videos $\checkmark$ onboard out-of-distribution tracking synthetic
SOS [249] 2022 20 videos $\checkmark$ onboard out-of-distribution tracking real
RoadSaW [250] 2022 720K scenes $\checkmark$ onboard Road surface and wetness estimation real
I see you [251] 2022 170 sequences, 340 trajectories $\checkmark$ V2X OD real
ASAP [252] 2022 1.2M images $\checkmark$ onboard online 3D OD real
OpenLane-V2 [253] 2023 466K images $\checkmark$ onboard lane detection, scene understanding real
Daimler Stereo Ped [254] 2011 21,790 frames $\checkmark$ onboard pedestrian detection real
Brain4Cars [255] 2015 2M frames $\checkmark$ onboard Maneuver Anticipation real
NEXET [256] 2017 91,190 frames $\times$ onboard OD real
DIML [257] 2017 470 videos $\checkmark$ onboard lane detection real
TABLE V: Autonomous driving dataset 2
Dataset Year Size Temporal Sensing domain Tasks Real/Synthetic
UAH-Driveset [258] 2016 500 mins $\checkmark$ onboard lane detection, detection real
SSCBENCH [259] 2023 66,913 frames $\times$ onboard semantic scene completion real
FLIR [260] - 26,442 thermal frames $\checkmark$ onboard thermal image OD real
VLMV [261] 2020 900 frames $\checkmark$ V2X lane merge real
CityFlow [262] 2019 200K bounding boxes $\checkmark$ V2X OD, MOT, re-identification real
PREVENTION [263] 2019 356 mins $\checkmark$ onboard onboard trajectory and intention prediction real
TABLE VI: Autonomous driving dataset 3
VI Annotations process
The success and reliability of AD algorithms rely not only on the numerous data but also on high-quality annotations. In this section, we first explain the methodology for annotating data VI-A. Additionally, we analyze the most important aspects for ensuring annotation quality VI-B.
VI-A How annotations are created
Different AD tasks require specific types of annotation. For example, object detection requires bounding box labels of instance, segmentation is based on pixel- or point-level annotations, and continually labeled trajectory is critical for trajectory prediction. On the other hand, as shown in Fig. 11, the annotation pipeline can be categorized into three types: manual annotation, semi-automatic annotation, and fully automatic annotation. We detail the labeling approaches for different types of annotation in this section.
Annotate Segmentation Data. The target of annotating segmentation data is to assign a label to each pixel in an image or each point in a LiDAR frame to indicate which object or region it belongs to. After labeling, all pixels belonging to an object are annotated with the same class. For the manual annotation process, the annotator first draws boundaries around an object and then fills in the area or paints over the pixels directly. However, generating pixel/point-level annotations in this way is costly and inefficient.
Figure 11: Annotation pipelines. We demonstrate (a) Manual annotation: The professional annotators label the raw data using annotation tools. (b) Semi-automatic annotation: After generating annotations using an automatic annotation algorithm, the low-quality labels are refined by annotators. (c) Fully automatic annotation: The framework annotates data without human correction.
Many studies have proposed fully or semi-automatic annotation methods to improve annotation efficiency. [264] presented a fully automatic annotation approach based on weakly supervised learning to segment proposed drivable paths in images. [265] is a semi-automatic annotation method utilizing the objectness priors to generate segmentation masks. After that, [266] offered a semi-automatic method considering 20 classes. Polygon-RNN++[267] presented an interactive segmentation annotation tool following the idea of [268]. Instead of using image information to generate pixel-level labels, [269] explored transferring 3D information into 2D image domains to generate semantic segmentation annotations. For labeling 3D data, [270] proposed an image-assisted annotation pipeline. [271] leveraged active learning to select a few points and to form a minimal training set to avoid labeling the whole point cloud scenes. [272] introduced an efficient labeling framework with semi/weakly supervised learning to label outdoor point clouds.
Annotate 2D/3D Bounding Boxes. The quality of the bounding box annotations directly impacts the effectiveness and robustness of the perception system (like object detection) of autonomous vehicles in real-world scenarios. The annotation process generally involves labeling images with rectangular boxes or point clouds with cuboids to precisely encompass the objects of interest.
Labelme [273] is a prior tool focusing on labeling images for object detection. However, generating bounding boxes by professional annotators faces the same issue as manual segmentation annotation. Wang et al. [274] presented a semi-automatic video labeling tool based on the open-source video annotation system VATIC [^7] https://www.cs.columbia.edu/~vondrick/vatic/. [275] is another automatic video annotation tool for AD scenes. Compared to daytime annotation, process bounding box annotations in the nighttime is more challenging. [276] introduced a semi-automatic approach leveraging the trajectory to solve this problem.
In contrast to 2D annotations, 3D bounding boxes contain richer spatial information, such as accurate location, the object’s width, length, height, and orientation in space. Hence, labeling high-quality 3D annotations requires a more sophisticated framework. Meng et al. [277] applied a two-stage weakly supervised learning framework using human-in-the-loop to label LiDAR point clouds. ViT-WSS3D [278] generated pseudo-bounding boxes by modeling global interactions between LiDAR points and corresponding weak labels. Apolloscape [17] employed a labeling pipeline similar to [269], which consists of a 3D labeling and a 2D labeling branch, to handle static background/objects and moving objects, respectively. 3D BAT [279] developed an annotation toolbox to assist in obtaining 2D and 3D labels in semi-automatic labeling.
Annotate Trajectories. A trajectory is essentially a series of points that map the path of an object over time, reflecting the spatial and temporal information. Labeling trajectory data for AD is a process that entails annotating the path or movement patterns of various entities within a driving environment, such as vehicles, pedestrians, and cyclists. Usually, the annotating process relies on object detection and tracking results.
As one of the prior works in trajectory annotation, [280] online generated actions for maneuvers and were annotated into the trajectory. [281] consists of a crowd-sourcing step followed by a precise process of expert aggregation. [282] developed an active learning framework to annotate driving trajectory. Precisely anticipating movement patterns of pedestrians is critical for driving safety. Styles et al. [283] introduced a scalable machine annotation scheme for pedestrian trajectory annotations without human effort.
Annotate on Synthetic Data. Due to the expensive and time-consuming manual annotations on real-world data, synthetic data generated by computer graphics and simulators provide an alternative to address this issue. Since the data generation process is controllable and the attributes of each object in the scene (like position, size, and movement) are known, synthetic data can be automatically and accurately annotated.
The generated synthetic scenarios are designed to mimic real-world conditions, including multiple objects, various landscapes, weather conditions, and lighting variations. To achieve this goal, some researchers utilized the Grand Theft Auto 5 (GTA5) game engine to build datasets [4, 5]. [284] presented a real-time system based on multiple games to generate annotations for various AD tasks. Instead of applying game videos, SHIFT [10], CAOS [7], and V2XSet [11] were created based on the CARLA [112] simulator. Compared to [11], V2X-Sim [81] studied employing multiple simulators [112, 285] to generate dataset for V2X perception tasks. CODD [205] further exploited using [112] to generate 3D LiDAR point clouds for cooperative driving. Other works [3, 9, 2, 133] leveraged the Unity development platform [286] to generate synthetic datasets.
VI-B The quality of annotations
Existing supervised learning-based AD algorithms build upon numerous labeled data. However, training on low-quality annotations can negatively affect the safety and reliability of autonomous vehicles. Therefore, ensuring the quality of annotations is fundamental for improving accuracy while driving in complex real-world environments.
Figure 12: Mislabeling example of KITTI [13] dataset. We show the ground truth in blue. The bounding box of a car (in red circle) is not precise. Two cars (in green cube) are not annotated, although sensors obviously capture them.
According to the study [287], the annotation quality is affected by several factors, such as consistency, correctness, precision, and validation. Consistency is the foremost criterion in evaluating annotation quality. It involves maintaining uniformity across the entire dataset and is crucial for avoiding confusion in models trained on this data. For example, if a particular type of vehicle is labeled as a car, it should be consistently annotated the same in all other instances. Annotation precision is another vital indicator, which refers to whether the labels match the actual state of the objects or scenarios. In contrast, correctness highlights that annotated data are appropriate and relevant for the dataset purpose and annotation guidelines. After annotation, it is essential to validate the annotated data to ensure its accuracy and completion. The process can be done through manual review by experts or algorithms. Validation helps effectively prevent issues in datasets before they infect the performance of autonomous vehicles, decreasing potential safety risks. [288] presented a data-agnostic validation method for expert annotated datasets.
A failure case of annotation from KITTI [13] is shown in Fig. 12. We illustrate the ground truth bounding boxes (blue) in the corresponding image and LiDAR point cloud. On the left side of the image, the annotation of a car (circled in red) is inaccurate because it does not contain the whole object car. Additionally, two cars (highlighted by the green cuboid) are not annotated, even though the camera and LiDAR capture them clearly.
VII Data analysis
In this section, we systematically analyze datasets from different perspectives in detail, such as the distribution of data around the world (VII-A), chronological trend VII-B, and the data distribution VII-C.
VII-A Worldwide Distribution
Figure 13: The distribution of datasets around the world. This figure illustrates the distribution of data collection locations of the datasets.
We demonstrate an overview of the global distribution of 191 autonomous driving datasets in Fig. 13. The chart indicates that the USA is at the forefront with 40 datasets (21% share), underscoring its leadership in the autonomous driving domain. Germany accounts for 24 datasets, reflecting its robust automotive industry and influence on the advancement of autonomous technology. China is closely followed with 16 datasets, suggesting the interest and investment in this field of China. Another notable point is that there are 11 datasets collected worldwide and 24 from the European region (except Germany). This diverse regional distribution enhances the robustness of the collected data and highlights the international efforts and collaborations in the research community and industry.
On the other hand, although smaller segments represent a range of other countries, including Canada, Korea, the UK, Japan, and Singapore, these countries are developed countries with solid technological backgrounds and accumulation—an extreme regional bias reflected by this statistic. The dominance of the USA, western Europe, and East Asia leads to the bias where autonomous driving systems are overfitted to environmental conditions typical of these regions. This bias could result in autonomous vehicles’ failure to perform under varied or unseen regions and cases. Hence, introducing data from a more expansive array of countries and regions, such as Africa, can assist in the comprehensive development of autonomous vehicles.
Furthermore, 35 synthetic datasets generated by simulators like CARLA [112] take up 18.32% percent. Due to the limitation of recording from real-world driving environments, these synthetic datasets overcome such drawbacks and are critical for exploiting more robust and reliable driving systems.
VII-B Chronological Trends in Perception Datasets
In Fig. 10, we introduce a chronological overview of perception datasets with the top 50 impact scores from 2007 to 2023 (until the writing of this paper). The datasets are color-coded according to their sourcing domain, and synthetic datasets are marked with an external red outline, clearly illustrating the progress toward the diverse data collection strategy. A noticeable trend shows the increase in the number and variety of datasets over the years, indicating the requirement for high-quality datasets with the growing advancements in the field of autonomous driving.
In general, most of the datasets provide a perception perspective from the sensors equipped on the ego vehicle (onboard) because of the importance of the capability of the autonomous vehicle to efficiently and precisely precept the surroundings. On the other hand, due to the high-cost real-world data, some researchers propose high-influence synthetic datasets like VirtualKITTI [3] (2016) to alleviate the reliance on real data. Facilitated by the effectiveness of simulators, there are many novel synthetic datasets [7] [10] published in recent years. In the timeline, V2X datasets like DAIR-V2X [27] (2021) also exhibit a trend toward cooperative driving systems. Furthermore, because of the non-occlusion perspective provided by UAV, drone-based datasets, such as UAVDT [87] published in 2018, take a crucial position in advancing perception systems.
VII-C Data Distribution
Figure 14: Comparison of the distribution of the number of objects per frame across several datasets: Argoverse 2 [28], KITTI [13], nuScenes [22], ONCE [29], Waymo [23], and ZOD [30]. The horizontal axis quantifies the number of objects detected in a single frame, while the vertical axis represents the count of frames containing that number of objects.
Figure 15: Comparison of the distribution of the number of objects detected at various distances across several datasets: Argoverse 2 [28], KITTI [13], nuScenes [22], ONCE [29], Waymo [23], and ZOD [30]. The horizontal axis measures the distance from the ego vehicle in meters, and the vertical axis quantifies the number of objects detected at that distance.
We introduce an insight into the number of objects per frame for these datasets in Fig. 14. Notably, Waymo [23] exhibits an extreme number of frames with less than 50 objects while maintaining a broad presence across the chart, illustrating a wide range of scenarios from low to high object density per frame. Contrastingly, KITTI [13] shows a more constrained distribution and limited data size. Argoverse 2 [28] features a substantial number of frames with a higher object count—its peak appears around 70, which indicates its complex environmental representations in general. For ONCE [29], its density of objects evenly distributes in the supported perception range. Datasets like nuScenes [22] and ZOD [30] demonstrate similar curves with a quick rise and slow decline, implying a moderate level of environmental complexity with a decent variability of object counts per frame.
Beyond the number of objects in a scene, the object distribution based on the distance to the ego vehicle is another essential point for revealing a dataset’s variety and significant differences, illustrated in Fig. 15. The Waymo dataset demonstrates numerous labeled objects in near-field to mid-field scenarios. In contrast, Argoverse 2 and ZOD show a wider detection range, with some frames even including bounding boxes out of 200 meters. The curve of the nuScenes means it is particularly rich in objects in a shorter range, which is typical for urban driving scenarios. Nevertheless, as the distance increases, the nuScenes dataset quickly tappers off for the number of objects with annotations. The ONCE dataset covers a more even distribution of objects across distances, while the KITTI dataset focuses more on close-range detection.
VIII Discussion and Future Works
Dataset Year Size Temporal Sensing domain Tasks Real/Synthetic
BDD-X [289] 2018 8.4M $\checkmark$ onboard reasoning, planning real
Cityscapes-Ref [290] 2018 5,000 stereo videos $\checkmark$ onboard object referring real
TOUCHDOWN [291] 2019 9,326 examples $\checkmark$ onboard reasoning, navigation real
Talk2Car [292] 2019 11,959 commands, 850 videos $\checkmark$ onboard object referring real
BDD-OIA [293] 2020 11,303 scenarios $\times$ onboard explainable decision-making real
CityFlow-NL [294] 2021 5,289 samples $\checkmark$ onboard OT real
CARLA-NAV [295] 2022 83K $\checkmark$ onboard navigation real
NuPrompt [296] 2023 34K frames, 35K prompts $\checkmark$ onboard MOT real
NuScenes-QA [297] 2023 34K scenes, 460K QA pairs $\checkmark$ onboard visual question answering real
Refer-KITTI [51] 2023 6,650 frames $\checkmark$ onboard referring MOT real
Driving LLMs [298] 2023 10K driving situations, 160K QA pairs $\times$ drone visual question answering synthetic
DRAMA [299] 2023 17,785 scenarios $\checkmark$ onboard reasoning, visual question answering real
Rank2Tell [300] 2023 116 scenarios $\checkmark$ onboard importance level ranking real
LamPilot [301] 2023 4,900 samples $\checkmark$ others planning synthetic
LangAuto CARLAR [302] 2023 64K data clips $\checkmark$ onboard closed-loop driving synthetic
NuScenes-MQA [303] 2023 34K scenarios, 1.4M QA pairs $\times$ onboard visual question answering real
DriveMLM [304] 2023 280 hours $\checkmark$ onboard planning, control synthetic
DriveLM-nuScenes [305] 2023 4,871 frames $\checkmark$ onboard end-to-end driving real
DriveCARLA [305] 2023 183,373 frames $\checkmark$ onboard end-to-end driving real
LiDAR-text [306] 2023 420K 3D captioning data, 280K 3D grounding data $\times$ onboard 3D scene understanding real
TABLE VII: VLM autonomous driving dataset. OT: object tracking, MOT: multi-object tracking, QA: question-answering
In this paper, we mainly focus on analyzing the existing datasets, which usually include rich visual data, and aim at tasks in the modular-based pipeline. However, with the rapid technological development, especially the excellent performance of the Large Language Models, many novel trends of the next-generation autonomous driving datasets have occurred while proposing new challenges and requirements.
End-to-End Driving Datasets. Compared to the modular-designed autonomous driving pipeline, the end-to-end architecture simplifies the overall design process and reduces integration complexities. The success of the UniAD [307] verifies the potential ability of end-to-end models. However, the number of datasets for end-to-end AD is limited [167, 113]. Therefore, introducing datasets focusing on end-to-end driving is crucial for advancing autonomous vehicles. On the other hand, implementing an automatic labeling pipeline in a data engine can significantly facilitate the development of end-to-end driving frameworks and data [67].
Introduce Language into AD Datasets. The vision language models (VLMs) have recently achieved impressive advancement in many fields. Its inherent advantage in providing language information to vision tasks makes autonomous driving systems more explainable and reliable. [308] highlighted the prominent role of Multimodal Large Language Models in various AD tasks, such as perception [290, 296], motion planning [301], and motion control [111]. The autonomous driving datasets including language labels are shown in Tab. VII. Overall, incorporating language into AD datasets is a trend of the future development of AD datasets.
Data Generation via VLMs. As mentioned by [309], the powerful capability of VLMs can be used to generate autonomous driving data. For example, DriveGAN [310] generated high-quality AD data by disentangling different components without supervision. Additionally, due to the capability of world models for comprehending driving environments, some works [311, 312, 313] explored world models to generate high-quality driving videos. DriveDreamer[312] as a pioneering work derived from real-world scenarios, addressing the limitation of the gaming environments or simulated settings.
Domain Adaptation. Domain adaptation is a critical challenge in developing autonomous vehicles [314], referring to the ability of a model trained on one dataset (the source domain) to perform stable on another dataset (the target domain). This challenge manifests in multiple aspects, such as diversity in environmental conditions [315], sensor settings [316], or synthetic-to-real transform [317].
IX Conclusion
In this paper, we exhaustively and systematically reviewed and analyzed more than 200 existing autonomous driving datasets. We started with the sensor types and modalities, sensing domains, and tasks relevant to autonomous driving datasets. We introduced a novel evaluation metric called impact score to validate the influence and importance of perception datasets. Afterward, we demonstrated several high-impact datasets involving perception, prediction, planning, control, and end-to-end autonomous driving. Additionally, we explained the annotation methodology for autonomous driving datasets and investigated the factors affecting annotation quality.
Moreover, we delineated the chronological and geographical distribution of the collected datasets, providing a comprehensive perspective for understanding the current development of autonomous driving datasets. Meanwhile, we studied the data distribution of several datasets, offering a specific viewpoint for comprehending the variances across different datasets. In the end, we discussed the development and trend of the next generation of autonomous driving datasets.

### v2

@@
-A Survey on Autonomous Driving Datasets: Data Statistic, Annotation, and Outlook
+A Survey on Autonomous Driving Datasets: Statistics, Annotation Quality, and a Future Outlook
 Abstract
-Autonomous driving has rapidly developed and shown promising performance with recent advances in hardware and deep learning methods.
+Autonomous driving has rapidly developed and shown promising performance due to recent advances in hardware and deep learning techniques.
 High-quality datasets are fundamental for developing reliable autonomous driving algorithms.
-Previous dataset surveys tried to review the datasets but either focused on a limited number or lacked detailed investigation of the characters of datasets.
-To this end, we present an exhaustive study of over 200 autonomous driving datasets from multiple perspectives, including sensor modalities, data size, tasks, and contextual conditions.
-We introduce a novel metric to evaluate the impact of each dataset, which can also be a guide for establishing new datasets.
-We further analyze the annotation process and quality of datasets.
-Additionally, we conduct an in-depth analysis of the data distribution of several vital datasets.
-Finally, we discuss the development trend of the future autonomous driving datasets.
+Previous dataset surveys either focused on a limited number or lacked detailed investigation of dataset characteristics.
+To this end, we present an exhaustive study of 265 autonomous driving datasets from multiple perspectives, including sensor modalities, data size, tasks, and contextual conditions.
+We introduce a novel metric to evaluate the impact of datasets, which can also be a guide for creating new datasets.
+Besides, we analyze the annotation processes, existing labeling tools, and the annotation quality of datasets, showing the importance of establishing a standard annotation pipeline.
+On the other hand, we thoroughly analyze the impact of geographical and adversarial environmental conditions on the performance of autonomous driving systems.
+Moreover, we exhibit the data distribution of several vital datasets and discuss their pros and cons accordingly.
+Finally, we discuss the current challenges and the development trend of the future autonomous driving datasets.
+https://github.com/MingyuLiu1/autonomous_driving_datasets.git
 Index Terms:
-Dataset, influence, annotation, autonomous driving.
+Dataset, impact score, data analysis, annotation quality, autonomous driving.
 I Introduction
 Autonomous driving (AD) aims to revolutionize the transportation system by creating vehicles that can accurately perceive their environment, make intelligent decisions, and drive safely without human intervention.
-Due to thrilling technical development, various autonomous driving products have been implemented in several fields, such as robotaxi [1].
+Due to thrilling technical development, various autonomous driving products have been implemented in several fields, such as robotaxis [1].
 These rapid advancements in autonomous driving rely heavily on extensive datasets, which help autonomous driving systems be robust and reliable in complex driving environments.
 In recent years, there has been a significant increase in the quality and variety of autonomous driving datasets.
-The first apparent phenomenon in the development of datasets is the various data collection strategies, including synthetic datasets [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12] generated by simulators and recorded from the real world [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], to name just a few.
-Secondly, the datasets vary in composition, including but not limited to multiple sensory data (like camera images and LiDAR point clouds) and different annotation types for various tasks in autonomous driving.
+The first apparent phenomenon in the development of datasets is the various data collection strategies, including synthetic datasets [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12] generated by simulators and recorded from the real world [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], to name just a few.
+Secondly, the datasets vary in composition, including but not limited to multiple sensor modalities like camera images and LiDAR point clouds, different annotation types for various tasks, and data distribution.
 Fig.
-1 depicts the statistic of the 3D object bounding box distribution of six real-world datasets (Argoverse 2[28], KITTI [13], nuScenes [22], ONCE [29], Waymo [23], and ZOD [30]) under a bird’s-eye-view perspective, showing each dataset’s distinguished annotation characters.
-According to the equipment position of sensors, the variety of datasets also reflects in the sensing domains, containing onboard, V2X, drone, and others.
-Furthermore, the geometrical diversity and weather conditions improve the generalization of the autonomous driving datasets.
-Figure 1: BEV object distribution of datasets.
-Each heatmap corresponds to a dataset plotted with X and Y coordinates.
+1 depicts the distribution of 3D object bounding box in six well-known real-world datasets (Argoverse 2[28], KITTI [13], nuScenes [22], ONCE [30], Waymo [23], and ZOD [31]) under a Bird’s-Eye View (BEV), highlighting each dataset’s unique annotation characteristics.
+The sensor mounting positions also reflect datasets’ various sensing domains, including onboard, Vehicle-to-Everything (V2X), or drone domain.
+The datasets’ geometric diversity and varying weather conditions also enhance the generalizability of autonomous driving datasets.
+Figure 1: Bird’s-Eye View object distribution of datasets.
+Each heatmap represents a dataset and is plotted using X and Y coordinates.
 Y is the driving direction of the ego-vehicle.
@@
 We demonstrate the yearly published number of perception datasets in Fig.
-2 to reflect the trend of autonomous driving datasets from one perspective.
-Since numerous and continually increasing numbers of publicly published datasets exist, an exhaustive survey on autonomous driving datasets is valuable for advancing academic and industrial research.
+2 to illustrate the growth trends in autonomous driving datasets.
+Given the vast and growing number of publicly available datasets, a comprehensive survey of these resources is valuable for advancing both academic and industrial research in autonomous driving.
 In prior work, Yin et al.
 [32] summarized 27 publicly available datasets containing data collected on public roads.
-As a sequential work of  [31], [32] extended the dataset number.
-[33] and [34] proposed systematical introductions to the existing datasets from an application perspective.
-Beyond describing existing datasets, [35] discussed the domain adaptation between synthetic and real data and automatic labeling methods.
+As a sequential work of  [32], [33] extended the number of datasets.
+Guo et al.
+[34] and Janai et al.
+[35] proposed systematical introductions to the existing datasets from an application perspective.
+Beyond describing existing datasets, Liu et al.
+[36] discussed the domain adaptation between synthetic and real data and automatic labeling methods.
+Li et al.
 [37] summarized existing datasets and undertook an exhaustive analysis of the characters of the next-generation datasets.
@@
 AD-Dataset [38] collected a large number of datasets while lacking detailed analysis for the attributes of these datasets.
-Compared to studies on all types of datasets, some researchers presented surveys on a particular type of autonomous driving dataset, such as anomaly detection [38], synthetic datasets [39], 3D semantic segmentation [40], and decision-making [41].
-To this end, we aim to propose a comprehensive and systematic study on a large number of datasets in autonomous driving, covering all tasks from perception to control, considering real-world as well as synthetic data, and providing insight into the data modality and quality of several crucial datasets.
-We illustrate the comparison between other dataset surveys and ours in Tab I.
+Compared to studies on versatile datasets, some researchers presented surveys on a particular type of autonomous driving dataset, such as anomaly detection [39], synthetic datasets [40], 3D semantic segmentation [41], or decision-making [42].
+Additionally, some task-specific surveys [43, 44] also organized the related AD datasets.
+In this work, we present a comprehensive and systematic survey of a large number of datasets in autonomous driving.
+We compare our survey to others in Table I.
+Our survey covers all tasks from perception to control, considers real-world as well as synthetic data, and provides insights into the data modality and quality of several crucial datasets.
+TABLE I: We compare our survey paper with other AD dataset surveys in the following perspectives: collected dataset number (#Dataset), relevant tasks, sensing domain (S.
+domain), sensor modality (S.
+moda.), geometric conditions (Geo.), environmental conditions (Env.), analyzing data distribution, introducing annotation quality and process.
+In the context of environmental conditions, we refer to the variability in weather conditions and illumination.
+Geometric conditions include scenario types and geographical scope.
+We describe the task types in a coarse granularity, including perception (Perc.), prediction (Pred.), planning (Pl.), control (C.), 0and end-to-end (E2E).
 Survey General Data Annotation
@@
 Open-sourced Data Ecosystem [37] 2023 70 Perc, Pred, Pl $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-Ours 235 Perc, Pred, Pl, C, E2E $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-TABLE I: We compare our survey paper with other AD dataset surveys in the following perspectives: collected dataset number (#Dataset), relevant tasks, sensing domain (S.
-domain), sensor modality (S.
-moda.), geometric conditions (Geo.), environmental conditions (Env.), analyzing data distribution, introducing annotation quality and process.
-We describe the task types in a coarse granularity, including perception (Perc.), prediction (Pred.), planning (Pl.), control (C.) and end-to-end (E2E).
-Figure 2: Overview of the trend of publishing datasets.
-The diagram exhibits a rapid growth of onboard datasets between 2015 and 2020 and then slowly down.
-In contrast, there is an increment trend for the V2X datasets, showing the research trend on the cooperative perception systems.
+Ours 2024 265 Perc, Pred, Pl, C, E2E $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+Figure 2: Overview of dataset publication trends from 2008 to 2024.
+The diagram demonstrates a significant increase in the publication of onboard datasets between 2015 and 2020, followed by a gradual decline thereafter.
+In contrast, there has been a rising trend in the publication of V2X datasets, indicating growing research interest in cooperative perception systems.
 I-B Main Contributions
 The main contributions of this paper can be summarized as follows:
-  • We present an exhaustive survey on autonomous driving datasets.
-We consider publicly available datasets as comprehensively as possible, recording their fundamental characteristics, such as published year, data size, sensor modalities, sensing domains, geometrical and environmental conditions, and support tasks.
-To the best of our knowledge, we provide an overview of the most extensive collection of autonomous driving datasets recorded to date.
-  • We systematically illustrate the sensors and sensing domains for collecting autonomous driving data.
+  • We present an overview of the most exhaustive survey on autonomous driving datasets recorded to date.
+We show publicly available datasets as comprehensively as possible, recording their fundamental characteristics, such as published year, data size, sensor modalities, sensing domains, geometrical and environmental conditions, and support tasks.
+  • We systematically illustrate the sensors and sensing domains for collecting AD data.
 Furthermore, we describe the main tasks in autonomous driving, including task goals, required data modalities, and evaluation metrics.
-  • We summarize and divide the datasets according to their sensing domains and supporting tasks, which assists researchers in efficiently selecting and gathering information for the goal datasets.
-Thereby facilitating more targeted and effective research and development efforts.
-  • Additionally, we introduce an impact score metric to evaluate the influence of published perception datasets in the community.
-This metric can also be leveraged as a guidance for developing future datasets.
-We deeply analyze datasets with the highest grades based on the impact score, highlighting their advantages and utility.
+  • We categorize datasets based on their sensing domains and relevant tasks, enabling researchers to efficiently identify and compile information for their target datasets.
+This approach facilitates more focused and productive research and development efforts.
+  • Additionally, we introduce an impact score metric to evaluate the influence of published perception datasets.
+This metric can also be a guide for developing future datasets.
+We deeply analyze datasets with high impact scores, highlighting their advantages and utility.
   • We investigate the annotation quality of datasets and the existing labeling procedures for various autonomous driving tasks.
-  • Detailed data statistics is conducted to demonstrate the data distribution of various datasets from different perspectives, exhibiting their inherent limitations and suitable use situations.
-  • We analyze the recent technology trend and demonstrate the development direction of the next generation of datasets.
-We also provide a prospect on the potential effect of the advancement of Large Language Models on future autonomous driving.
+  • Our detailed data statistics demonstrate the data distribution of various datasets from different perspectives, exhibiting their inherent limitations and suitable use cases.
+  • We analyze recent technology trends and the development direction of next-generation datasets, such as integrating language into AD data, generating AD data using Vision Language Models, standardizing data creation, and promoting an open data ecosystem.
 I-C Scope & Limitations
-We aimed to conduct an exhaustive survey on the existing autonomous driving datasets to provide assistance and guidance for developing future algorithms and datasets in the domain.
-We collected datasets focusing on the four basic autonomous driving tasks: perception, prediction, planning, and control.
-Because several versatile datasets support multiple tasks, we only explained them in the main scope they support to avoid repeat introduction.
+We aim to conduct an exhaustive survey on the existing autonomous driving datasets to facilitate the development of future algorithms and datasets in this field.
+We collected datasets relevant to the five fundamental tasks of autonomous driving: perception, prediction, planning, control, and end-to-end (E2E) driving.
+To maintain clarity and prevent redundancy, we describe versatile datasets only within the main scope they support.
 Additionally, we collected a large number of datasets and exhibited them with their primary characters in tables.
-Nevertheless, a detailed explanation of all collected datasets could not highlight the most popular ones, and it can hurt assisting researchers in finding valuable datasets through this survey.
-Hence, we only describe the most impactful datasets in detail.
-I-D Structure of the Article
-The rest of the article is structured as follows: Section II introduces the approach we leveraged to obtain public datasets and the evaluation metrics for datasets.
+However, to ensure this survey aids researchers effectively, we focus our detailed discussions on the most impactful datasets rather than providing extensive descriptions of all datasets.
+I-D Survey Structure
+The rest of the survey is structured as follows: Section II introduces the approach leveraged to source public datasets and the evaluation metrics for them.
 Section III demonstrates the primary sensors used in autonomous driving and their modalities.
 Section IV discusses autonomous driving tasks, related challenges, and required data.
-We further discuss several important datasets in Section V.
-The process of annotations and factors affecting annotation quality are exhibited in Section VI.
-Moreover, we statistic the data distribution of several datasets in section VII.
-In section VIII, we investigate the developing trend and future works of autonomous driving datasets.
-In the end, we conclude our work in Section IX.
-The taxonomy of this survey is shown in Fig.
+In-depth discussions of several important datasets are presented in Section V.
+The process of annotations and factors affecting annotation quality are addressed in Section VI.
+Section VII provides statistical analysis of data distribution across various datasets.
+Future trends and potential research directions in autonomous driving datasets are explored in Section VIII.
+The paper concludes with Section IX.
+The survey’s taxonomy is shown in Fig.
 3.
-Figure 3: This survey’s primary taxonomy includes impact score, sensors and modalities, autonomous driving tasks, important datasets, and annotation process.
+Figure 3: This survey’s primary taxonomy includes impact score, sensors and modalities, autonomous driving tasks, high-influence datasets, and annotation process.
 II Methodology
-This section consists of 1) How we collect and filter datasets (II-A), and 2) How to evaluate the impact of a dataset on the autonomous driving domain (II-B).
+This section comprises two parts: 1) the collection and filtering of datasets, as detailed in Section II-A, and 2) the evaluation metric of a dataset’s impact on the autonomous driving domain, described in Section II-B.
 II-A Datasets Collection
 Following [45], we conducted a systematic review to exhaustively collect published autonomous driving datasets.
-To ensure source diversity, we utilized well-known search engines such as Google [^1] https://www.google.com/, Google Scholar [^2] https://scholar.google.com/ and Baidu [^3] https://www.baidu.com/ to search datasets.
-To ensure a thorough dataset collection from various countries and regions, we conducted searches in English, Chinese, and German using keywords such as ”autonomous driving datasets,” ”intelligent vehicle datasets,” and terms related to object detection, classification, tracking, segmentation, prediction, planning, and control.
-Furthermore, we explored IEEE Xplore [^4] https://ieeexplore.ieee.org/Xplore/home.jsp and pertinent conferences in autonomous driving and intelligent transportation systems to collate datasets from journals and conference proceedings.
+To ensure source diversity, we utilized well-known search engines such as Google, Google Scholar and Baidu to search datasets.
+To ensure a thorough dataset collection from various countries and regions, we conducted searches in English, Chinese, and German using keywords such as “autonomous driving dataset/benchmark”, “intelligent vehicle dataset/benchmark” and terms related to object detection, classification, tracking, segmentation, prediction, planning, control, and end-to-end driving.
+Furthermore, we explored IEEE Xplore, Paperswithcode, and pertinent conferences in autonomous driving and intelligent transportation systems to collate datasets from journals and conference proceedings.
 We verified datasets from these sources through keyword searches and manual title reviews.
-Finally, to ensure the inclusion of specialized or lesser-known datasets, we searched through Github [^5] https://github.com/ repositories and Paperwithcodes [^6] https://paperswithcode.com/.
+Finally, to ensure the inclusion of specialized or lesser-known datasets, we searched through Github repositories.
 Similar to databases, we performed both manual and keyword-based searches for datasets.
@@
 First, we calculate the citation scores from the total citation number and average annual citation.
-To gain fair citation counts, we choose the time of the earliest version of a dataset as its publish time.
-Moreover, all citation counts were collected as of September 20, 2023, to ensure the comparison is based on a consistent timeframe.
-The total citation number $\mathbf{c}^{t}$ reflects the overall influence of a dataset.
+To gain fair citation counts, we choose the time of the earliest version of a dataset as its publication time.
+Moreover, all citation counts were collected as of March 05, 2024, to ensure the comparison is based on a consistent timeframe.
+The total citation number $c^{t}$ reflects the overall influence of a dataset.
 A higher count of this metric means that the dataset has been widely recognized and utilized by researchers.
 However, datasets published in earlier years can accumulate more citations.
-To address this unfairness, we leverage average annual citation $\mathbf{c}^{a}$, which describes the yearly citation increase speed of a dataset.
-The calculate function is shown in Eq.
+To address this unfairness, we leverage average annual citation $c^{a}$, which describes a dataset’s yearly citation increase speed.
+The function is shown in Eq.
 1.
-$\mathbf{c}^{a}=\begin{cases}\mathbf{c}^{t}/\left(\mathbf{y}_{curr}-\mathbf{y}_{pub}\right)&\text{if }\mathbf{y}_{curr}\neq\mathbf{y}_{pub}\\ \mathbf{c}^{t}&\text{if }\mathbf{y}_{curr}=\mathbf{y}_{pub}\end{cases}$ (1)
-where $\mathbf{y}_{curr}$ and $\mathbf{y}_{pub}$ represent the current year and the dataset published year, respectively.
+$c^{a}=\begin{cases}c^{t}/\left(y_{curr}-y_{pub}\right)&\text{if }y_{curr}\neq y_{pub}\\ c^{t}&\text{if }y_{curr}=y_{pub}\end{cases}$ (1)
+where $y_{curr}$ and $y_{pub}$ represent the current year and the dataset published year, respectively.
 On the other hand, the citation number of distastes has a wide distribution range from single digits to tens of thousands.
-To alleviate the extreme unbalance and highlight the differences of each dataset, we apply logarithmic transformation followed by Min-Max normalization to both $\mathbf{c}^{t}$ and $\mathbf{c}^{a}$, described in Eq.
+To alleviate the extreme imbalance and highlight the differences of each dataset, we apply a logarithmic transformation followed by Min-Max normalization to both $c^{t}$ and $c^{a}$, described in Eq.
 2.
-$\mathbf{c}_{norm}=\min\text{-}\max\left(\log(\mathbf{c})\right)$ (2)
-Finally, the citation score $\mathbf{c}_{score}$ is the summation of $\mathbf{c}^{t}_{norm}$ and $\mathbf{c}^{a}_{norm}$:
-$\mathbf{c}_{score}=0.5\*\mathbf{c}^{t}_{norm}+0.5\*\mathbf{c}^{a}_{norm}$ (3)
+$c_{norm}=\min\text{-}\max\left(\log(c)\right)$ (2)
+Finally, the citation score $c_{score}$ is the summation of $c^{t}_{norm}$ and $c^{a}_{norm}$:
+$c_{score}=0.5\*c^{t}_{norm}+0.5\*c^{a}_{norm}$ (3)
 Data Dimension Score.
 We measure the data dimension across four perspectives: dataset size, temporal information, task number, and labeled categories.
-Dataset size $\mathbf{f}$ is represented by the frame number of a dataset, reflecting its volume and comprehensiveness.
-To get the dataset size score $\mathbf{f_{normal}}$, we leverage the same method as the citation score to process the frame number to overcome the extreme imbalance between different datasets.
+Dataset size $f$ is represented by the frame number of a dataset, reflecting its volume and comprehensiveness.
+To get the dataset size score $f_{norm}$, we leverage the same method as the citation score to process the frame number to overcome the extreme imbalance between different datasets.
 Temporal information is essential for autonomous driving as it enables the vehicle to understand how the surrounding environment changes over time.
-We use $\mathbf{t}\in\left\{0,1\right\}$ to indicate whether a dataset includes temporal information.
+We use $t\in\left\{0,1\right\}$ to indicate whether a dataset includes temporal information.
 Regarding the task number, we only consider datasets related to the six fundamental tasks in the autonomous driving perception domain, such as 2D object detection, 3D object detection, 2D semantic segmentation, 3D semantic segmentation, tracking, and lane detection.
-Therefore, the task number score is recorded as $\mathbf{t_{n}}\in\left\{1,2,3,4,5,6\right\}$.
+Therefore, the task number score is recorded as ${t^{n}}\in\left\{1,2,3,4,5,6\right\}$.
 A large number of categories is critical for the robustness and versatility of a dataset.
 During the statistic, if a dataset supports multiple tasks and includes various types of annotation, we choose the largest number of categories.
-Afterward, the categories are divided into five levels, $\mathbf{l}=\left\{1,2,3,4,5\right\}$, based on quintiles.
-To simplify the calculation, we normalize $\mathbf{t_{n}}$ and $\mathbf{l}$ before the following process.
-In order to reflect the data dimension score $\mathbf{d}_{score}$ as objectively as possible, we give different weights to the four components, as shown in Eq. 4.
-$\mathbf{d}_{score}=0.5\*\mathbf{f}_{norm}+0.1\*\mathbf{t}+0.2\*\mathbf{t_{n}}_{norm}+0.2\*\mathbf{l}_{norm}$ (4)
+Afterward, the categories are divided into five levels, $cat=\left\{1,2,3,4,5\right\}$, based on quintiles.
+We normalize $t_{n}$ and $l$ before the following process to simplify the calculation.
+In order to reflect the data dimension score $d_{score}$ as objectively as possible, we give different weights to the four components, as shown in Eq. 4.
+$d_{score}=0.5\*f_{norm}+0.1\*t+0.2\*t^{n}_{norm}+0.2\*cat_{norm}$ (4)
 Environmental Diversity Score.
-We evaluate the environmental diversity of a dataset according to the following factors: 1) weather conditions, such as rain or snow.
-2) time of day refers to when the data is collected, like morning or dusk.
-3) the types of driving scenarios, e.g., urban or rural.
-4) the geometric scope means the number of countries or cities where the data is recorded.
-It is worth noting that we treat the geometric scope for synthetic datasets as missing.
+We evaluate the environmental diversity of a dataset according to the following factors: 1) weather conditions (such as rain or snow), 2) time of day (e.g., morning or dusk), 3) types of driving scenarios (like urban or rural), and 4) geometric scope refers to the number of countries or cities where the data is recorded.
+It is worth noting that we treat the geographical scope for synthetic datasets as undefined.
 We follow the granularity with which a paper categorizes its data to quantify the diversity.
-Moreover, for the missing value, if a dataset announces that the data is recorded under diversity conditions, we use the median value as the missing value.
-Otherwise, we set the missing value of this attribute to one.
+Moreover, for the missing value, if a dataset announces that the data is recorded under diversity conditions, we assign the median value; otherwise, the missing value is set to one.
 We apply quintiles to quantify each factor into five distinct levels.
-After that, the environmental diversity score $\mathbf{e}_{score}$ is the sum of these four factors.
-In the end, we leverage Eq. 5 to calculate the impact score $\mathbf{i}_{score}$.
-$\mathbf{i}_{score}=60\*\mathbf{c}_{score}+20\*\mathbf{d}_{score}+20\*\mathbf{e}_{score}$ (5)
-The total impact score is 100, and 60 percent belongs to the citation score $\mathbf{c}_{score}$, data dimension score $\mathbf{d}_{score}$ and environmental diversity score $\mathbf{e}_{score}$ takes 40 percent.
-III Data sources and Cooperative Perception in Autonomous Driving
+After that, the environmental diversity score $e_{score}$ is the sum of these four factors.
+In the end, we leverage Eq. 5 to calculate the impact score $i_{score}$.
+$i_{score}=60\*c_{score}+20\*d_{score}+20\*e_{score}$ (5)
+The total impact score is 100, and 60 percent belongs to the citation score $c_{score}$. Data dimension score $d_{score}$ and environmental diversity score $e_{score}$ takes 40 percent.
+III Sensors and Perception Technology in Autonomous Driving
 In this section, we introduce the sensors and their modalities mainly used in autonomous driving (III-A).
-Furthermore, in III-B, we analyze the data acquisition and communication domain, such as onboard, drone, and vehicle-to-everything.
-III-A Sensors and Modalities of Data
+Subsequently, in III-B, we analyze the data acquisition methods and cooperative perception technologies.
+III-A Sensor Modalities
 (a) Camera
@@
 (f) Thermal camera
-Figure 4: Sensors on autonomous driving vehicle.
+Figure 4: Sensors on autonomous driving vehicles.
 The type of each sensor is (a) Camera: Basler ace acA1600-20uc, (b) LiDAR: Velodyne Puck LITE, (c) Radar: Ainstein Launches K-79, (d) Event-based camera: Evaluation Kit 4 HD, (e) IMU: IMU383_Aceinna-W and (f) Thermal camera: FLIR_2nd_Gen_ADK.
 All figures are extracted from the websites hosting the sensors.
-Figure 5: Modalities of sensors. We demonstrate each type of sensor’s modality to intuitively understand each sensor’s characters.
+Figure 5: We present the sensor modalities to provide an intuitive understanding of each sensor’s characteristics.
+(a) is from nuScenes [22], (b) is from KITTI [13], (c) is from [46], (d) is from [47], (e) is from [48].
+All figures are collected from the open-source data of datasets.
 Efficiently and precisely collecting data from the surrounding environment is the key to a reliable perception system in autonomous driving.
@@
 4.
-The most used sensors are cameras, LiDAR, and radars.
-Event-based and thermal cameras are also installed in vehicles or roadsides to improve the perception capability further.
+The most used sensors are cameras, LiDARs, and radars.
+Event-based and thermal cameras are also installed on vehicles or roadside infrastructure to improve the perception capability further.
 RGB Images.
@@
 Monocular cameras offer a 2D view without depth; stereo cameras, with their dual lenses, provide depth perception; fisheye cameras use wide-angle lenses to capture a broad view.
-All these cameras channel light through a lens onto an image sensor, e.g., CMOS, converting this light into electronic signals representing an image.
 As shown in Fig.
@@
 LiDAR Point Clouds.
-LiDAR uses laser beams to measure the distance between the sensor and an object, creating a 3D environment representation [44].
-The LiDAR point clouds (Fig.
+LiDARs use laser beams to measure the distance between the sensor and an object, creating a 3D environment representation [50].
+LiDAR point clouds (Fig.
 5 (b)) provide precise spatial information with high resolution and can detect objects over long distances.
 However, the density of these points can decrease with increasing distance, leading to sparser representations for distant objects.
-The weather conditions, e.g., fog, also limit the performance of LiDAR.
-In general, LiDAR is suitable in cases that require 3D concise information.
+Weather conditions, e.g., fog, also limit the performance of LiDAR.
+In general, LiDARs are suitable in cases that require 3D concise information.
 Radar Point Clouds.
@@
 In autonomous driving, it is used to track the movement and orientation of the vehicle.
-Although IMU does not include visual information of the surrounding environment, the perception systems can achieve more accurate and robust tracking of a vehicle’s motion and orientation by fusing the data from an IMU with data from other sensors.
-Figure 6: Sensor number distribution. We statistic the distribution of different sensors. Overall, RGB cameras and LiDAR are the most used sensors in autonomous driving datasets.
 We analyze the sensor distribution from the collected datasets, shown in Fig.
 6.
-More than half of the sensors are monocular cameras (53.85%) due to their low price and reliable performance.
-Additionally, 93 datasets include LiDAR data, valued for its high resolution and precise spatial information.
+More than half of the sensors are monocular cameras (52.79%) due to their low price and reliable performance.
+Additionally, 93 datasets (25.98%) include LiDAR data, valued for its high resolution and precise spatial information.
 However, its high-cost limits LiDAR’s widespread use.
 Beyond LiDAR point clouds, 29 datasets leverage stereo cameras to capture depth information.
-Furthermore, 5.41%, 3.42%, and 1.71% datasets include radar, thermal camera, and fisheye camera, respectively.
-Given the temporal efficiency of capturing dynamic scenes, three datasets generate data based on event-based cameras.
+Furthermore, 5.31%, 3.35%, and 1.68% datasets include radar, thermal camera, and fisheye camera.
+Given the temporal efficiency of capturing dynamic scenes, ten datasets generate data based on event-based cameras (2.79%).
+Figure 6: Sensor type distribution. We show the distribution of different sensors. Overall, RGB cameras and LiDARs are the most used sensors in autonomous driving datasets.
 III-B Sensing Domains and Cooperative Perception Systems
-The sensory data and communication between the ego vehicle and other entities in the surrounding environment play a pivotal role in ensuring self-driving systems’ safety, efficiency, and overall functionality.
+Sensory data and cooperation between the ego vehicle and surrounding entities are pivotal for ensuring autonomous driving systems’ safety, efficiency, and overall functionality.
 Therefore, the positioning of sensors is crucial as it determines the quality, angle, and scope of data that can be collected.
-Generally, sensors in the autonomous driving environment can be categorized into the following domains: onboard, Vehicle-to-Everything (V2X), drone, and others.
+Generally, sensing domains in autonomous driving can be categorized as onboard, Vehicle-to-Everything (V2X), drone-based, and others.
 Figure 7: Overview of cooperative perception systems in autonomous driving.
-A complete autonomous driving perception system consists of the ego vehicle and the cooperation between vehicles, infrastructure, and networks.
+A complete autonomous driving perception system consists of the ego vehicle along with collaborative interactions between vehicles, infrastructure, and networks.
 Onboard.
-Onboard sensors are installed directly on the autonomous driving vehicle and usually consist of cameras, LiDAR, radars, and IMUs.
+Onboard sensors are installed directly on the autonomous vehicle and usually consist of cameras, LiDARs, radars, and IMUs.
 These sensors provide a direct perspective from the vehicle’s standpoint, offering immediate feedback on the surroundings.
-Nevertheless, due to the limitation of the vehicle detection scope, onboard sensors may have limitations in providing advanced warning about obstacles in blind spots or detecting hazards around sharp bends.
+Nevertheless, due to the limited detection scope, onboard sensors may have limitations in providing advanced warnings about obstacles in blind spots or detecting hazards around sharp bends.
 V2X.
-Vehicle-to-Everything (V2X) encompasses communications between a vehicle and any other components in the transport system [49], including V2V, V2I, and V2N (as shown in Fig.
+Vehicle-to-Everything encompasses communications between a vehicle and any other components in the transport system [55], including Vehicle-to-Vehicle, Vehicle-to-Infrastructure, and Vehicle-to-Network (as shown in Fig.
 7).
@@
   • Vehicle-to-Vehicle (V2V)
-    V2V enables nearby vehicles to share data, including their positions and velocity and sensory data, like camera images or LiDAR scans, which assists in creating a more comprehensive understanding of the driving scenarios.
+    V2V enables nearby vehicles to share essential data, such as position, velocity, and captured sensory data (e.g., camera images or LiDAR scans).
+This shared information contributes to a more comprehensive understanding of driving scenarios.
   • Vehicle-to-Infrastructure (V2I)
     V2I facilitates communications between the autonomous vehicle and infrastructure components such as traffic lights, signs, or roadside sensors.
-The sensors embedded in the road infrastructure, including cameras, LiDAR, radars, or event-based cameras, work synergistically to extend the perception range and improve the situational awareness of autonomous vehicles.
-In this survey, we consider both perceptions through infrastructure or V2I belonging to V2I.
+The sensors implemented in road infrastructure collaborate to enhance the perception range and situational awareness of autonomous vehicles.
+In this survey, we categorize interactions involving single or multiple vehicles with single or multiple infrastructure components, as well as collaborations among multiple infrastructure elements, under V2I.
   • Vehicle-to-Network (V2N)
@@
 Drone.
-Drones, or UAVs, offer an aerial perspective, providing data essential for trajectory prediction and route planning [18] .
+Drones, or Unmanned Aerial Vehicles (UAVs), offer an aerial perspective, providing data essential for trajectory prediction and route planning [18].
 For example, the real-time data from drones can be integrated into traffic management systems to optimize the traffic flow and alert autonomous vehicles of accidents ahead.
@@
 Fig.
-9 illustrates examples of several main tasks in autonomous driving.
+9 highlights several main tasks in autonomous driving.
 Figure 8: The overview of autonomous driving pipeline.
 Autonomous driving systems can be categorized into two types: modular-based and end-to-end.
-Both rely on the data collected by various sensors installed on the vehicles or infrastructures and interact with the surrounding driving experiments.
+Both types rely on data collected by various sensors installed on vehicles or infrastructures.
+These systems interact with and respond to the surrounding environment during driving scenarios.
 IV-A Perception and Localization
@@
 Figure 9: Examples of various autonomous driving tasks.
-(a) is from KITTI [13], (b) is from Cityscapes [14], (c) is from V2X-Seq [50], (d) is from BDD100K [24], (e) is from Refer-KITTI [51], (f) is from KITTI-360 [52], (g) is from Dr(eye)ve [53], (h) is from Argoverse 2 [28].
+(a) is from KITTI [13], (b) is from Cityscapes [14], (c) is from V2X-Seq [56], (d) is from BDD100K [24], (e) is from Refer-KITTI [57], (f) is from KITTI-360 [58], (g) is from Dr(eye)ve [59], (h) is from TUMTraf [60].
 All figures are collected from the open-source data of datasets or the websites hosting the datasets.
@@
 2D/3D Object Detection.
-2D or 3D object detection aims to identify and classify other entities within the driving environment.
-While 2D object detection identifies objects in image space, 3D object detection further incorporates precise depth information, often provided by LiDAR.
+2D or 3D object detection aims to identify the locations and classification of other entities within the driving environment.
 Although the detection technologies have significantly advanced, several challenges remain, such as object occlusions, varying light conditions, and diverse object appearances.
@@
 And the mean Intersection over Union (mIoU):
-$mIoU=\frac{1}{k=1}\sum_{i=0}^{k}\frac{p_{ii}}{\sum_{j=0}^{k}p_{ij}+\sum_{j=0}^{k}p_{ji}-p_{ii}}$ (8)
+$mIoU=\frac{1}{k+1}\sum_{i=0}^{k}\frac{p_{ii}}{\sum_{j=0}^{k}p_{ij}+\sum_{j=0}^{k}p_{ji}-p_{ii}}$ (8)
 Where $k\in\mathbb{N}$ is the number of classes, and $p_{ii}$, $p_{ij}$, and $p_{ji}$ represent true positives, false positives, and false negatives, respectively.
@@
 According to [28], HD map automation [65] and HD map change detection [66] have received more and more attention.
-Normally, the HD map quality is estimated using the accuracy metric.
+Usually, the HD map quality is estimated using the accuracy metric.
 SLAM.
@@
 Hence, data from cameras, IMUs for position tracking, and real-time LiDAR point clouds are vital.
+Sturm et al.
 [67] introduces two evaluation metrics, relative pose error (RPE) and absolute trajectory error (ATE), for evaluating the quality of the estimated trajectory from the input RGB-D images.
@@
 Trajectory Prediction.
-Using time-series data from sensors like cameras and LiDAR, trajectory prediction pertains to anticipating the future paths or movement patterns of other entities [61], such as pedestrians, cyclists, or other vehicles.
+Based on time-series data from sensors like cameras and LiDARs, trajectory prediction pertains to anticipating the future paths or movement patterns of other entities [68], such as pedestrians, cyclists, or other vehicles.
 Behavior Prediction.
-Behavior prediction anticipates the potential actions of other road users [62], e.g., whether a vehicle will change lanes.
+Behavior prediction anticipates the potential actions of other road users [69], e.g., whether a vehicle will change the lane.
 Training behavior prediction models rely on extensive annotated data due to entities’ vast range of potential actions within various scenarios.
 Intention Prediction.
-Intention prediction focuses on inferring the higher-level goals of the intention behind the actions of objects, involving a deeper semantic comprehension of the physical or mental activities of humans [64].
-Because of the task’s complexity, it requires data not only from perception sensors like cameras but also additional information, such as traffic signals and hand gestures, to infer the intentions of other agents.
+Intention prediction focuses on inferring the higher-level goals behind the actions of objects, involving a deeper semantic comprehension of the physical or mental activities of humans [71].
+Because of the task’s complexity, it requires data from perception sensors like cameras, traffic signals, and hand gestures.
 IV-C Planning and Control
@@
 Planning represents the decision-making process in reaction to the perceived environment and predictions.
-A classic three-level hierarchical planning framework consists of path, behavioral, and motion planning [53].
+A classic three-level hierarchical planning framework comprises path, behavioral, and motion planning [59].
 Path Planning Path planning, also known as route planning, involves setting long-term objectives. It is a high-level process of determining the best path to the destination.
@@
 Their inherent promise is in their simplicity and efficiency by reducing the need for hand-crafted components [73].
-However, implementing end-to-end models faces critical limitations such as extensive training data requirements, low interpretability, and inflexible modular tuning.
+However, implementing end-to-end models faces limitations such as extensive training data requirements, low interpretability, and inflexible modular tuning.
 Large-scale benchmarking for end-to-end AD can be categorized into closed-loop and open-loop evaluation [74].
@@
 We also exhibit datasets for end-to-end autonomous driving (V-C).
-Figure 10: Chronological overview. We show the top 50 datasets with the highest impact scores. This diagram records datasets from 2007 to 2023 (until the writing of this paper).
-Dataset Year Size Temp Tasks Categories Weather Time of day Scenario Geometrical Impact
+Figure 10: Chronological overview. We show the top 50 datasets ranked by their impact scores. This diagram covers datasets released from 2009 through March 17, 2024.
+TABLE II: High-impact perception datasets.
+For a more comprehensive demonstration, we exhibit 50 perception datasets from different sensing domains instead of those with the highest scores.
+The citation number of each dataset is not shown in the table due to the fast change for this point.
+Dataset Year Size Temp Tasks Categories Weather Time of day Scenario Geographical Impact
 2D Det 3D Det 2D Seg 3D Seg Tracking Lane Det number conditions type scope score
 Onboard
-nuScenes [22] 2019 40K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23 3 2 4 2 81.20
-Cityscapes [14] 2016 25K $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 30 1 1 3 1 79.00
-BDD100K [24] 2020 12M $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 40 5 2 4 1 77.82
-Waymo [23] 2019 230K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23 2 3 5 1 77.81
-KITTI [13] 2012 41K $\checkmark$ $\checkmark$ $\checkmark$ 8 1 1 2 1 77.30
-SYNTHIA [2] 2016 13.4K $\times$ $\checkmark$ 13 2 3 5 - 70.44
-Apolloscape [17] 2018 143,906 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 28 4 3 3 1 68.97
-Virtual KITTI [3] 2016 21,260 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 8 3 2 5 - 67.61
-VIPER [5] 2017 254,064 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 11 4 3 4 - 67.07
-SemanticKITTI [6] 2019 43,552 $\checkmark$ $\checkmark$ 28 1 1 4 1 66.66
-GTA5 [4] 2016 24,966 $\times$ $\checkmark$ 19 2 2 2 - 66.10
-Argoverse 2 [28] 2023 6M $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 30 2 1 1 6 61.11
-Lane Det [68] 2017 133,235 $\times$ $\checkmark$ 1 1 3 3 1 59.62
-CityPersons [69] 2017 25K $\times$ $\checkmark$ $\checkmark$ 30 2 1 - 27 58.84
-CamVid [70] 2009 701 $\times$ $\checkmark$ 32 2 1 2 1 58.69
-IDD [71] 2019 10,004 $\times$ $\checkmark$ 34 3 3 5 1 58.12
-Foggy Cityscapes [72] 2018 20,550 $\times$ $\checkmark$ $\checkmark$ 19 1 1 2 1 58.09
-A2D2 [25] 2020 41,277 $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 38 2 1 3 3 58.06
-Virtual KITTI 2 [9] 2020 20,992 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 8 4 2 3 1 56.74
-RADIATE [26] 2021 44,140 $\checkmark$ $\checkmark$ $\checkmark$ 8 5 2 4 1 56.70
-GTSDB [73] 2013 900 $\times$ $\checkmark$ 4 2 2 3 1 56.68
-Caltech Ped [74] 2009 250K $\times$ $\checkmark$ 1 1 1 1 2 56.64
-SHIFT [10] 2022 2,5M $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23 5 5 3 - 55.79
-CAOS [7] 2022 13K $\times$ $\checkmark$ 13 3 2 3 - 55.74
-STF [75] 2020 13,500 $\times$ $\checkmark$ $\checkmark$ 1 4 2 3 4 54.91
-KITTI-360 [52] 2021 150K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 37 1 1 1 1 54.77
-ACDC [76] 2021 4,006 $\times$ $\checkmark$ 19 4 2 3 1 54.62
-VPGNet [77] 2017 20K $\times$ $\checkmark$ 17 3 2 1 1 54.37
-ONCE [29] 2021 1M $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ 5 3 2 5 1 54.09
+nuScenes [22] 2019 40K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23 3 2 4 2 83.03
+Cityscapes [14] 2016 25K $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 30 1 1 3 1 79.87
+BDD100K [24] 2020 12M $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 40 5 2 4 1 79.79
+Waymo [23] 2019 230K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23 2 3 5 1 79.66
+KITTI [13] 2012 41K $\checkmark$ $\checkmark$ $\checkmark$ 8 1 1 2 1 78.04
+SYNTHIA [2] 2016 13.4K $\times$ $\checkmark$ 13 2 3 5 - 71.16
+Apolloscape [17] 2018 143,906 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 28 4 3 3 1 69.91
+SemanticKITTI [6] 2019 43,552 $\checkmark$ $\checkmark$ 28 1 1 4 1 68.40
+Virtual KITTI [3] 2016 21,260 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 8 3 2 5 - 68.27
+VIPER [5] 2017 254,064 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 11 4 3 4 - 67.96
+GTA5 [4] 2016 24,966 $\times$ $\checkmark$ 19 2 2 2 - 66.87
+Argoverse 2 [28] 2023 6M $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 30 2 1 1 6 65.28
+Lane Det [75] 2017 133,235 $\times$ $\checkmark$ 1 1 3 3 1 60.71
+SHIFT [10] 2022 2,5M $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23 5 5 3 - 60.67
+CityPersons [76] 2017 25K $\times$ $\checkmark$ $\checkmark$ 30 2 1 - 27 59.75
+A2D2 [25] 2020 41,277 $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 38 2 1 3 3 59.42
+Foggy Cityscapes [77] 2018 20,550 $\times$ $\checkmark$ $\checkmark$ 19 1 1 2 1 59.38
+CamVid [78] 2009 701 $\times$ $\checkmark$ 32 2 1 2 1 59.27
+IDD [79] 2019 10,004 $\times$ $\checkmark$ 34 3 3 5 1 59.18
+CAOS [7] 2022 13K $\times$ $\checkmark$ 13 3 2 3 - 58.98
+RADIATE [26] 2021 44,140 $\checkmark$ $\checkmark$ $\checkmark$ 8 5 2 4 1 58.78
+Virtual KITTI 2 [9] 2020 20,992 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 8 4 2 3 1 58.67
+KITTI-360 [58] 2021 150K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 37 1 1 1 1 58.48
+Dr(eye)ve [59] 2018 555000 $\times$ $\checkmark$ 10 3 3 3 - 58.13
+ACDC [80] 2021 4,006 $\times$ $\checkmark$ 19 4 2 3 1 57.64
+GTSDB [81] 2013 900 $\times$ $\checkmark$ 4 2 2 3 1 57.48
+ONCE [30] 2021 1M $\times$ $\checkmark$ $\checkmark$ 5 3 2 5 1 57.03
+Caltech Ped [82] 2009 250K $\times$ $\checkmark$ 1 1 1 1 2 56.98
+STF [83] 2020 13,500 $\times$ $\checkmark$ $\checkmark$ 1 4 2 3 4 56.85
 V2X
-DAIR-V2X [27] 2021 71,254 $\checkmark$ $\checkmark$ $\checkmark$ 10 2 2 2 1 49.21
-V2XSet [11] 2022 11,447 $\checkmark$ $\checkmark$ $\checkmark$ 1 1 1 1 - 45.39
-V2VNet [78] 2020 51,2K $\checkmark$ $\checkmark$ 1 1 1 1 - 44.94
-T&J [79] 2019 100 $\checkmark$ $\checkmark$ $\checkmark$ 1 1 1 2 1 44.58
-Rope3D [80] 2022 50K $\times$ $\checkmark$ $\checkmark$ 12 3 3 2 1 44.39
-V2X-Sim [81] 2022 10K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23 1 1 1 - 41.00
+TUMTraf [84, 85, 86, 60] 2022 50,253 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 10 6 5 10 1 55.51
+DAIR-V2X [27] 2021 71,254 $\checkmark$ $\checkmark$ $\checkmark$ 10 2 2 2 1 54.58
+V2XSet [11] 2022 11,447 $\checkmark$ $\checkmark$ $\checkmark$ 1 1 1 1 - 49.84
 V2V4Real [87] 2023 40K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 5 2 1 2 1 40.28
-Co-Percep [83] 2020 10K $\times$ $\checkmark$ $\checkmark$ 1 1 1 1 1 40.06
-A9-Dataset [84] 2022 1,098 $\checkmark$ $\checkmark$ $\checkmark$ 9 4 2 1 1 39.12
-LUMPI [85] 2022 200K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 3 6 3 1 1 35.87
-DOLPHINS [86] 2022 42,376 $\checkmark$ $\checkmark$ $\checkmark$ 2 3 2 4 - 35.21
+Rope3D [88] 2022 50K $\times$ $\checkmark$ $\checkmark$ 12 3 3 2 1 48.96
+V2X-Sim [89] 2022 10K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23 1 1 1 - 48.50
+V2VNet [90] 2020 51,2K $\checkmark$ $\checkmark$ 1 1 1 1 - 48.20
+T&J [91] 2019 100 $\checkmark$ $\checkmark$ $\checkmark$ 1 1 1 2 1 46.63
+Co-Percep [92] 2020 10K $\times$ $\checkmark$ $\checkmark$ 1 1 1 1 1 42.99
+DeepAccident [12] 2023 285K $\checkmark$ 6 4 3 2 - 41.79
+LUMPI [93] 2022 200K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 3 6 3 1 1 40.42
 Drone
@@
 STCrowd [102] 2022 10,891 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 1 3 1 1 1 35.19
-TABLE II: High-impact perception datasets. For a more comprehensive demonstration, we exhibit 50 perception datasets from different sensing domains instead of those with the highest scores.
 V-A Perception Datasets
@@
 It is worth noting that the datasets in the table are sorted by impact score within their respective categories and do not represent the overall top 50.
-In the following part, we choose several datasets with the highest impact scores within each sensing source and exhibit them considering their published year.
+In the following section, we choose several datasets with the highest impact scores within each sensing source and exhibit them, considering their published year.
 V-A1 Onboard
 KITTI.
-KITTI [13] has profoundly influenced the autonomous driving domain since its release in 2012.
-KITTI encompasses a variety of real-world driving scenarios recorded by various sensors, including cameras, LiDAR, and GPS/IMU.
-Its rich annotations and high-resolution sensor data have facilitated the development and benchmarking of numerous algorithms across a range of autonomous driving tasks, such as object detection, tracking, optical flow, depth estimation, and visual odometry.
+KITTI [13] has been instrumental in advancing the autonomous driving domain since its release in 2012.
+KITTI data is recorded by various sensors, including cameras, LiDAR, and GPS/IMU, facilitating the development of algorithms for object detection, tracking, optical flow, depth estimation, and visual odometry.
+However, KITTI data was mainly recorded under ideal weather conditions in German urban areas.
+The relatively small geographic and environmental conditions scope limit its real-world applications.
 Cityscapes.
-Cityscapes [14] comprises a vast collection of images captured explicitly in intricate urban environments.
-Through meticulous labeling, Cityscapes provides pixel-level segmentation for 30 different object classes, which include various vehicle types, pedestrians, roads, and traffic sign information.
-Given its complexity and richness, Cityscapes has become a standard benchmark for tasks such as semantic segmentation in urban scenarios.
-SYNTHIA.
-SYNTHIA [2] is a synthetic dataset in the realm of autonomous driving.
-The dataset includes 13,400 images with pixel-wise annotations for semantic segmentation.
-One of the standout attributes of SYNTHIA is its capability to bridge the gap between real-world and synthetic data, fostering the development of robust and transferable approaches across different domains.
-Virtual KITTI.
-Virtual KITTI [3] crafts to closely mimic the original KITTI dataset through a virtual environment, distinguishing itself by delivering high-resolution video sequences encapsulating various traffic situations and environmental conditions.
-Similar to [13], Virtual KITTI supports crucial autonomous driving tasks, including object detection, semantic segmentation, and object tracking.
+Cityscapes [14] comprises a vast collection of images captured explicitly in intricate urban environments and has become a standard benchmark semantic segmentation.
+Cityscapes provides pixel-level segmentation through meticulous labeling for 30 object classes, including various vehicle types, pedestrians, roads, and traffic sign information.
+However, Cityscapes primarily focuses on German cities and lacks diverse weather conditions.
+These limitations hinder its usefulness for developing robust and generalized autonomous driving solutions.
 VIPER.
-VIPER [5] is a synthetic dataset collected from driving, riding, and walking perspectives in a realistic virtual world, addressing the challenge of data scarcity and the high cost of labeling real-world data.
-VIPER contains over 250,000 video frames annotated with ground-truth data for low- and high-level vision tasks while encapsulating various weather conditions, lighting scenarios, and complex urban landscapes.
-Overall, VIPER presents a valuable and cost-effective tool for researchers to expedite the development of reliable and safe autonomous driving.
-Apolloscapes.
-Apolloscapes [17] offers more than 140,000 high-resolution frames with accurate bounding boxes and pixel-wise semantic labels, critical for training and validating autonomous vehicles’ perception and navigation systems.
-Apolloscapes supports images and point clouds semantic segmentation, 2D/3D object detection, multi-object tracking, and lane segmentation, enabling the creation and evaluation of advanced and safe autonomous driving systems.
+VIPER [5] is a synthetic dataset collected based on driving, riding, and walking perspectives in a realistic virtual world.
+VIPER contains over 250K video frames annotated with ground-truth data for low- and high-level vision tasks.
+It encapsulates various weather conditions, lighting scenarios, and complex urban landscapes, making it an invaluable resource for testing the robustness of autonomous driving algorithms.
+Nonetheless, the transition from synthetic data to real-world application remains challenging, as algorithms must bridge the domain gap to maintain their effectiveness in real environments.
 SemanticKITTI.
-SemanticKITTI [6] is a notable extension of the KITTI family, focusing on semantic segmentation in the autonomous driving field.
-SemanticKITTI consists of over 43,000 LiDAR point cloud frames, making it one of the largest datasets for 3D semantic segmentation in outdoor environments.
-SemanticKITTI provides precise labels for 28 categories, such as car, road, building, etc., achieving a robust benchmark for evaluating the performance of point cloud semantic segmentation methods, underpinning numerous studies and innovations in related domains.
+SemanticKITTI [6] consists of over 43,000 LiDAR point cloud frames, making it one of the most extensive datasets for 3D semantic segmentation in outdoor environments.
+SemanticKITTI provides precise labels for 28 categories, such as car, road, building, etc., achieving a robust benchmark for semantic segmentation.
+However, similar to the previous datasets, SemanticKITTI faces limitations in real-world applicability due to its restricted environmental diversity and geographical scope.
 nuScenes.
-nuScenes [22] stands as an essential contribution to the field of autonomous driving, offering a rich repository of data that addresses the diverse needs of perception systems.
-nuScenes leverages LiDAR, radars, and cameras to record data from different urban scenes from Boston and Singapore.
-It is worth mentioning that its six cameras provide a comprehensive perspective of the surrounding environment, making them widely utilized in multi-view object detection tasks.
-In conclusion, the nuScenes dataset is a cornerstone in developing autonomous driving technologies, supporting multi-tasks and applications, and setting new benchmarks in the field.
+nuScenes [22] stands as an essential contribution to the field of autonomous driving, providing multimodal sensor setup, including LiDAR, radars, and cameras.
+It addresses the diversity in urban scenes and environmental conditions.
+The data recorded from Boston and Singapore cover varied driving behaviors and urban layouts, enhancing generalizability.
+Its six cameras provide a comprehensive perspective of the surrounding environment, making them widely utilized in multi-view object detection tasks.
+Yet, the dataset’s coverage of rare driving cases, such as accidents, could be expanded to better assess algorithm performance under challenging conditions.
 Waymo.
-The Waymo Open Dataset [23], introduced in 2019, significantly influences research and advancement in autonomous driving by providing an extensive size of multimodal sensory data with high-quality annotations.
-Key contributions of the Waymo dataset include its comprehensive coverage of driving conditions and geographics, which are pivotal for the robustness and generability of different tasks, such as detection, tracking, and segmentation.
+The Waymo Open Dataset [23], introduced in 2019, significantly influences research and advancement in autonomous driving.
+Waymo provides an extensive size of multimodal sensory data with high-quality annotations compared to others.
+Key contributions of the Waymo dataset include its comprehensive coverage of driving conditions and geographics, which are pivotal for the robustness and generability of different tasks.
+The real-world applicability of Waymo’s data is enhanced by its diversity, although exploring its limitations in specific adverse conditions could provide deeper insights into areas needing improvement.
 BDD100K.
-BDD100K [24] dataset, released by the Berkeley DeepDrive center in 2018, is a substantial and diverse driving dataset renowned for its size and diversity.
+BDD100K [24] dataset is renowned for its size and diversity.
 It comprises 100,000 videos, each about 40 seconds in duration.
-Meanwhile, it offers various annotated labels for object detection, tracking, semantic segmentation, and lane detection.
-This extensive compilation of data prompts advancements in the autonomous driving community, establishing itself as a challenging and versatile platform for researchers and engineers to propose and refine algorithms.
-RADIATE.
-RADIATE [26] is the first public radar dataset, which contains 44,140 annotated frames gathered under several adversarial weather conditions, such as rain, fog, overcast, and snow.
-It also incorporates LiDAR and camera data, allowing full perception and understanding of the driving surroundings.
+Furthermore, it includes different times of day and weather conditions, offering a solid foundation for testing and improving the robustness of algorithms.
+Meanwhile, it provides various annotated labels for object detection, tracking, semantic segmentation, and lane detection.
+However, the varying quality of video annotations challenges the dataset’s real-world applicability, highlighting the need for consistent, high-quality annotations to ensure algorithm effectiveness across all conditions.
 Argoverse 2.
@@
 Argoverse 2 supports a wide range of essential tasks, including but not limited to 3D object detection, semantic segmentation, and tracking.
-In summary, the Argoverse 2 dataset offers numerous multimodal data of real-world driving scenarios, fostering innovations and advancements of algorithms and showing its substantial potential as a vital resource in autonomous driving.
+The limitation of Argoverse 2 in the real world is similar to [22] and [23], considering more adversarial conditions and edge cases can extend its application field.
 V-A2 V2X
-V2VNet.
-The dataset introduced by V2VNet [78] focuses on leveraging V2V communication to allow autonomous vehicles to share information and perceive the environment from multiple viewpoints, which is crucial for detecting occluded objects and predicting the behavior of other traffic participants.
-This dataset was created using a high-fidelity LiDAR simulator named Lidarsim [96] that utilizes real-world data to generate realistic LiDAR point clouds for various traffic scenes.
-In conclusion, this work brings attention to the V2V as a promising avenue for improving the capabilities of autonomous vehicles.
+TUMTraf.
+The TUM Traffic Dataset family (TUMTraf) is a cutting-edge real-world dataset comprising 50,253 labeled frames (9,545 point clouds and 40,708 images) across five releases, capturing diverse traffic scenarios in Munich, Germany [84, 85, 86, 60].
+It integrates multiple data modalities like RGB camera, event-based camera, LiDAR, GPS, and IMU.
+TUMTraf also provides perspectives from the infrastructure and vehicle, enabling research in cooperative perception.
+TUMTraf stands out for including edge cases like accidents, near-miss events, and traffic violations, providing a great resource to improve perception systems.
 DAIR-V2X.
@@
 The dataset is designed to tackle challenges such as the temporal asynchrony between vehicle and infrastructure sensors and the data transmission costs involved in such cooperative systems.
-The impact of the DAIR-V2X dataset on autonomous driving is significant because it sets a benchmark for complexities of vehicle-infrastructure cooperation thanks to its multiple scenarios from the real world.
-Rope3D.
-Rope3D presented in [80] is an essential contribution to the perception systems, which addresses the critical gap in autonomous driving by leveraging data gathered from roadside cameras.
-Rope3D consists of 50,000 images under various environmental conditions, including different illumination (daytime, night, dusk) and weather scenarios (rainy, sunny, cloudy), ensuring a high diversity of data that reflects real-world complexities.
-Overall, the Rope3D dataset is a pioneering work that accelerates advancements in roadside perception for autonomous driving while being a critical tool for researchers and engineers to develop more robust and intelligent autonomous driving systems.
-V2V4Real.
-V2V4Real [82] is the first large-scale, real-world dataset to address V2V cooperative perception.
-The data was collected from two vehicles with multimodal sensors like LiDAR and cameras.
-V2V4Real focuses on a range of perception tasks, such as cooperation 3D object detection, cooperative 3D object tracking, and Sim2Real domain adaptation.
-This versatility makes it an invaluable resource for developing and benchmarking autonomous driving algorithms.
+DAIR-V2X has been instrumental in advancing vehicle-infrastructure cooperation, setting benchmarks for addressing V2X perception tasks.
 V-A3 Drone
@@
 Pasacal3D+ [97] is an extension of the PASCAL VOC 2022 [103], overcoming the limitations of previous datasets by providing a richer and more varied set of annotations for images.
-Pasacal3D+ augments 12 rigid object categories, such as cars, buses, and bicycles, with 3D pose annotations and adds more images from ImageNet [98], resulting in a high degree of variability.
-TT 100K.
-Tsinghua-Tencent 100K [15] solves the challenges of detecting and classifying traffic signs in realistic driving conditions.
-It provides 100K images, including 30,000 traffic-sign instances.
-Beyond its large data size, the high-resolution images encompass a wide range of illumination and weather conditions, making it robust for training and validating traffic sign recognition.
+Pasacal3D+ augments 12 rigid object categories, such as cars, buses, and bicycles, with 3D pose annotations and adds more images from ImageNet [104], resulting in high variability.
+However, Pascal3D+ only focuses on rigid objects and may not fully address the dynamic driving environments where pedestrians and other non-rigid objects are present.
 Mapillary Vistas.
@@
 V-B Prediction, Planning, and Control Datasets
-Prediction, planning, and control datasets serve as the foundation for facilitating the training and evaluation of driving systems to forecast traffic dynamics, pedestrian movements, and other essential factors that influence driving decisions.
-By simulating myriad driving scenarios, they empower autonomous vehicles to make informed decisions, navigate complex environments, and maintain safety and efficiency on the road.
+Prediction, planning, and control datasets serve as the foundation for facilitating the development of driving systems.
+These datasets are critical for forecasting traffic dynamics, pedestrian movements, and other essential factors that influence driving decisions.
 Hence, we demonstrate in detail several high-impact datasets related to these tasks according to the data size, modalities, and citation number.
-We summarize the prediction, planning, and control datasets into task-specific and multi-task groups.
+We summarize these datasets into task-specific and multi-task groups.
+TABLE III: Prediction, planning, and control datasets.
+We demonstrate several crucial datasets related to prediction, planning, and control.
+BP: behavior prediction, DAP: driver’s attention prediction, IP: intention prediction, MP: motion prediction, TP: trajectory prediction, MPlan: motion planning, DM: decision-making, OT: object tracking, QA: question-answering, DBP: driver behavior recognition
 Dataset Year Sensing domain Size Tasks Weather conditions Time of day Scenario conditions
 Task-Specific
+Brain4Cars [105] 2015 others 2M frames Maneuver Anticipation
 JAAD [16] 2017 onboard 75K frames pedestrian IT sunny, rainy, cloudy, snowy day, afternoon, night urban
-Dr(eye)ve [53] 2018 onboard 500K frames driver’s attention prediction sunny, rainy, cloudy day, night urban, countryside, highway
+Dr(eye)ve [59] 2018 onboard 500K frames DAP sunny, rainy, cloudy day, night urban, countryside, highway
 highD [18] 2018 drone 45K km distance TP sunny 8 am to 5 pm highway
@@
 Argoverse [19] 2019 onboard 300K trajectories TP variety variety urban
+Drive&\&&Act [107] 2019 others 9.6M images DBR
+DbNet [108] 2019 onboard 100 km distance DBR various road types
+D2CAV [109] 2020 onboard behavioral strategy
 inD [110] 2020 drone 11.5K trajectories road user prediction sunny day 4 urban intersections
@@
 nuPlan [113] 2021 drone 1.5K hours data MPlan 4 cities
+DriPE [114] 2021 others 10K pictures DBR daytime
+Speak2label [115] 2021 others 586 videos DAP sunny, cloudy daytime
+CoCAtt [116] 2022 onboard 11.9 hours DAP countryside
 exiD [117] 2022 drone 16 hours data TP sunny daytime 7 locations on highway
 MONA [118] 2022 drone 702K trajectories TP sunny, overcast, rain 8 am to 5 pm urban
+Occ3D-nuScenes [119] 2024 onboard 40K frames occupancy prediction
+Occ3D-Waymo [119] 2024 onboard 200K frames occupancy prediction
 Multi-Task
+HDD [120] 2018 onboard 104 hours data driver behavior, causal reasoning suburban, urban, highway
 INTERACTION [20] 2019 drone, V2X 110K trajectories MPlan and MP, DM (un)signalized intersection
@@
 rounD [122] 2019 drone 13,746 road users scenario classification, BP sunny daytime (sub-)urban
+PREVENTION [123] 2019 onboard 356 mins video IP, TP highway, urban areas
+DrivingStereo [124] 2019 onboard 180K images trajectory planning sunny, rainy, cloudy daytime suburban, urban, highway
+foggy, dusky elevated road, country road
 Lyft Level 5 [125] 2021 drone 1.1K hours data MPlan, MP suburban
-LOKI [109] 2021 onboard 644 scenarios TP, BP (sub-)urban
+LOKI [126] 2021 onboard 644 scenarios TP, BP variety variety (sub-)urban
 SceNDD [127] 2022 onboard 68 driving scenes MPlan, MP urban
 DeepAccident [12] 2023 V2X 57K frames MP, accident prediction sunny, rainy, cloudy, wet noon, sunset, night synthetic
-Talk2BEV [111] 2023 onboard 20K QA pairs DM, IT, spatial reasoning
 V2X-Seq (forecasting) [56] 2023 V2X 50K scenarios online/offline VIC TP 28 urban intersections
-TABLE III: Prediction, planning, and control datasets.
-We demonstrate several crucial datasets related to prediction, planning, and control.
-BP: behavior prediction, IP: intention prediction, MP: motion prediction, TP: trajectory prediction, MPlan: motion planning, DM: decision-making, OT: object tracking, QA: question-answering
 V-B1 Task-Specific Datasets
 highD.
-The drone-based highD [18] dataset provides a large-scale collection of naturalistic vehicle trajectories on German highways, containing post-processed trajectories of 110,000 cars and trucks.
-The dataset aims to overcome the limitations of existing measurement methods for scenario-based safety validation, which often fail to capture the naturalistic behavior of road users or to include all relevant data with sufficient quality.
+The drone-based highD [18] dataset provides a large-scale collection of naturalistic vehicle trajectories on German highways, containing post-processed trajectories of 110K cars and trucks.
+It is designed to address the shortcomings of traditional measurement techniques in scenario-based safety validation, which often miss capturing authentic road user behaviors or lack comprehensive, high-quality data.
+But highD is recorded under ideal weather conditions, limiting its application under adversarial weather conditions.
 PIE.
@@
 It encompasses over 6 hours of driving footage recorded in downtown Toronto under various lighting conditions.
-The PIE dataset offers rich annotations for perception and visual reasoning, including bounding boxes with occlusion flags, crossing intention confidence, and text labels for pedestrian actions.
-The long continuous sequences and annotations facilitate multiple tasks like trajectory prediction and pedestrian intention prediction.
-USyd.
-USyd [99] pushes the progress of the driver intention prediction in the context of urban intersections without traffic signals, which are common in urban settings and represent a challenge due to the lack of clear road rules and signals.
-The dataset incorporates over 23,000 vehicles traversing five different intersections, collected using a vehicle-mounted LiDAR-based tracking system.
-The data modalities included are exhaustive, providing vehicle tracks with lateral and longitudinal coordinates, heading, and velocity.
-This information is vital to predict driver behavior, accounting for the uncertainty inherent in human driving patterns.
+PIE offers rich annotations for perception and visual reasoning, including bounding boxes with occlusion flags, crossing intention confidence, and text labels for pedestrian actions.
 Argoverse.
 Argoverse [19] is a crucial dataset in 3D object tracking and motion forecasting.
-Argoverse provides ${}^{\circ}$ images from 7 cameras, forward-facing stereo imagery, and LiDAR point clouds.
+Argoverse provides 360 images from 7 cameras, forward-facing stereo imagery, and LiDAR point clouds.
 The recorded data covers over 300K extracted vehicle trajectories from 290km of mapped lanes.
 With the assistance of rich sensor data and semantic maps, Argoverse is pivotal in advancing research and development in prediction systems.
-inD.
-The importance of the inD [100] lies in its large-scale, high-quality, and diverse set of trajectory data crucial for several applications, including road user prediction models and scenarios-based safety validation for autonomous vehicles in urban intersection environments.
-It covers around 11,500 different road user trajectories, e.g., vehicles, bicyclists, and pedestrians.
-These trajectories have a positioning error of less than 0.1 meters, which is pivotal for the reliability of the data.
-PePscenes.
-PePscenes [101] tackles the requirement for understanding and anticipating pedestrian actions in dynamic driving surroundings.
-This dataset enhances the nuScenes [22] dataset by adding per-frame 2D/3D bounding box and behavior annotations, focusing on pedestrian crossing actions.
-One of the key attributes of [101] is incorporating various data types involving semantic maps, scene images, trajectories, and ego-vehicle states, which are necessary for creating robust models capable of understanding complex traffic scenarios.
-openDD.
-The openDD [102] dataset focuses on analyzing and predicting traffic scenes around roundabouts, which are complex and unregulated by traffic lights.
-It is unique in leveraging drone-captured imagery with high resolution (4K), which spans over 62 hours of trajectory data from 501 separate flights.
-The dataset contains not only trajectories but also shapefiles and an extensible markup language (XML) file describing the road topology, along with a reference image for each underlying intersection.
+Real-world implementation has verified the effectiveness of Argoverse in urban environments, though its performance in different geographical areas may be constrained.
 nuPlan.
 nuPlan [113] is the world’s first closed-loop machine learning-based planning benchmark in autonomous driving.
-This multimodal dataset comprises around 1,500 hours of human driving data from four cities across America and Asia, featuring different traffic patterns, such as merges, lane changes, interactions with cyclists and pedestrians, and driving in construction zones.
+This multimodal dataset comprises around 1,500 hours of human driving data from four cities across America and Asia.
+It features different traffic patterns, such as merges, lane changes, interactions with cyclists and pedestrians, and driving in construction zones.
 These characters of the nuPlan dataset take into account the dynamic and interactive nature of actual driving, allowing for a more realistic evaluation.
+Its real-world impact is seen in developing more adaptive and context-aware planning systems.
 exiD.
@@
 It takes advantage of drones to record traffic without occlusion, minimizing the influence on traffic and ensuring high data quality and efficiency.
-This drone-based dataset surpasses previous datasets in terms of the diversity of interactions captured, especially those involving lane changes at highway entries and exits.
+This drone-based dataset surpasses previous datasets regarding the diversity of interactions captured, especially those involving lane changes at highway entries and exits.
+Extending the data from weather conditions and nighttime could be an improvement direction for this dataset.
 MONA.
@@
 This dataset boasts an average overall position accuracy of 0.51 meters, which exhibits the quality of the data collected using highly accurate localization and LiDAR sensors.
+However, only recording the data in a particular city may limit its generalizability to other geographic locations.
 V-B2 Multi-Task Datasets
 INTERACTION.
-The INTERACTION [20] dataset covers diverse, complex, and critical driving scenes, coupled with its comprehensive semantic map, allowing it to be a versatile platform for a multitude of tasks, such as motion prediction, imitation learning, and validation of decision and planning.
-Its inclusion of different countries further improves the robustness of analyzing the driving behavior across different cultures, which is critical for the global development of autonomous driving.
-BLVD.
-The BLVD [106] benchmark facilitates tasks such as dynamic 4D (3D+temporal) tracking, 5D (4D+interactive) interactive event recognition, and intention prediction, which are essential for a deeper understanding of traffic scenes.
-BLVD offers around 120K frames from different traffic scenes, comprising object density (low and high) and illumination conditions (daytime and nighttime).
-These frames are fully annotated with a large number of 3D labels incorporating vehicles, pedestrians, and riders.
+The INTERACTION [20] dataset is a versatile platform that offers diverse, complex, and critical driving scenes, along with a comprehensive semantic map, making it suitable for a multitude of tasks, such as motion prediction, imitation learning, and validation of decision and planning.
+Its inclusion of different countries and continents further improves the robustness of analyzing the driving behavior across different cultures.
+A potential shortcoming of the dataset is that the impact of environmental conditions has not been explicitly addressed.
 rounD.
-The rounD dataset presented by [107] is pivotal for scenario classification, road user behavior prediction, and driver modeling because of the large number of collections of road user trajectories at roundabouts.
+The rounD dataset presented by [122] is pivotal for scenario classification, road user behavior prediction, and driver modeling.
+It provides a large number of road user trajectories at roundabouts.
 The dataset utilizes a drone equipped with a 4K resolution camera to collect over six hours of video, recording more than 13K road users.
 The broad recorded traffic situations and the high-quality recordings make rounD an essential dataset in autonomous driving, facilitating the study of naturalistic driving behaviors in public traffic.
+Yet, similar to all datasets only collected under clear weather conditions, developing models capable of performing under challenging situations could be restricted.
 Lyft Level 5.
@@
 It supports multiple tasks like motion forecasting, motion planning, and simulation.
-The numerous multimodal data with detailed annotations make the Lyft Level 5 dataset a vital benchmark for prediction and planning.
+The numerous multimodal data with detailed annotations make it a vital benchmark for prediction and planning.
+The dataset shows limitations in accurately representing scenes with uncommon traffic conditions or rare pedestrian behaviors.
 LOKI.
@@
 This dataset makes a multidimensional view of traffic scenarios available by utilizing camera images with corresponding LiDAR point clouds, making it a highly flexible resource for the community.
-SceNDD.
-SceNDD [110] introduces real driving scenarios that feature diverse trajectories and driving behaviors for developing efficient motion planning and path-following algorithms.
-It is also adaptable to different configurations of the ego-vehicle and contains a time horizon of prediction that can be broken down into timestamps for detailed analysis.
-In conclusion, the SceNDD dataset is a significant addition to autonomous driving prediction and planning research.
 DeepAccident.
-The synthetic dataset, DeepAccident [12], is the first work that provides direct and explainable safety evaluation metrics for autonomous vehicles.
-The extensive dataset with 57K annotated frames and 285K annotated samples supports end-to-end motion and accident prediction, which is vital for improving the predictive capabilities of autonomous driving systems in avoiding collisions and ensuring safety.
-Moreover, this multimodal dataset is versatile for various V2X-based perception tasks, such as 3D object detection, tracking, and bird’s-eye-view (BEV) semantic segmentation.
-Talk2Bev.
-The innovative dataset, Talk2BEV [111], promotes the shift from traditional autonomous driving tasks to incorporate large vision-language models with BEV maps in the context of autonomous driving.
-Talk2BEV utilizes recent advances in vision-language models, allowing for a more flexible and comprehensive understanding of road scenarios.
-This dataset comprises over 20,000 diverse question categories, all human-annotated and derived from [22].
-The proposed Talk2BEV-Bench benchmark can be leveraged across multiple tasks, covering decision-making, visual and spatial reasoning, and intent prediction.
-V2X-Seq (Forecasting).
-The trajectory forecasting dataset is a substantial part of the real-world dataset V2X-Seq [50], containing about 80K infrastructure-view and 80K vehicle-view scenarios, and a further 50K cooperative-view scenarios.
-This diversity of sensing domains creates a more holistic view of the traffic environment, exhibiting huge potentiality for research and analysis on vehicle-infrastructure-cooperative (VIC) trajectory prediction.
+The synthetic dataset, DeepAccident [12] is the first work that provides direct and explainable safety evaluation metrics for autonomous vehicles.
+The extensive dataset with 57K annotated frames and 285K annotated samples supports end-to-end motion and accident prediction, which is vital in avoiding collisions and ensuring safety.
+Moreover, this multimodal dataset is versatile for various V2X-based perception tasks, such as 3D object detection, tracking, and BEV semantic segmentation.
+The various environmental conditions also enhance the generalizability of this dataset.
+Due to the domain adaptation issue, the performance of models trained on DeepAccident in real driving situations needs further study.
 V-C End-to-end Datasets
@@
 DDD17.
-The DDD17 [113] dataset is notable for its use of event-based cameras, which provide a concurrent stream of standard active pixel sensor (APS) images and dynamic vision sensors (DVS) temporal contrast events, offering a unique blend of visual data.
+The DDD17 [129] dataset is notable for using event-based cameras.
+The dataset provides a concurrent stream of standard active pixel sensor (APS) images and dynamic vision sensors (DVS) temporal contrast events, offering a unique blend of visual data.
 Additionally, DDD17 captures diverse driving scenarios, including highway and city driving and various weather conditions, thus providing exhaustive and realistic data for training and testing end-to-end autonomous driving algorithms.
-Other datasets summarized in this survey are shown in Tab. IV Tab. V, Tab. VI.
-Dataset Year Size Temporal Sensing domain Tasks Real/Synthetic
-KAIST MPD [114] 2015 95K color-thermal pair frames $\times$ onboard pedestrian detection real
-BelgiumTS [115] 2011 13K traffic sign annotations $\times$ onboard traffic sign detection real
-nighttime drive [116] 2018 35K $\times$ onboard SS real
-[…]
-Caltech Lanes [118] 2019 1,224 frames $\times$ onboard lane detection real
-A*3D [119] 2020 39,179 point cloud frames $\times$ onboard (3D) OD real
-PreSIL [8] 2019 50K frames $\checkmark$ onboard OD, 3D SS synthetic
-H3D [120] 2019 27,721 frames $\checkmark$ onboard (3D) OD, MOT real
-ROAD [121] 2021 122K frames $\checkmark$ onboard OD, SS real
-ETH Ped [122] 2007 2,293 frames $\checkmark$ onboard pedestrian detection real
-CADC [123] 2020 56K $\times$ onboard (3D) OD, OT real
-RadarScenes [124] 2021 40K radar frames $\checkmark$ onboard OD, classification real
-All-in-One Drive [125] 2021 100K $\checkmark$ onboard (3D) OD, (3D) SS, trajectory prediction real
-CURE-TSR [126] 2017 2.2M annotated images $\times$ onboard traffic sign detection real
-Paris-Lille-3D [127] 2018 2,479 frames $\times$ onboard 3D SS, classification real
-PandaSet [128] 2021 8,240 frames $\checkmark$ onboard (3D) OD, SS, OT real
-NightOwis [129] 2018 56K frames $\checkmark$ onboard pedestrian detection, tracking real
-SynWoodScape [130] 2022 80K frames $\times$ onboard (3D) OD, segmentation synthetic
-Boreas [131] 2023 7,111 frames $\checkmark$ onboard (3D) OD, localization real
-TUD-Brussels [132] 2009 1,600 frames $\times$ onboard (3D) OD real
-VEIS [133] 2018 61,305 frames $\times$ onboard OD, SS synthetic
-CCTSDB 2021 [134] 2021 16,356 frames $\times$ onboard traffic sign detection real
-SemanticPOSS [135] 2020 2,988 point cloud frames $\checkmark$ onboard 3D SS synthetic
-IDDA [136] 2020 1M frames $\times$ onboard segmentation synthetic
-TJ4RadSet [137] 2022 7,757 frames $\checkmark$ onboard OD, OT real
-TME Motorway [138] 2012 30K frames $\checkmark$ onboard OD, OT real
-Stanford Tack [139] 2011 14K tracks $\checkmark$ onboard classification real
-CARRADA [140] 2020 7,193 radar frames $\checkmark$ onboard SS real
-SODA10M [141] 2021 20K labeled images $\checkmark$ onboard OD real
-Robo3D [142] 2023 476K frames $\checkmark$ onboard (3D) OD, 3D SS real
-LostAndFound [143] 2016 2,104 frames $\checkmark$ onboard obstacle detection real
-Titan [144] 2020 75,262 frames $\checkmark$ onboard OD, action recognition real
-CODA [145] 2022 1,500 frames $\times$ onboard corner case detection real
-PixSet [146] 2021 29K point cloud frames $\times$ onboard (3D) OD real
-ZOD [30] 2023 100K frames $\checkmark$ onboard (3D) OD, segmentation real
-K-Radar [147] 2023 35K radar frames $\times$ onboard 3D OD, OT real
-RoadObstacle21 [148] 2021 327 scenes $\times$ onboard anomaly segmentation synthetic
-VIL-100 [149] 2021 10K frames $\times$ onboard lane detection real
-EuroCity Persons [150] 2018 47,300 frames $\times$ onboard OD real
-TuSimple [151] 2017 6,408 frames $\times$ onboard lane detection, velocity estimation real
-OpenMPD [152] 2021 15K frames $\times$ onboard (3D) OD, 3D OT, semantic segmentation real
-WADS [153] 2021 1K point cloud frames $\checkmark$ onboard SS real
-NightCity [154] 2020 4,297 frames $\times$ onboard nighttime SS real
-LiDAR Snowfall [155] 2022 7,385 point cloud frames $\checkmark$ onboard 3D OD synthetic
-LLMAS [156] 2019 100K frames $\checkmark$ onboard lane detection real
-Fishyscapes [157] 2021 1,030 frames $\times$ onboard SS, anomaly detection real
-Toronto-3D [158] 2020 4 scenarios $\times$ onboard 3D SS real
-MIT-AVT [159] 2020 1.15M 10s video clips $\checkmark$ onboard SS, anomaly detection real
-LISA [160] 2014 6,610 frames $\checkmark$ onboard traffic sign detection real
-SynthCity [161] 2019 367.9M points in 30 scans $\times$ onboard (3D) OD, (3D) SS synthetic
-SemanticUSL [162] 2021 1.2K frames $\times$ onboard domain adaptation 3D SS real
-RANUS [163] 2018 4K frames $\times$ onboard SS, scene understanding real
-PePScenes [101] 2020 40K frames $\checkmark$ onboard (3D) OD, pedestrian action prediction real
-MUAD [164] 2022 10.4K frames $\checkmark$ onboard OD. SS, depth estimation synthetic
-AUTOCASTSIM [165] 2022 52K frames $\checkmark$ V2X (3D) OD, OT, SS real
-Mcity [166] 2019 1,7500 frames $\checkmark$ onboard SS real
-comma2k19 [167] 2018 2M images $\times$ onboard pose estimation, end-to-end driving real
-DET [168] 2019 5,424 event-based camera images $\times$ onboard lane detection real
-TRoM [169] 2017 712 frames $\times$ onboard road marking detection real
-Cirrus [170] 2020 6,285 frames $\checkmark$ onboard (3D) OD real
-KITTI InstanceMotSeg [171] 2020 12,919 frames $\checkmark$ onboard moving instance segmentation real
-CARTI [172] 2022 11K frames $\checkmark$ V2X cooperative perception synthetic
-Multifog KITTI [173] 2021 15K frames $\times$ onboard 3D OD synthetic
-UNDD [174] 2019 7.2K frames $\checkmark$ onboard SS real
-CeyMo [175] 2021 2,887 frames $\times$ onboard road marking detection real
-RaidaR [176] 2021 58,542 rainy street scenes $\times$ onboard SS real
-K-Lane [177] 2022 15,382 frames $\times$ onboard lane detection real
-aiMotive [178] 2023 26,583 frames $\checkmark$ onboard 3D OD, MOT real
-SAP [179] 2016 19K frames $\times$ drone OD, OT real
-Astyx [180] 2019 500 radar frames $\times$ onboard 3D OD real
-Comap [181] 2021 4,391 frames $\checkmark$ V2X 3D OD synthetic
-Ithaca365 [182] 2022 7K frames $\times$ onboard 3D OD, SS, depth estimation real
-Small Obstacles [183] 2020 3K frames $\times$ onboard small obstacle segmentation real
-GLARE [184] 2022 2,157 frames $\times$ onboard traffic sign detection real
-WIBAM [185] 2021 33,092 frames $\checkmark$ V2X 3D OD real
-MIT DriveSeg [186] 2019 5K frames $\times$ onboard SS real
-SUPS [187] 2023 5K frames $\checkmark$ onboard SS, depth estimation, SLAM synthetic
-R3 [188] 2021 369 scenes $\times$ onboard out-of-distribution detection real
-TABLE IV: Autonomous driving dataset 1
-Dataset Year Size Temporal Sensing domain Tasks Real/Synthetic
-V2X-Seq (perception) [50] 2023 15K frames $\checkmark$ V2X cooperative perception real
-DRIV100 [189] 2021 100K frames $\times$ onboard domain adaptation SS real
-NEOLIX [190] 2021 30K frames $\checkmark$ onboard 3D OD, OT real
-CARLANE [191] 2022 118K frames $\checkmark$ onboard lane detection synthetic
-IPS3000+ [192] 2021 14,198 frames $\checkmark$ V2X 3D OD real
-TUM-Traffic [193] 2023 4.8K frames $\checkmark$ V2X 3D OD, OT real
-Amodal Cityscapes [194] 2022 5K frames $\times$ onboard amodal SS real
-WEDGE [195] 2023 3,360 frames $\times$ others OD, classification synthetic
-HEV [196] 2019 230 video clips $\checkmark$ onboard object localization real
-Aachen Day-Night [197] 2018 4,328 images, 1.65M points $\checkmark$ onboard visual localization real
-HAD [198] 2019 5,675 video clips $\checkmark$ onboard end-to-end driving real
-CARLA-100 [199] 2019 100 hours driving $\checkmark$ onboard path planning, behavior cloning synthetic
-Collective activity [200] 2009 44 short videos $\checkmark$ others human activity recognition real
-DA4AD [201] 2020 9 sequences $\checkmark$ onboard visual localization real
-Bosch STL [202] 2017 13,334 images $\checkmark$ onboard traffic light detection and classification real
-UrbanLaneGraph [203] 2023 around 5,220 km lane spans $\checkmark$ drone lane graph estimation real
-CrashD [204] 2022 15,340 scenes $\times$ onboard 3D OD synthetic
-CODD [205] 2022 108 sequences $\checkmark$ V2X multi-agent SLAM synthetic
-MIT DSS [186] 2020 10K video frames $\checkmark$ onboard SS real
-CPIS [83] 2020 10K frames $\times$ V2X cooperative 3D OD synthetic
-CarlaScenes [206] 2022 7 sequences $\checkmark$ onboard (3D) SS, SLAM, depth estimation synthetic
-FlyingThings3D [207] 2015 26,066 frames $\checkmark$ others scene flow estimation synthetic
-San Francisco Landmark [208] 2011 150K panoramic images $\times$ others landmark identification real
-DDAD [209] 2020 21,200 frames $\times$ onboard depth estimation real
-Ua-detrac [210] 2015 140K frames $\checkmark$ V2X OD, MOT real
-NCLT [211] 2015 34.9 hours $\checkmark$ onboard odometry real
-NVSEC [212] 2018 around 28 km distances $\checkmark$ others SLAM real
-MSLU [213] 2013 36.8 km distances $\checkmark$ onboard SLAM real
-Oxford Radar RobotCar [214] 2020 240K scans $\checkmark$ onboard odometry real
-A2D2 [25] 2020 433,833 frames $\checkmark$ onboard 3D OD, SS, SLAM real
-VERI-Wild [215] 2019 416,314 images $\checkmark$ V2X onboard re-identification real
-OTOH [108] 2020 170K scenes $\checkmark$ drone trajectory prediction, planning real
-TRANCOS [216] 2015 1.2K images $\times$ V2X onboard number estimation real
-Complex Urban [217] 2017 around 190km paths onboard SLAM real
-Syncapes [218] 2018 25K images $\times$ onboard OD, SS synthetic
-SydneyUrbanObject [219] 2013 588 object scans $\times$ onboard classification real
-ApolloCar3D [220] 2019 5,277 driving images $\times$ onboard 3D instance understanding real
-ACDC [76] 2021 4,006 images $\times$ onboard SS on adverse conditions real
-MulRan [221] 2020 41.2km paths $\checkmark$ onboard place recognition real
-Paris-rue-Madame [222] 2014 643 objects $\times$ onboard OD, SS real
-Ground Truth SitXel [223] 2013 78,500 frames $\checkmark$ onboard stereo confidence real
-LiVi-Set [224] 2018 10K frames $\checkmark$ onboard driving behavior prediction real
-Newer College [225] 2020 290M points, 2300 seconds $\checkmark$ others SLAM real
-CADP [226] 2018 1,416 scenes $\times$ V2X traffic accident analysis real
-LIBRE [227] 2020 40 frames $\times$ onboard LiDAR performance benchmark real
-OPV2V [228] 2022 11,464 frames $\times$ V2X onboard-to-onboard perception synthetic
-NYC3DCars [229] 2013 2K images $\times$ onboard OD real
-RUGD [230] 2019 37K frames $\checkmark$ others SS real
-EU LTD [231] 2020 around 37 hours $\checkmark$ onboard odometry real
-MOTSynth [232] 2021 768 driving sequences $\checkmark$ onboard Pedestrian detection and tracking synthetic
-LLamas [156] 2019 100,042 images $\times$ onboard lane marker detection, lane segmentation real
-PedX [233] 2018 5K images $\checkmark$ onboard pedestrian detection and tracking real
-CCD [234] 2020 4.5K videos $\checkmark$ onboard accident prediction real
-MAVD [235] 2021 113,283 images $\checkmark$ onboard OD and OT with sound real
-Gated2Depth [236] 2020 17,686 frames $\checkmark$ onboard depth estimation real
-DDD20 [237] 2017 51 hours event frames $\checkmark$ onboard end-to-end driving real
-TCGR [238] 2020 839,350 frames $\checkmark$ others traffic control gesture recognition real
- 4Seasons [238] 2020 350km recordings $\checkmark$ onboard SLAM real
-CCSAD [239] 2015 96K frames $\checkmark$ onboard scene understanding real
-TAF-BW [240] 2018 2 scenarios $\checkmark$ V2X MOT, V2X communication real
-Boxy [241] 2019 200K frames $\checkmark$ onboard OD real
-AMUSE [242] 2013 117,440 frames $\checkmark$ onboard SLAM real
-Brno Urban [243] 2019 375.7 km $\checkmark$ onboard recognition real
-AUTOMATUM [244] 2021 30 hours $\checkmark$ drone trajectory prediction real
-DurLAR [245] 2021 100K frames $\checkmark$ onboard depth estimation real
-Reasonable-Crowd [246] 2021 92 scenarios $\checkmark$ onboard driving behavior prediction synthetic
-Daimler Ped [247] 2013 12,485 frames $\checkmark$ onboard pedestrian path prediction real
-PVDN [248] 2021 59,746 frames $\checkmark$ onboard nighttime OD, OT real
-CARLA-WildLife [249] 2022 26 videos $\checkmark$ onboard out-of-distribution tracking synthetic
-SOS [249] 2022 20 videos $\checkmark$ onboard out-of-distribution tracking real
-RoadSaW [250] 2022 720K scenes $\checkmark$ onboard Road surface and wetness estimation real
-I see you [251] 2022 170 sequences, 340 trajectories $\checkmark$ V2X OD real
-ASAP [252] 2022 1.2M images $\checkmark$ onboard online 3D OD real
-OpenLane-V2 [253] 2023 466K images $\checkmark$ onboard lane detection, scene understanding real
-Daimler Stereo Ped [254] 2011 21,790 frames $\checkmark$ onboard pedestrian detection real
-Brain4Cars [255] 2015 2M frames $\checkmark$ onboard Maneuver Anticipation real
-NEXET [256] 2017 91,190 frames $\times$ onboard OD real
-DIML [257] 2017 470 videos $\checkmark$ onboard lane detection real
-TABLE V: Autonomous driving dataset 2
-Dataset Year Size Temporal Sensing domain Tasks Real/Synthetic
-UAH-Driveset [258] 2016 500 mins $\checkmark$ onboard lane detection, detection real
-SSCBENCH [259] 2023 66,913 frames $\times$ onboard semantic scene completion real
-FLIR [260] - 26,442 thermal frames $\checkmark$ onboard thermal image OD real
-VLMV [261] 2020 900 frames $\checkmark$ V2X lane merge real
-CityFlow [262] 2019 200K bounding boxes $\checkmark$ V2X OD, MOT, re-identification real
-PREVENTION [263] 2019 356 mins $\checkmark$ onboard onboard trajectory and intention prediction real
-TABLE VI: Autonomous driving dataset 3
 VI Annotations process
@@
 Additionally, we analyze the most important aspects for ensuring annotation quality VI-B.
-VI-A How annotations are created
+VI-A Annotation Generation
 Different AD tasks require specific types of annotation.
@@
 11, the annotation pipeline can be categorized into three types: manual annotation, semi-automatic annotation, and fully automatic annotation.
-We detail the labeling approaches for different types of annotation in this section.
-Annotate Segmentation Data.
-The target of annotating segmentation data is to assign a label to each pixel in an image or each point in a LiDAR frame to indicate which object or region it belongs to.
-After labeling, all pixels belonging to an object are annotated with the same class.
-For the manual annotation process, the annotator first draws boundaries around an object and then fills in the area or paints over the pixels directly.
-However, generating pixel/point-level annotations in this way is costly and inefficient.
+In this section, we detail the labeling approaches for different annotation types.
 Figure 11: Annotation pipelines.
@@
 (c) Fully automatic annotation: The framework annotates data without human correction.
-Many studies have proposed fully or semi-automatic annotation methods to improve annotation efficiency.
-[264] presented a fully automatic annotation approach based on weakly supervised learning to segment proposed drivable paths in images.
-[265] is a semi-automatic annotation method utilizing the objectness priors to generate segmentation masks.
-After that, [266] offered a semi-automatic method considering 20 classes.
-Polygon-RNN++[267] presented an interactive segmentation annotation tool following the idea of [268].
-Instead of using image information to generate pixel-level labels, [269] explored transferring 3D information into 2D image domains to generate semantic segmentation annotations.
-For labeling 3D data, [270] proposed an image-assisted annotation pipeline.
-[271] leveraged active learning to select a few points and to form a minimal training set to avoid labeling the whole point cloud scenes.
-[272] introduced an efficient labeling framework with semi/weakly supervised learning to label outdoor point clouds.
 Annotate 2D/3D Bounding Boxes.
@@
 Wang et al.
-[274] presented a semi-automatic video labeling tool based on the open-source video annotation system VATIC [^7] https://www.cs.columbia.edu/~vondrick/vatic/.
-[275] is another automatic video annotation tool for AD scenes.
-Compared to daytime annotation, process bounding box annotations in the nighttime is more challenging.
+[131] presented a semi-automatic video labeling tool based on the open-source video annotation system VATIC[132].
+Manikandan et al.
+[133] propose another automatic video annotation tool for AD scenes.
+Process bounding box annotations in the nighttime is more challenging than daytime annotation.
+Schorkhuber et al.
 [134] introduced a semi-automatic approach leveraging the trajectory to solve this problem.
@@
 3D BAT [138] developed an annotation toolbox to assist in obtaining 2D and 3D labels in semi-automatic labeling.
+Annotate Segmentation Data.
+The target of annotating segmentation data is to assign a label to each pixel in an image or each point in a LiDAR frame to indicate which object or region it belongs to.
+After labeling, all pixels belonging to an object are annotated with the same class.
+For the manual annotation process, the annotator first draws boundaries around an object and then fills in the area or paints over the pixels directly.
+However, generating pixel/point-level annotations in this way is costly and inefficient.
+Many studies have proposed fully or semi-automatic annotation methods to improve annotation efficiency.
+Barnes et al.
+[139] presented a fully automatic annotation approach based on weakly supervised learning to segment proposed drivable paths in images.
+The semi-automatic annotation method [140] utilizes the objectness priors to generate segmentation masks.
+After that, [141] offered a semi-automatic method considering 20 classes.
+Polygon-RNN++[142] presents an interactive segmentation annotation tool following the idea of [143].
+Instead of using image information to generate pixel-level labels, [137] explores transferring 3D information into 2D image domains to generate semantic segmentation annotations.
+For labeling 3D data, [144] proposes an image-assisted annotation pipeline.
+Luo et al.
+[145] leverage active learning to select a few points and to form a minimal training set to avoid labeling the whole point cloud scenes.
+Liu et al.
+[146] introduce an efficient labeling framework with semi/weakly supervised learning to label outdoor point clouds.
 Annotate Trajectories.
@@
 Usually, the annotating process relies on object detection and tracking results.
-As one of the prior works in trajectory annotation, [280] online generated actions for maneuvers and were annotated into the trajectory.
-[281] consists of a crowd-sourcing step followed by a precise process of expert aggregation.
-[282] developed an active learning framework to annotate driving trajectory.
+As one of the prior works in trajectory annotation, [147] online generates actions for maneuvers and is annotated into the trajectory.
+The annotation approach [148] consists of a crowd-sourcing step followed by a precise process of expert aggregation.
+Jarl et al.
+[149] develop an active learning framework to annotate driving trajectory.
 Precisely anticipating movement patterns of pedestrians is critical for driving safety.
 Styles et al.
-[283] introduced a scalable machine annotation scheme for pedestrian trajectory annotations without human effort.
+[150] introduce a scalable machine annotation scheme for pedestrian trajectory annotations without human effort.
 Annotate on Synthetic Data.
@@
 The generated synthetic scenarios are designed to mimic real-world conditions, including multiple objects, various landscapes, weather conditions, and lighting variations.
-To achieve this goal, some researchers utilized the Grand Theft Auto 5 (GTA5) game engine to build datasets [4, 5].
-[284] presented a real-time system based on multiple games to generate annotations for various AD tasks.
-Instead of applying game videos, SHIFT [10], CAOS [7], and V2XSet [11] were created based on the CARLA [112] simulator.
-Compared to [11], V2X-Sim [81] studied employing multiple simulators [112, 285] to generate dataset for V2X perception tasks.
-CODD [205] further exploited using [112] to generate 3D LiDAR point clouds for cooperative driving.
-Other works [3, 9, 2, 133] leveraged the Unity development platform [286] to generate synthetic datasets.
-VI-B The quality of annotations
-Existing supervised learning-based AD algorithms build upon numerous labeled data.
-However, training on low-quality annotations can negatively affect the safety and reliability of autonomous vehicles.
+Selecting a suitable simulation tool is crucial to achieve this goal.
+Torcs [151] and DeepDriving [152] are two prior works for simulating autonomous vehicles while lacking multi-modality information and other types of objects like pedestrians.
+Recently, the open source simulators, such as CARLA [128], SUMO [153], and AirSim [154], have been widely used in data generation.
+They provide transparent platforms where researchers and developers can freely access and modify the source code, tailoring the simulator to specific requirements.
+In contrast, because of non-open sources, commercial simulation tools like NVIDIA’s Drive Constellation [155] can make it difficult for users to create special driving environments.
+Rather than professional simulators, the game engines, including Grand Theft Auto 5 (GTA5) and Unity [156], are also popular for creating synthetic autonomous driving data.
+Specifically, some researchers utilize the GTA5 game engine to build datasets [4] and [5].
+Krahenbuhl et al.
+[157] present a real-time system based on multiple games to generate annotations for various AD tasks.
+Instead of applying game videos, SHIFT [10], CAOS [7], FedBEVT [158], and V2XSet [11] are created based on the CARLA simulator.
+Compared to [11], V2X-Sim [89] studies employing multiple simulators [128, 153] to generate dataset for V2X perception tasks.
+CODD [159] further exploits using [128] to generate 3D LiDAR point clouds for cooperative driving.
+Other works [3, 9, 2, 160] leverage the Unity development platform to generate synthetic datasets.
+VI-B Quality of Annotations
+Supervised learning-based autonomous driving (AD) algorithms rely heavily on extensive, well-labeled datasets.
+High-quality datasets ensure that these systems can accurately perceive and interpret complex driving environments, enhancing safety and reliability on the road.
+This in turn fosters user trust, a crucial factor for the widespread adoption of autonomous vehicles.
+Conversely, poor dataset quality can result in system errors and safety risks, undermining user confidence, hindering acceptance, and failing to meet the criteria for trustworthy AI as discussed by [161].
 Therefore, ensuring the quality of annotations is fundamental for improving accuracy while driving in complex real-world environments.
+It is even more important to fine-tune the data quality and use active learning methods for dataset curation to get robust performance results on a test set than fine-tuning the model architecture [162].
 Figure 12: Mislabeling example of KITTI [13] dataset.
@@
 Annotation precision is another vital indicator, which refers to whether the labels match the actual state of the objects or scenarios.
-In contrast, correctness highlights that annotated data are appropriate and relevant for the dataset purpose and annotation guidelines.
+In contrast, correctness highlights that annotated data are appropriate and relevant for the purpose of the dataset and annotation guidelines.
 After annotation, it is essential to validate the annotated data to ensure its accuracy and completion.
@@
 Validation helps effectively prevent issues in datasets before they infect the performance of autonomous vehicles, decreasing potential safety risks.
+Inel et al.
 [164] presented a data-agnostic validation method for expert annotated datasets.
@@
 On the left side of the image, the annotation of a car (circled in red) is inaccurate because it does not contain the whole object car.
-Additionally, two cars (highlighted by the green cuboid) are not annotated, even though the camera and LiDAR capture them clearly.
+Additionally, two cars (highlighted by the green rectangle) are not annotated, even though the camera and LiDAR capture them clearly.
+Furthermore, datasets like the IPS300+ [165] contain many labeled objects within a frame (319.84 labels per frame), but on the other hand, the annotations are of bad quality.
+Many large datasets like the Pandaset [166], Oxford [167], CADC [168], nuScenes [22], and Lyft Level 5 [22] were labeled by specialized labeling companies like Scale AI and provide high-quality annotations.
+Labeling the nuScenes dataset took about 7,937 hours and cost 100k USD.
+Another way to label datasets is to use custom labeling tools like 3D BAT [138, 60], which was used to create the TUMTraf dataset.
+The Waymo and KITTI datasets were both labeled with custom labeling tools.
+V2V4Real [87] has used the SUSTechPoints [169] labeling tool to generate the dataset.
 VII Data analysis
@@
 The chart indicates that the USA is at the forefront with 40 datasets (21% share), underscoring its leadership in the autonomous driving domain.
-Germany accounts for 24 datasets, reflecting its robust automotive industry and influence on the advancement of autonomous technology.
-China is closely followed with 16 datasets, suggesting the interest and investment in this field of China.
-Another notable point is that there are 11 datasets collected worldwide and 24 from the European region (except Germany).
+Germany accounts for 24 datasets, and China closely follows with 16 datasets.
+On the other hand, developed countries, including Canada, Korea, the UK, Japan, and Singapore, cover the rest smaller segments.
+Although 11 datasets are collected worldwide and 24 are from the European region (except Germany), all these countries or regions are considered high-income areas.
+The dominance of the USA, western Europe, and East Asia reflects an extremely unbalanced development of autonomous driving around the world.
+Specifically, one of the most classic datasets, KITTI [13], is collected in the metropolitan area of Karlsruhe, Germany.
+In comparison, both Waymo [23] and Argoverse 2 [28] datasets are collected from a wide variety of places, including six different cities in the USA individually.
+Apolloscapes [17] and DAIR-V2X [27] are recorded in China.
+Instead of collecting data solely within a country, nuScenes [22] is established based on the data from America (Boston) and Singapore, both known for complex and highly challenging traffic situations.
+Other well-known autonomous driving datasets [24, 88, 95, 18, 21, 118] are collected in those previously mentioned countries.
+It is noted that thanks to the geographic diversity, [22, 23] have been widely used in transfer learning to verify the generalizability of autonomous driving algorithms.
+Moreover, autonomous driving faces unique challenges across different geographical regions.
+However, relying solely on data from a single source can introduce bias that could result in autonomous vehicles’ failure to perform under varied or unseen regions and cases.
+For instance, the diversity and quantity of electric scooters in China far exceed those in Germany, meaning an algorithm trained exclusively on German data may struggle to accurately recognize objects in China.
+Hence, Recording data from different continents and countries can assist in addressing unique challenges caused by geographical locations.
 This diverse regional distribution enhances the robustness of the collected data and highlights the international efforts and collaborations in the research community and industry.
-On the other hand, although smaller segments represent a range of other countries, including Canada, Korea, the UK, Japan, and Singapore, these countries are developed countries with solid technological backgrounds and accumulation—an extreme regional bias reflected by this statistic.
-The dominance of the USA, western Europe, and East Asia leads to the bias where autonomous driving systems are overfitted to environmental conditions typical of these regions.
-This bias could result in autonomous vehicles’ failure to perform under varied or unseen regions and cases.
-Hence, introducing data from a more expansive array of countries and regions, such as Africa, can assist in the comprehensive development of autonomous vehicles.
 Furthermore, 35 synthetic datasets generated by simulators like CARLA [128] take up 18.32% percent.
 Due to the limitation of recording from real-world driving environments, these synthetic datasets overcome such drawbacks and are critical for exploiting more robust and reliable driving systems.
+However, the domain adaptation from synthetic data to real-world data is still a challenging research topic, which limits a broader application of synthetic data and relevant simulators.
 VII-B Chronological Trends in Perception Datasets
 In Fig.
-10, we introduce a chronological overview of perception datasets with the top 50 impact scores from 2007 to 2023 (until the writing of this paper).
+10, we introduce a chronological overview of perception datasets with the top 50 impact scores from 2009 to 2024 (until the writing of this paper).
 The datasets are color-coded according to their sourcing domain, and synthetic datasets are marked with an external red outline, clearly illustrating the progress toward the diverse data collection strategy.
@@
 Facilitated by the effectiveness of simulators, there are many novel synthetic datasets [7] [10] published in recent years.
-In the timeline, V2X datasets like DAIR-V2X [27] (2021) also exhibit a trend toward cooperative driving systems.
+In the timeline, V2X datasets like DAIR-V2X [27] and TUMTraf ([84, 85, 86, 60]) also exhibit a trend toward cooperative driving systems.
 Furthermore, because of the non-occlusion perspective provided by UAV, drone-based datasets, such as UAVDT [94] published in 2018, take a crucial position in advancing perception systems.
@@
 Nevertheless, as the distance increases, the nuScenes dataset quickly tappers off for the number of objects with annotations.
-The ONCE dataset covers a more even distribution of objects across distances, while the KITTI dataset focuses more on close-range detection.
+The ONCE dataset covers a more even distribution of objects across distances, while the KITTI dataset focuses more on close-range objects.
+VII-D Influence of Adversarial Environmental Conditions.
+TABLE IV: 3D object detection performance under adversarial environmental conditions on the nuScenes [22] dataset.
+We report mAP (%) as the evaluation metric.
+’L’ and ’C’ refer to LiDAR and cameras, respectively.
+The term ’Normal’ indicates that the model was tested across the entire validation set, encompassing all environmental conditions.
+Method Modality Adversarial Conditions mAP $\uparrow$
+VoxelNext [170] L normal 60.5
+nighttime 32.4
+rainy 58.7
+UVTR [171] L normal 60.8
+nighttime 33.9
+rainy 56.5
+Transfusion [172] L+C normal 68.9
+nighttime 37.5
+rainy 65.8
+We further study the influence of adversarial environmental conditions (low lightning and rain) on the performance of 3D object detectors in the autonomous driving system.
+The experiment results are reported in Tab.
+IV.
+We utilize the nuScenes [22] dataset and choose three state-of-the-art methods, VoxelNext [170], UTVR [171], and Transfusion [172].
+For a fair comparison, we directly utilize the pre-trained checkpoints provided in the repository of each approach.
+The rainy and nighttime data are also manually collected from the nuScenes validation set.
+All three methods demonstrate similar trends under different environmental conditions.
+Specifically, the detection accuracy declines significantly under low illumination conditions compared with evaluations of the whole validation set.
+Besides, all models exhibit slight decreases under the non-heavy rainy situations recorded by nuScenes.
+The reliability of detectors in the real world could be much worse once the rain is heavy.
+Therefore, taking image restoration into account is a promising way for camera or sensor fusion-based approaches to overcome these challenges [173, 174].
+In conclusion, increasing the amount of data recorded on various types and intensities of weather conditions is critical for training a robust and reliable autonomous driving system.
 VIII Discussion and Future Works
+TABLE V: VLM autonomous driving dataset.
+OT: object tracking, MOT: multi-object tracking, QA: question-answering, DM: decision-making, IP: intention prediction, VR: visual reasoning, SR: spatial reasoning
 Dataset Year Size Temporal Sensing domain Tasks Real/Synthetic
@@
 NuPrompt [182] 2023 34K frames, 35K prompts $\checkmark$ onboard MOT real
-NuScenes-QA [297] 2023 34K scenes, 460K QA pairs $\checkmark$ onboard visual question answering real
+NuScenes-QA [183] 2023 1K scenarios, 460K QA pairs $\checkmark$ onboard visual QA real
 Refer-KITTI [57] 2023 6,650 frames $\checkmark$ onboard referring MOT real
-Driving LLMs [298] 2023 10K driving situations, 160K QA pairs $\times$ drone visual question answering synthetic
-DRAMA [299] 2023 17,785 scenarios $\checkmark$ onboard reasoning, visual question answering real
+Driving LLMs [184] 2023 10K driving situations, 160K QA pairs $\times$ drone visual QA synthetic
+DRAMA [185] 2023 17,785 scenarios $\checkmark$ onboard reasoning, visual QA real
 Rank2Tell [186] 2023 116 scenarios $\checkmark$ onboard importance level ranking real
@@
 LangAuto CARLAR [188] 2023 64K data clips $\checkmark$ onboard closed-loop driving synthetic
-NuScenes-MQA [303] 2023 34K scenarios, 1.4M QA pairs $\times$ onboard visual question answering real
+NuScenes-MQA [189] 2023 34K scenarios, 1.4M QA pairs $\times$ onboard visual QA real
 DriveMLM [190] 2023 280 hours $\checkmark$ onboard planning, control synthetic
@@
 LiDAR-text [192] 2023 420K 3D captioning data, 280K 3D grounding data $\times$ onboard 3D scene understanding real
-TABLE VII: VLM autonomous driving dataset. OT: object tracking, MOT: multi-object tracking, QA: question-answering
-In this paper, we mainly focus on analyzing the existing datasets, which usually include rich visual data, and aim at tasks in the modular-based pipeline.
-However, with the rapid technological development, especially the excellent performance of the Large Language Models, many novel trends of the next-generation autonomous driving datasets have occurred while proposing new challenges and requirements.
+Talk2BEV [193] 2023 1K BEV scenarios, 20K QA pairs $\checkmark$ onboard DM, IP, VR, SR real
+NuScenes-MQA [194] 2024 1.4M QA pairs, 34,149 scenarios $\checkmark$ onboard visual QA real
+VLAAD [195] 2024 10,379 scenarios $\checkmark$ onboard visual QA, reasoning real
+With rapid technological development, powerful computation resources, and excellent artificial intelligent algorithms, many novel trends in next-generation autonomous driving datasets have occurred while proposing new challenges and requirements.
 End-to-End Driving Datasets.
 Compared to the modular-designed autonomous driving pipeline, the end-to-end architecture simplifies the overall design process and reduces integration complexities.
-The success of the UniAD [307] verifies the potential ability of end-to-end models.
-However, the number of datasets for end-to-end AD is limited [167, 113].
+The success of UniAD [196] verifies the potential ability of end-to-end models.
+However, the number of datasets for end-to-end AD is limited [129].
 Therefore, introducing datasets focusing on end-to-end driving is crucial for advancing autonomous vehicles.
 On the other hand, implementing an automatic labeling pipeline in a data engine can significantly facilitate the development of end-to-end driving frameworks and data [74].
+Potential Applications of AD Datasets Future autonomous driving datasets should provide extensive real-world environmental and traffic data, supporting a wide range of applications beyond the ego-vehicle and basic vehicle-infrastructure cooperation.
+For instance, the interaction data between autonomous vehicles and intelligent infrastructure components can guide the advancement of Internet-of-Thing (IoT)-enabled devices like smart traffic signals.
+Moreover, the detailed insights into traffic patterns, congestion, and vehicle behavior across different times and conditions can facilitate urban planning, optimize the traffic flow, and enhance overall traffic management strategies.
 Introduce Language into AD Datasets.
-The vision language models (VLMs) have recently achieved impressive advancement in many fields.
+Vision language models (VLMs) have recently achieved impressive advancement in many fields.
 Its inherent advantage in providing language information to vision tasks makes autonomous driving systems more explainable and reliable.
-[308] highlighted the prominent role of Multimodal Large Language Models in various AD tasks, such as perception [290, 296], motion planning [301], and motion control [111].
-The autonomous driving datasets including language labels are shown in Tab.
-VII.
+The survey [197] highlights the prominent role of Multimodal Large Language Models in various AD tasks, such as perception [176, 182], motion planning [187], and motion control [193].
+Autonomous driving datasets including language labels are shown in Tab.
+V.
 Overall, incorporating language into AD datasets is a trend of the future development of AD datasets.
 Data Generation via VLMs.
-As mentioned by [309], the powerful capability of VLMs can be used to generate autonomous driving data.
+The powerful capability of VLMs can be used to generate data [198, 199, 200, 201].
 For example, DriveGAN [202] generated high-quality AD data by disentangling different components without supervision.
 Additionally, due to the capability of world models for comprehending driving environments, some works [203, 204, 205] explored world models to generate high-quality driving videos.
-DriveDreamer[312] as a pioneering work derived from real-world scenarios, addressing the limitation of the gaming environments or simulated settings.
+DriveDreamer [204] as a pioneering work derived from real-world scenarios, addressing the limitation of gaming environments or simulated settings.
+The most recent advancements in text-video generation, exemplified by Sora [206], have opened a new avenue for generating autonomous driving data.
+With a brief description, Sora can generate high-quality synthetic data that is very close to real scenes.
+This capability significantly enhances dataset augmentation, especially by extending the data available for rare events like traffic accidents.
+The increased data availability provided by VLMs is poised to enhance the training and evaluation of autonomous driving systems, potentially improving their safety and reliability.
 Domain Adaptation.
 Domain adaptation is a critical challenge in developing autonomous vehicles [207], referring to the ability of a model trained on one dataset (the source domain) to perform stable on another dataset (the target domain).
-This challenge manifests in multiple aspects, such as diversity in environmental conditions [315], sensor settings [316], or synthetic-to-real transform [317].
+This challenge manifests in multiple aspects, such as diversity in driving conditions [208], sensor settings [209], or synthetic-to-real transform [210].
+Therefore, a trend of developing the next-generation datasets is incorporating data from heterogeneous sources.
+First of all, datasets should not only cover a wide range of environmental conditions (various weather conditions and day-night illumination) but also include diverse geographic locations.
+Second, combining data from various sensor types is also pivotal to overcoming the domain adaptation issue.
+Another solution is taking into account blending high-quality synthetic data and real-world data in a balanced way to improve generalization.
+Uncertainty Issues in Autonomous Driving.
+Usually, uncertainty is modeled in a probabilistic way in the field of machine learning.
+The uncertainty can be referred to two types: a) aleatoric uncertainty, where the uncertainty like noise stems from the data, and b) epistemic uncertainty refers to uncertainty caused by unawareness about the best model [211].
+For autonomous driving, one of the major reasons resulting in uncertainty is the incompleteness of training data [212].
+The insufficient training data can not wholly represent the driving environment, causing autonomous vehicles to problematically deal with rare cases during operation.
+Therefore, improving the diversity of the dataset for training reliable autonomous driving systems is crucial.
+Moreover, datasets include rare events and edge cases, allowing models to better understand and quantify the uncertainty associated with unexpected situations and safely handle them.
+Standardization of Data Creation.
+Addressing the standardization of data creation is critical aspect for developing new datasets, as it directly impacts the accuracy, efficiency, and reliability of autonomous vehicle models.
+In general, data standardization refers to the attributes of the data, terminology and structure of the dataset, and data storage [213].
+For the data attributes, unifying data formats across different sensor types and source domains is valuable to facilitate integrated processing and analysis.
+Developing comprehensive guidelines for dataset labeling is crucial to guarantee consistent, high-quality annotations, enhancing model training across diverse datasets and boosting performance reliability.
+Moreover, establishing standardized data storage and access protocols is vital, enabling seamless dataset sharing and integration across multiple sources, and fostering collaboration within the autonomous driving research community.
+Data Privacy.
+The advancement of autonomous vehicles require a lot of data to guarantee the driving safety.
+However, the more the available data, the greater the concern that private data will be abused.
+For the earlier autonomous driving datasets like KITTI [13], there is no anonymization progress for the recorded data, especially for images, which leads to a potential risk for leak private information.
+With the introduction and refinement of relevant regulations in various countries, data anonymization has been widely adopted in recent datasets [84, 22, 23].
+Nevertheless, the information provided by a large dataset is more than individual biographic features or vehicle plates.
+Professional institutes can indirectly extract external information by analyzing the existing information in the data.
+For instance, by analyzing the types of vehicles and the attire of pedestrians in the dataset, one can infer information about the infrastructure, basic construction, and other characteristics of the surrounding area in which the data were collected.
+Open Data Ecosystems.
+The primary objective of open data ecosystems (ODE) for autonomous driving is to foster innovation, enhance transparency, and facilitate collaboration across governments, companies, and research communities.
+This is achieved through the free exchange of datasets, breaking down the traditional barriers that have restricted access to corporations and research institutions.
+By doing so, ODE empowers a wider range of innovators to participate in autonomous driving development, thereby boosting a more diverse and inclusive innovation ecosystem.
+Additionally, ODE establishes dynamic feedback loops where users can report issues, propose improvements, and contribute to enhancing datasets.
+Nevertheless, unrestricted access to data raises strong concerns about data security and privacy.
+Addressing these concerns necessitates carefully crafting and continuously refining relevant legal frameworks to safeguard sensitive information while prompting the growth of ODE.
+The balance is pivotal for maintaining the integrity of open data ecosystems and ensuring their sustainable development in autonomous driving.
 IX Conclusion
-In this paper, we exhaustively and systematically reviewed and analyzed more than 200 existing autonomous driving datasets.
-We started with the sensor types and modalities, sensing domains, and tasks relevant to autonomous driving datasets.
+In this paper, we exhaustively and systematically reviewed 265 existing autonomous driving datasets.
+We started with the sensor types and modalities, sensing domains, and tasks relevant to autonomous driving.
 We introduced a novel evaluation metric called impact score to validate the influence and importance of perception datasets.
-Afterward, we demonstrated several high-impact datasets involving perception, prediction, planning, control, and end-to-end autonomous driving.
-Additionally, we explained the annotation methodology for autonomous driving datasets and investigated the factors affecting annotation quality.
-Moreover, we delineated the chronological and geographical distribution of the collected datasets, providing a comprehensive perspective for understanding the current development of autonomous driving datasets.
-Meanwhile, we studied the data distribution of several datasets, offering a specific viewpoint for comprehending the variances across different datasets.
-In the end, we discussed the development and trend of the next generation of autonomous driving datasets.
+Our analysis covered the attributes and significance of high-impact datasets across perception, prediction, planning, control, and end-to-end driving tasks.
+Additionally, we investigated the annotation methodologies and the factors affecting annotation quality.
+We also analyzed datasets from chronological and geographical perspectives to understand current trends and conducted experiments demonstrating the crucial need for data diversity under adversarial conditions.
+With the study on data distribution, we offered a specific viewpoint for comprehending the variances across different datasets.
+Our findings underscore the essential role of diverse, high-quality datasets in shaping the future of autonomous driving.
+Looking forward, we highlighted key challenges and future directions for autonomous driving datasets, including the adoption of VLM, domain adaptation, uncertainty challenges, standardization, data privacy, and open data ecosystems.
+These areas represent promising avenues for future research and are pivotal for advancing autonomous driving technology, setting a clear roadmap for further innovation in this rapidly evolving field.
+Supplementary Tables of AD Datasets
+TABLE VI: Part I of Autonomous Driving Datasets. SS: semantic segmentation, OD: object detection, OT: object tracking, MOT: multi-object tracking
+Dataset Year Size Temporal Sensing domain Tasks Real/Synthetic
+ETH Ped [214] 2007 2,293 frames $\checkmark$ onboard pedestrian detection real
+TUD-Brussels [215] 2009 1,600 frames $\times$ onboard (3D) OD real
+Collective activity [216] 2009 44 short videos $\checkmark$ others human activity recognition real
+San Francisco Landmark [217] 2011 150K panoramic images $\times$ others landmark identification real
+Daimler Stereo Ped [218] 2011 21,790 frames $\checkmark$ onboard pedestrian detection real
+BelgiumTS [219] 2011 13K traffic sign annotations $\times$ onboard traffic sign detection real
+Stanford Tack [220] 2011 14K tracks $\checkmark$ onboard classification real
+TME Motorway [221] 2012 30K frames $\checkmark$ onboard OD, OT real
+MSLU [222] 2013 36.8 km distances $\checkmark$ onboard SLAM real
+SydneyUrbanObject [223] 2013 588 object scans $\times$ onboard classification real
+Ground Truth SitXel [224] 2013 78,500 frames $\checkmark$ onboard stereo confidence real
+NYC3DCars [225] 2013 2K images $\times$ onboard OD real
+AMUSE [226] 2013 117,440 frames $\checkmark$ onboard SLAM real
+Daimler Ped [227] 2013 12,485 frames $\checkmark$ onboard pedestrian path prediction real
+LISA [228] 2014 6,610 frames $\checkmark$ onboard traffic sign detection real
+Paris-rue-Madame [229] 2014 643 objects $\times$ onboard OD, SS real
+TRANCOS [230] 2015 1.2K images $\times$ V2X onboard number estimation real
+FlyingThings3D [231] 2015 26,066 frames $\checkmark$ others scene flow estimation synthetic
+Ua-detrac [232] 2015 140K frames $\checkmark$ V2X OD, MOT real
+NCLT [233] 2015 34.9 hours $\checkmark$ onboard odometry real
+CCSAD [234] 2015 96K frames $\checkmark$ onboard scene understanding real
+KAIST MPD [235] 2015 95K color-thermal pair frames $\times$ onboard pedestrian detection real
+SAP [236] 2016 19K frames $\times$ drone OD, OT real
+LostAndFound [237] 2016 2,104 frames $\checkmark$ onboard obstacle detection real
+UAH-Driveset [238] 2016 500 mins $\checkmark$ onboard lane detection, detection real
+CURE-TSR [239] 2017 2.2M annotated images $\times$ onboard traffic sign detection real
+TuSimple [240] 2017 6,408 frames $\times$ onboard lane detection, velocity estimation real
+TRoM [241] 2017 712 frames $\times$ onboard road marking detection real
+NEXET [242] 2017 91,190 frames $\times$ onboard OD real
+DIML [243] 2017 470 videos $\checkmark$ onboard lane detection real
+Bosch STL [244] 2017 13,334 images $\checkmark$ onboard traffic light detection and classification real
+Complex Urban [245] 2017 around 190km paths onboard SLAM real
+DDD20 [246] 2017 51 hours event frames $\checkmark$ onboard end-to-end driving real
+PedX [247] 2018 5K images $\checkmark$ onboard pedestrian detection and tracking real
+LiVi-Set [248] 2018 10K frames $\checkmark$ onboard driving behavior prediction real
+Syncapes [249] 2018 25K images $\times$ onboard OD, SS synthetic
+NVSEC [250] 2018 around 28 km distances $\checkmark$ others SLAM real
+Aachen Day-Night [251] 2018 4,328 images, 1.65M points $\checkmark$ onboard visual localization real
+CADP [252] 2018 1,416 scenes $\times$ V2X traffic accident analysis real
+TAF-BW [253] 2018 2 scenarios $\checkmark$ V2X MOT, V2X communication real
+comma2k19 [254] 2018 2M images $\times$ onboard pose estimation, end-to-end driving real
+nighttime drive [255] 2018 35K $\times$ onboard SS real
+VEIS [160] 2018 61,305 frames $\times$ onboard OD, SS synthetic
+Paris-Lille-3D [256] 2018 2,479 frames $\times$ onboard 3D SS, classification real
+NightOwis [257] 2018 56K frames $\checkmark$ onboard pedestrian detection, tracking real
+EuroCity Persons [258] 2018 47,300 frames $\times$ onboard OD real
+RANUS [259] 2018 4K frames $\times$ onboard SS, scene understanding real
+SynthCity [260] 2019 367.9M points in 30 scans $\times$ onboard (3D) OD, (3D) SS synthetic
+D2-City [261] 2019 700K annotated frames $\checkmark$ onboard OD, MOT real
+Caltech Lanes [262] 2019 1,224 frames $\times$ onboard lane detection real
+Mcity [263] 2019 1,7500 frames $\checkmark$ onboard SS real
+DET [264] 2019 5,424 event-based camera images $\times$ onboard lane detection real
+PreSIL [8] 2019 50K frames $\checkmark$ onboard OD, 3D SS synthetic
+H3D [265] 2019 27,721 frames $\checkmark$ onboard (3D) OD, MOT real
+LLAMAS [266] 2019 100K frames $\checkmark$ onboard lane segmentation real
+MIT DriveSeg [267] 2019 5K frames $\times$ onboard SS real
+Astyx [268] 2019 500 radar frames $\times$ onboard 3D OD real
+UNDD [269] 2019 7.2K frames $\checkmark$ onboard SS real
+Boxy [270] 2019 200K frames $\checkmark$ onboard OD real
+RUGD [271] 2019 37K frames $\checkmark$ others SS real
+ApolloCar3D [272] 2019 5,277 driving images $\times$ onboard 3D instance understanding real
+HEV [273] 2019 230 video clips $\checkmark$ onboard object localization real
+HAD [274] 2019 5,675 video clips $\checkmark$ onboard end-to-end driving real
+CARLA-100 [275] 2019 100 hours driving $\checkmark$ onboard path planning, behavior cloning synthetic
+Brno Urban [276] 2019 375.7 km $\checkmark$ onboard recognition real
+VERI-Wild [277] 2019 416,314 images $\checkmark$ V2X onboard re-identification real
+CityFlow [278] 2019 200K bounding boxes $\checkmark$ V2X OD, MOT, re-identification real
+VLMV [279] 2020 306K frames $\checkmark$ V2X lane merge real
+Small Obstacles [280] 2020 3K frames $\times$ onboard small obstacle segmentation real
+Cirrus [281] 2020 6,285 frames $\checkmark$ onboard (3D) OD real
+KITTI InstanceMotSeg [282] 2020 12,919 frames $\checkmark$ onboard moving instance segmentation real
+A*3D [283] 2020 39,179 point cloud frames $\times$ onboard (3D) OD real
+Toronto-3D [284] 2020 4 scenarios $\times$ onboard 3D SS real
+MIT-AVT [285] 2020 1.15M 10s video clips $\checkmark$ onboard SS, anomaly detection real
+CADC [168] 2020 56K $\times$ onboard (3D) OD, OT real
+SemanticPOSS [286] 2020 2,988 point cloud frames $\checkmark$ onboard 3D SS synthetic
+IDDA [287] 2020 1M frames $\times$ onboard segmentation synthetic
+TABLE VII: Part II of Autonomous Driving Datasets
+Dataset Year Size Temporal Sensing domain Tasks Real/Synthetic
+CARRADA [288] 2020 7,193 radar frames $\checkmark$ onboard SS real
+Titan [289] 2020 75,262 frames $\checkmark$ onboard OD, action recognition real
+NightCity [290] 2020 4,297 frames $\times$ onboard nighttime SS real
+PePScenes [111] 2020 40K frames $\checkmark$ onboard (3D) OD, pedestrian action prediction real
+DDAD [291] 2020 21,200 frames $\times$ onboard depth estimation real
+MulRan [292] 2020 41.2km paths $\checkmark$ onboard place recognition real
+Oxford Radar RobotCar [167] 2020 240K scans $\checkmark$ onboard odometry real
+OTOH [125] 2020 170K scenes $\checkmark$ drone trajectory prediction, planning real
+DA4AD [293] 2020 9 sequences $\checkmark$ onboard visual localization real
+CPIS [92] 2020 10K frames $\times$ V2X cooperative 3D OD synthetic
+EU LTD [294] 2020 around 37 hours $\checkmark$ onboard odometry real
+Newer College [295] 2020 290M points, 2300 seconds $\checkmark$ others SLAM real
+CCD [296] 2020 4.5K videos $\checkmark$ onboard accident prediction real
+LIBRE [297] 2020 40 frames $\times$ onboard LiDAR performance benchmark real
+Gated2Depth [298] 2020 17,686 frames $\checkmark$ onboard depth estimation real
+TCGR [299] 2020 839,350 frames $\checkmark$ others traffic control gesture recognition real
+DSEC [47] 2021 53 sequences (3193 seconds in total) $\checkmark$ onboard dynamic perception real
+4Seasons [300] 2021 350km recordings $\checkmark$ onboard SLAM real
+PVDN [301] 2021 59,746 frames $\checkmark$ onboard nighttime OD, OT real
+ACDC [80] 2021 4,006 images $\times$ onboard SS on adverse conditions real
+DRIV100 [302] 2021 100K frames $\times$ onboard domain adaptation SS real
+NEOLIX [303] 2021 30K frames $\checkmark$ onboard 3D OD, OT real
+IPS3000+ [165] 2021 14,198 frames $\checkmark$ V2X 3D OD real
+AUTOMATUM [304] 2021 30 hours $\checkmark$ drone trajectory prediction real
+DurLAR [305] 2021 100K frames $\checkmark$ onboard depth estimation real
+Reasonable-Crowd [306] 2021 92 scenarios $\checkmark$ onboard driving behavior prediction synthetic
+MOTSynth [307] 2021 768 driving sequences $\checkmark$ onboard Pedestrian detection and tracking synthetic
+MAVD [308] 2021 113,283 images $\checkmark$ onboard OD and OT with sound real
+Multifog KITTI [309] 2021 15K frames $\times$ onboard 3D OD synthetic
+Comap [310] 2021 4,391 frames $\checkmark$ V2X 3D OD synthetic
+R3 [311] 2021 369 scenes $\times$ onboard out-of-distribution detection real
+WIBAM [312] 2021 33,092 frames $\checkmark$ V2X 3D OD real
+CeyMo [313] 2021 2,887 frames $\times$ onboard road marking detection real
+RaidaR [314] 2021 58,542 rainy street scenes $\times$ onboard SS real
+Fishyscapes [315] 2021 1,030 frames $\times$ onboard SS, anomaly detection real
+RadarScenes [316] 2021 40K radar frames $\checkmark$ onboard OD, classification real
+ROAD [317] 2021 122K frames $\checkmark$ onboard OD, SS real
+All-in-One Drive [46] 2021 100K $\checkmark$ onboard (3D) OD, (3D) SS, trajectory prediction real
+PandaSet [166] 2021 8,240 frames $\checkmark$ onboard (3D) OD, SS, OT real
+SODA10M [318] 2021 20K labeled images $\checkmark$ onboard OD real
+PixSet [319] 2021 29K point cloud frames $\times$ onboard (3D) OD real
+RoadObstacle21 [320] 2021 327 scenes $\times$ onboard anomaly segmentation synthetic
+VIL-100 [321] 2021 10K frames $\times$ onboard lane detection real
+OpenMPD [322] 2021 15K frames $\times$ onboard (3D) OD, 3D OT, semantic segmentation real
+WADS [323] 2021 1K point cloud frames $\checkmark$ onboard SS real
+CCTSDB 2021 [324] 2021 16,356 frames $\times$ onboard traffic sign detection real
+SemanticUSL [325] 2021 1.2K frames $\times$ onboard domain adaptation 3D SS real
+CARLANE [326] 2022 118K frames $\checkmark$ onboard lane detection synthetic
+CrashD [327] 2022 15,340 scenes $\times$ onboard 3D OD synthetic
+CODD [159] 2022 108 sequences $\checkmark$ V2X multi-agent SLAM synthetic
+CarlaScenes [328] 2022 7 sequences $\checkmark$ onboard (3D) SS, SLAM, depth estimation synthetic
+OPV2V [329] 2022 11,464 frames $\times$ V2X onboard-to-onboard perception synthetic
+CARLA-WildLife [330] 2022 26 videos $\checkmark$ onboard out-of-distribution tracking synthetic
+SOS [330] 2022 20 videos $\checkmark$ onboard out-of-distribution tracking real
+RoadSaW [331] 2022 720K frames $\checkmark$ onboard Road surface and wetness estimation real
+I see you [332] 2022 170 sequences, 340 trajectories $\checkmark$ V2X OD real
+ASAP [333] 2022 1.2M images $\checkmark$ onboard online 3D OD real
+Amodal Cityscapes [334] 2022 5K frames $\times$ onboard amodal SS real
+SynWoodScape [335] 2022 80K frames $\times$ onboard (3D) OD, segmentation synthetic
+TJ4RadSet [336] 2022 7,757 frames $\checkmark$ onboard OD, OT real
+CODA [337] 2022 1,500 frames $\times$ onboard corner case detection real
+LiDAR Snowfall [338] 2022 7,385 point cloud frames $\checkmark$ onboard 3D OD synthetic
+MUAD [339] 2022 10.4K frames $\checkmark$ onboard OD. SS, depth estimation synthetic
+AUTOCASTSIM [340] 2022 52K frames $\checkmark$ V2X (3D) OD, OT, SS real
+CARTI [341] 2022 11K frames $\checkmark$ V2X cooperative perception synthetic
+K-Lane [342] 2022 15,382 frames $\times$ onboard lane detection real
+Ithaca365 [343] 2022 7K frames $\times$ onboard 3D OD, SS, depth estimation real
+GLARE [344] 2022 2,157 frames $\times$ onboard traffic sign detection real
+SUPS [345] 2023 5K frames $\checkmark$ onboard SS, depth estimation, SLAM synthetic
+Boreas [346] 2023 7,111 frames $\checkmark$ onboard (3D) OD, localization real
+Robo3D [347] 2023 476K frames $\checkmark$ onboard (3D) OD, 3D SS real
+ZOD [31] 2023 100K frames $\checkmark$ onboard (3D) OD, segmentation real
+K-Radar [348] 2023 35K radar frames $\times$ onboard 3D OD, OT real
+aiMotive [349] 2023 26,583 frames $\checkmark$ onboard 3D OD, MOT real
+UrbanLaneGraph [350] 2023 around 5,220 km lane spans $\checkmark$ drone lane graph estimation real
+WEDGE [351] 2023 3,360 frames $\times$ others OD, classification synthetic
+OpenLane-V2 [352] 2023 466K images $\checkmark$ onboard lane detection, scene understanding real
+V2X-Seq (perception) [56] 2023 15K frames $\checkmark$ V2X cooperative perception real
+SSCBENCH [353] 2023 66,913 frames $\times$ onboard semantic scene completion real
+RoadSC [354] 2023 90,759 images $\times$ onboard road snow coverage estimation real
+V2X-Real [355] 2024 171K images, 33K LiDAR point clouds - V2X 3D OD real
+RCooper [356] 2024 50K images, 30K LiDAR point clouds $\checkmark$ V2X 3D OD, OT real
+FLIR [48] - 26,442 thermal frames $\checkmark$ onboard thermal image OD real